A Complete Bibliography of Publications in Canadian Journal of Mathematics = Journal canadien de

mathématiques for the decade 1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.07

Title word cross-reference

(0, 1) [457]. (AcapB)^{n} = A^{n}capB^{n} [340]. (F, d_{n}) [258]. (G, σ) [47]. (L^{2}X, DX) [491]. (modK) [733]. (p; , q) [269]. −d^{2}/dx^{2} + q [125]. 0, 1 [875]. 0 − 2 [928, 1022]. 0 < p < 1 [1017]. 1 [832, 1084, 1038]. 1 − ALG [459]. 2 [858, 520, 985, 1044, 143]. 2 < p < ∞ [422]. 2p [864]. 2 × 2 [707]. 3 [997, 1040, 595, 223, 229, 513, 994]. 4 [997]. 5 [595, 151]. 6 [463]. [f, g] [603]. A [538, 441, 189, 884]. A_{λ} [603]. α [1228, 451]. αn^{2} + βn [1038]. AP [209]. ∗ [1148, 1129]. B [651]. B_{p}(G) [739]. β [451, 344]. bfC^{n} [1162]. C [538, 203]. C(S) [539, 1131, 9]. C(X) [908, 305, 847]. C[0, 1] [643]. C^{∗} [1062, 651, 836]. C^{0} [485]. C^{∗} [308, 1280, 535, 548, 1253, 194, 589]. C^{∞} [468]. C^{n} [1149, 1207]. c_{0} [310]. F [1006]. L [1001]. D_{1} [899]. δ [322]. Δu = P u [866]. E [265]. E^{3} [475, 459, 398]. E^{n} [178]. η [211]. exp(−midxmid) [1119]. F [1058, 654, 747, 638, 425]. F K [496]. G [1226, 610]. G_{δ} [806]. GA [966]. Γ_{0}, (16p) [250]. Γ_{0}, (n) [250]. K, CV_{k}(η), S [498]. M [368]. U [769]. U_{α} [498]. GF(2^{y}) [538]. GF(q) [711]. GL(n), [494]. GL_{2}(F) [1124]. H [762, 1305, 1298, 1244, 1273]. H − B [505]. H^{1} [251, 355]. H^{2}(G, Z_{p}) [624]. H^{∞} [1162]. H^{∞} + C [1207]. H^{p} [856]. H^{p}(R_{+}^{n+1}) [452]. H^{p}(S) [422]. H^{p}(U^{n}) [401]. H_{1} [899, 45]. H_{∞} [45]. HNN [511, 217]. Hom(G, H) [22]. Hom(V, V) [328]. I [630]. i, d/dx [125]. J_{p} [83, 84]. K [446, 255, 344, 1285, 934, 1281, 128]. k[X, Y] [1106]. K_{0} [1274]. K_{1} [1274]. k_{3} [774]. KB [933]. L [1222, 1268, 1199, 779, 4]. L^{2}(SO(n)/SO(n − m)) [714]. L^{p} [360, 931, 694]. L^{p}(0, ∞) [433]. L^{p}(G) [1096]. L^{p}(G), [1017]. L_{n} [679]. L_{p} [790, 644, 938]. L_{p}(G) [798]. lambda [1260]. lambda = 0 [1040]. lambda = 1 [437]. lambda = 2 [437]. lambda = 3 [474]. lambda = au + 1 [460]. λ = ∞ [493]. leq2 [427]. lnc [818]. M [43, 438, 480, 41, 819, 923]. M^{2} × S^{1} [66]. M^{3} [657]. M_{22} [911]. mod2 [760]. N [1093, 563, 620, 818, 949, 1133, 397, 740, 454, 728, 1241, 1262, 177]. n + 1 [818]. O [318]. O(n) [494]. ω [605]. ω_{1} [1194]. OP [131]. P [984, 998, 563, 1013, 864, 267, 395, 991, 591, 253, 254, 732, 348, 832, 188, 205, 583, 1229, 1279, 1231]. p < 1 [452]. p^{2} [1104]. P^{n} [1154]. P^{n+1} [528]. Φ [1284, 255]. pi [1248, 676]. π_{1}(M^{3}) [301]. pq^{2} + 1 [1116]. PSL(2, C) [195]. PSL(2, m) [382]. PSL(3, q) [443]. PSL(n, q) [157]. PSp_{4}(3) [449]. PSU(3, q^{2}) [443]. q [1026]. Q_{2}(G) [428]. Q_{p} [576]. QF − 1 [427]. R [666, 426, 900, 287, 538]. r(F, K_{m}) [747]. R(X) [893]. R^{2} [685, 878]. R^{d} [923]. R^{n} [1105]. R^{n} × L^{p} [1165]. R_{0} [24]. S [347, 301, 311, 865]. S(X) [1016]. S(Y) [1016]. S^{1} [878]. S^{3} [156]. §^{p} × §^{q} [299]. S_{n}^{−1} [1036]. S_{s}, [494]. SC_{n}P [528]. Σ [735, 1056, 937, 1269, 1194]. SJightarrowSF [760]. SL(2) [1268]. SL(2, C) [1208]. SL(2, m) [382]. SL(2, R) [331]. SL(3, 3^{n}) [778]. SL(3, q) [443]. SL(4, 3^{n}) [778]. Sp(4, 3^{n}) [778]. Sp(n) [494]. Spec [900]. SU(2) [331]. SU(3, 3^{n}) [778]. SU(3, q^{2}) [443]. ∑ A_{k}XB_{k} ≡ 0 [1278]. T [514]. T^{n} [1087]. T_{0} [516, 715]. T_{0} + T_{5} [517]. θn^{2} [91]. U [52, 648]. ULC [398]. V [468, 482]. V_{p} [892]. W^{∗} [525]. X [847]. XAX^{T} = C [538]. y′′ + p(t)f(y, y′) = 0 [425]. z [889]. Z/4 [1160]. zeta [484]. |βΦ − Φ| = 1 [1284].

-additive [1269]. -adic [563, 205, 1229]. -algebra [1280, 535, 194, 589, 651, 525, 836].

-algebras [1056, 1093, 308, 1062, 548, 1253]. -algèbres [1253]. -ary [1262]. -closed [24, 287]. -connected [595]. -construction [344]. -continuous [322]. -convex [819, 923]. -cubes [934]. -degenerate [128]. -determined [468]. -dimensional [728, 1084, 143, 318]. -dual [651]. -embedded [889, 1274]. -equivalence [762]. -extension [188]. -finite [1298, 735]. -Fourier [931]. -function [610]. -functions [605, 484, 1222, 311, 865]. -functors [265]. -geometries [1260]. -graded [985]. -group [1199, 348, 779]. -groups [998, 395, 991, 591, 253, 254, 4, 583, 769]. -high [563]. -homogeneous [928, 1022]. -ideals [1194]. -index [83, 84]. -indistinguishability [1268]. -integral [528, 209, 1154]. -invariant [648]. -irreducible [994]. -isomorphisms [426]. -Leviflat [1281]. -manifolds [997, 223, 229, 513]. -mappings [468]. -matrices [457, 203]. -maximal [301]. -metric [397]. -multiplication [482]. -norms [255]. -ominoes [454]. -orthogonality [514]. -points [937]. -polytopes [997, 595]. -primary [438, 43]. -proper [441, 189]. -pure [451]. -pure-injective [1013]. -ring [638]. -rings [1248, 1148, 1026, 347, 52]. -semigroups [1129]. -separating [666]. -series [1279, 620]. -space [1241]. -spaces [732, 774, 790, 908, 446, 1058, 1305, 715, 984, 480, 654, 1244, 1273]. -speedable [1228]. -spline [1001]. -stabilité [1226]. -stability [1226]. -stable [676]. -Stone [368]. -subgroup [1044]. -subgroups [267]. -sufficiency [485]. -th [1133, 832]. -theory [344, 1285]. -valued [875]. -varieties [858]. -ways [463].

1 [1019]. 17 [737].

22 [500]. 23 [556, 513]. 24 [959]. 26

[759, 802, 1070]. 27 [1164]. 28 [1022, 977].

29 [1145, 1146].

3-dimensional [965].

= [286].

Aardenne [1252]. abelian [524, 138, 980, 532, 767, 192, 1168, 1129, 537, 12, 152, 779, 232, 174, 365, 366, 1041, 1166, 472, 677, 698, 608, 132, 64, 451, 320, 945]. abélien [767]. Absolute [680, 334, 533, 71, 29, 30, 240, 25]. abstract [39, 1112, 216, 768, 1163, 771, 101]. abstraite [39]. action [1148, 1074, 874, 1312]. Actions [190, 268, 36, 1256, 370]. acyclic [720]. Addendum [1070, 240, 1022, 959]. addition [12, 139]. Additive [586, 638, 986, 197, 411, 1269, 559]. Additivity [1154]. adic [563, 205, 1229]. adjoint [1075, 213]. adjunction [1066]. admissible [179]. admitting [1044, 1041]. AF [1291]. AF-algebras [1291]. Affine [1108, 275, 705, 76, 835, 1220]. affinities [1238]. Alexandroff [957]. algebra [407, 1280, 974, 181, 535, 574, 194, 589, 571, 214, 651, 1157, 525, 213, 204, 836, 170, 775, 776, 756, 326, 361]. Algebraic [135, 915, 278, 1108, 354, 19, 124, 1244, 1273]. Algebraically [627]. Algebras [983, 1309, 481, 697, 770, 1056, 1093, 1286, 843, 118, 783, 844, 960, 308, 1062, 647, 1294, 839, 329, 281, 89, 362, 985, 1198, 567, 486, 548, 1253, 634, 373, 1174, 1101, 333, 2, 415, 982, 126, 585, 367, 1298, 693, 826, 462, 582, 824, 317, 1004, 17, 526, 98, 391, 1276, 874, 1312, 1291, 1157, 169, 179, 698, 976, 379, 316, 577, 1231, 786, 687, 107]. Algèbres [126, 1253]. algebroidal [1111]. algorithms [1257]. Almost [1287, 569, 984, 885, 1308, 1310, 1142, 1147, 423, 709, 642, 1269]. Almost- [984]. almost-fixed-point [1147]. Almost-Riemannian [885]. along [50, 942]. alternate [538]. alternating [940, 877]. alternative [1232, 46, 332, 112, 391]. altitude [1172]. amalgamated [249]. amarts [1293]. ambivalent [708]. Amicable [920]. Amitsur [283]. analog [988]. Analogues [713, 576]. analysis [830]. Analytic [349, 1264, 1162, 688, 713, 270, 530, 59, 794, 593, 441, 324, 1216, 79, 146]. analytiques [530]. André [82]. Angle [420]. anneau [237]. anneaux [492, 805]. annuli [657]. Anosov [380]. any [1220]. Apollonius [857]. application [830, 1257, 125, 422, 812, 372, 671, 559, 558, 1098]. Applications [980, 776, 656, 93, 794, 1219, 1167, 551, 585, 534, 321, 1113, 21, 717, 958, 504, 49, 139, 456]. applied [846, 746]. approach [1232, 6, 75, 241]. approximate [1045, 473, 837]. approximately [31]. Approximating [1045]. Approximation [172, 261, 324, 279, 1035, 914, 1119, 834, 1028, 853, 1204, 414, 972, 1190]. approximations [833]. arbitrary [1133, 367, 1078, 471, 741, 615, 203]. arboricity [176]. arc [412, 1239, 92]. Archimedean [859, 1199, 289]. arcs [94]. Arf [63]. Arhangel’skii [732]. arithmetic [201]. arithmetical [671]. Aronszajn [1197]. arrays [1203]. Artin [1157]. Artinian [1255, 421, 584, 1158, 312]. ary [1262]. ascendant [1057]. ascending [682]. Ascoli [774]. Associated [343, 921, 111, 110, 430, 168, 323, 562]. associative [669, 729]. Asymptotic [942, 177, 50, 214, 791, 1203, 444]. asymptotiques [444]. atriodic [1090]. augmentation [282, 620]. Automatic [1134]. Automorphism [373, 1004, 536, 1010, 1044, 846, 519]. automorphisms [41, 928, 1022, 991, 1198, 220, 950, 1210, 109]. Averaging [908, 396]. avoidance [769]. axe [914]. axiom [1110]. Axioms [397, 25]. axis [914, 38, 972]. axisymmetric [1204]. b. [671]. Bachmuth [950]. backs [850, 958]. Baer [103, 275, 705, 1176]. Baire [1275, 1264]. balanced [11, 1, 26]. ball [1149, 1089, 1207]. Banach [1042, 110, 539, 649, 353, 826, 462, 734, 1051, 116, 1261, 165, 885, 786]. base [957]. based [1052]. bases [733, 1193, 982, 376, 1197, 793]. Basic [278]. basically [1284]. basis [144, 1058, 323, 654]. Baskakov [1225]. Bass [158]. be [532, 943, 1158, 213]. behavior [477, 1138, 799]. behaviour [50, 1203, 177]. Bergman [1149, 1303]. Bernoulli [1135, 837]. Bernstein [886, 1237]. Bessel [1073]. Best [1035, 206]. between [603, 1053, 1096, 1250, 4, 502, 185]. biaxially [894]. Bicommutators [167, 556]. Biconnected [196]. biorthogonal [255, 597]. bipartite [442, 1205]. Biplanar [1242]. Bireflectionality [1069]. bisection [420]. Blaschke [336]. Bloch [731, 987]. Block [1288, 11, 1, 26]. Block-finite [1288]. Blocking [1161]. blocks [1012, 1205]. Blow [949]. Bochner [701]. Bohr [219]. Bonnet [472]. Boolean [839, 181, 1004, 607]. Boolean-valued [607]. border [530]. Borel [367, 221, 589, 241, 559]. Borel-like [367]. Bott [804]. boule [1149]. bound [52, 454, 624]. Boundaries [826]. Boundary [703, 477, 871, 1112, 225, 658, 581, 702, 613, 1138, 503, 1190, 522, 146, 872, 504, 1025]. Bounded [1079, 753, 719, 882, 499, 1206, 1014, 110, 866, 1082, 418, 121, 856, 99, 658, 581, 702, 722, 1138, 503, 496, 324]. bounded-homomorphic [1206]. Boundedness [1099, 510]. Bounding [1306]. Bounds [948, 1243, 1178]. Bovdi [248]. BP [1091]. BP-coaction [1091]. branched [156]. Brandt [114]. Brauer [1312, 874, 599, 785]. Britton [884]. Brooks [342]. Brown [737, 959, 737, 211]. Brownian [77]. Bruijn [1252]. bundles [990, 277, 472, 1120]. Burnside [1104].

C [737, 959, 809]. cône [469]. calculus [1009, 803]. Calkin [974]. can [532]. Canad [556, 759, 737, 1022, 802, 1070, 513, 1145, 1146, 959, 1164, 977, 500]. canonical [1031, 51]. Canonically [866]. Cantor [639]. Cantor-like [639]. Capacity [639]. Caplygin [504]. caractère [298]. Carathéodory [1034]. Cardan [684]. Cardinal [852, 1254, 89, 1001]. cardinality [346]. Cartan [783]. Cartesian [817, 626]. Cartier [354]. case [460, 474, 493]. cases [437]. categories [596, 1111, 817, 432]. Category [1011, 1275, 134, 12, 715, 716]. catenoid [706]. Cech [321]. Cell [33]. cells [1305]. cellular [1182]. Center [727]. Centered [1197]. Centers [499]. Central [85, 265, 486, 591, 835, 729]. centralizer [983]. centralizers [1249, 1286]. centres [585]. Cercles [50]. Certain [751, 1072, 312, 1039, 345, 806, 733, 41, 1103, 915, 237, 795, 1148, 762, 991, 282, 161, 635, 685, 119, 1162, 693, 145, 293, 483, 178, 399, 37, 1290, 175, 1227, 394, 510, 694, 502, 508, 242, 519, 1137, 786, 631]. certains [237, 739]. Cesàro [240, 924, 29, 30]. Chain [1256, 162, 1177, 405]. Chains [549, 6, 557, 682]. Changes [1032, 1056]. Character [696, 876, 389, 1251, 443, 298]. characteristic [676, 1133, 971, 1301, 367, 545, 1231]. Characteristically [169]. Characterization [806, 56, 559, 770, 246, 947, 557, 193, 449, 700, 31, 327, 1024, 1029, 793]. Characterizations [1206, 215, 1100, 860, 1293, 54, 289]. characterized [1209]. characterizes [1307]. characterizing [1054, 1311]. characters [940, 123, 599, 785, 243, 512]. Chebyshev [886]. Chevalley [863]. choices [734]. Chromatic [437, 460, 474, 493, 614, 306]. circle [909, 475]. circle-like [475]. circles [882]. circuitous [50]. circuits [6, 782]. class [909, 859, 58, 457, 864, 795, 265, 95, 796, 430, 148, 798, 1127, 1184, 606, 887, 944, 1020, 642, 799, 559, 1048, 251, 14, 1216, 205, 967, 892, 941, 458, 789, 7]. Classes [430, 806, 435, 1149, 90, 1035, 1042, 762, 530, 951, 24, 1002, 281, 1053, 104, 1293, 245, 1117, 545, 303, 242, 631]. classical [1069, 792]. Classification [1233, 855, 858, 831, 1253, 814, 990, 597, 402, 107]. classifications [1253]. Clifford [243, 512]. closed [338, 459, 627, 881, 24, 110, 1028, 383, 1114, 78, 1100, 1250, 833, 1190, 1178, 287, 1213]. closedness [817]. closure [1081, 686, 1176, 433]. closures [679]. cluster [383]. clusters [761]. coaction [1091]. coalgebras [976]. cobordism [263]. cocategory [716]. codes [228, 1150]. codimension [520, 133]. Coding [837]. Coefficient [58, 799, 736]. Coefficients [658, 561, 759, 1014, 892, 227]. cofaithful [167, 556]. cofibre [1068]. Cohen [376]. cohereditary [488]. cohomological [417]. Cohomology [367, 1062, 283, 869, 491]. coincidence [196]. Coinitial [1210]. colimits [738]. collections [153, 399]. collectionwise [1128, 883]. collective [206]. collineation [705, 113]. collineations [835]. colocalization [750]. Colorful [1254]. colourable [954]. Coloured [69, 274]. colourings [1126]. combination [498]. combinations [944, 227]. Combinatorial [1152, 562, 1289, 997, 70, 905]. combinatorics [1232]. Comments [1208]. common [445, 552, 800, 1164]. commutant [1214, 519]. Commutative [130, 154, 1093, 297, 1122, 42, 230, 938, 46, 332, 959, 462, 880, 642, 245, 252, 135, 495, 126]. commutatives [126]. commutativity [917, 1181]. commutator [542, 803, 588]. Commutators [737, 813]. commute [910]. commuting [341]. Compact [697, 266, 554, 195, 574, 48, 1003, 300, 421, 190, 370, 586, 975, 1145, 593, 109, 542, 821, 105, 758, 1166, 1125, 1051, 1110, 298, 617, 966]. Compactification [1076, 653, 1092, 966].

Compactifications

[568, 1066, 321, 827, 957, 617].

compactifying [584]. Compactness [416, 22, 1056, 953, 820, 847, 956, 867]. comparison [1060, 561, 759, 937]. Compatible [928, 845, 1230, 570, 1022]. complement [176]. complementary [310, 325]. complemented [104]. complements [1192, 1213]. completability [1197]. Complete [773, 464, 1046, 952, 1073, 616, 379]. completely [42, 1100, 351, 1072]. completeness [601, 953, 847, 140, 608]. completion [779]. completions [592]. Complex [1028, 485, 1071, 940, 1281, 400, 731, 1089, 668, 822, 227, 987]. complexes [529, 33, 508]. components [1097, 1269]. composed [159]. composés [159]. composite [1068]. composition [647, 349, 643]. Compositions [32]. concept [101]. concerning [732, 2, 415, 91, 73]. Concordance [299]. Condition [300, 917, 854, 88, 1208, 1158, 1041, 213, 509]. conditions [162, 1112, 1256, 1053, 637, 225, 1293, 619, 4, 660, 522, 258, 781]. Cone [1078, 469, 286, 956]. cones [1078]. conference [1116]. configurations [413]. confined [810]. Confluent [1097, 1090, 814]. conformal [1184]. Conformally [520, 341, 973]. Congruence [168, 1002, 502]. congruence-distributive [1002]. Congruence-preserving [168]. conics [652]. Conjecture [82, 1223, 566, 1109, 950, 73, 356, 403]. conjugacy [1267]. Conjugate [267, 699, 90, 110, 1099, 937, 1029]. conjugates [1262, 786]. connected [595, 1032, 400, 78, 922, 369, 679, 1076]. connectedly [78]. Connectivity [231]. Connexions [763]. consequence [415]. consequences [2]. constants [533, 1001]. constraint [1182]. constraints [1009]. Constructing [1185]. Construction [552, 709, 809, 344, 129, 888]. Constructions [1167, 771]. constructive [645]. contact [133]. contain [705, 300]. containing [94]. continua [475, 1219, 814]. continuable [1005, 1247]. Continuity [999, 364, 334, 1134]. Continuous [857, 872, 1120, 688, 875, 1282, 686, 952, 318, 31, 369, 206, 672, 1190, 324, 322, 115, 1134, 307]. continuum [1084]. contractible [143]. contribution [1067]. control [703]. Convergence [777, 187, 1142, 569, 1087, 1293, 734, 701]. converging [304]. Convex [579, 18, 838, 819, 818, 724, 923, 256, 1058, 59, 1018, 448, 1125]. convexes [1018]. convexity [498]. convexoid [1053]. coordinate [496]. coordinates [405]. copies [1151]. coprime [846]. copurity [907]. correcting [228]. Correction [1312, 69]. corresponding [1165]. Corrigenda [1310]. Corrigendum [556, 759, 259, 429, 1311, 261, 513, 415, 438, 1145, 1164, 500, 26, 160, 802]. Cosine [1129]. Cotorsion [750]. countability [1265]. countable [820, 392, 96, 970, 1006, 141, 793]. Countably [1110, 466]. Counterexample [356, 831]. Counterexamples [605, 621]. Counting [6, 274, 1124, 1224]. cover [769]. cover-avoidance [769]. Covering [435, 882, 1086, 974, 637]. Coverings [1168, 156]. covers [501, 1196, 1128, 546]. CPI [1030]. CPI-extensions [1030]. Creation [440]. Critères [1226]. Criteria [943, 742, 618, 909, 1226, 425, 288, 565, 1099, 218, 413, 661]. criterion [1115, 598]. cross [1188]. Crowns [650]. cubes [934]. cubic [1249]. Curvature [503, 1178, 341, 59, 590]. curve [459, 1186]. curved [1141]. Curves [961, 338, 398, 1090]. cusps [250]. cycles [6]. Cyclic [1033, 339, 922, 998, 563, 1209, 302, 803]. cyclically [595].

D. [82]. d.r.e [721]. dans [530, 812].

Darboux [885]. Dauns [768]. Davies [1127]. decay [136, 429]. decision [465]. decomposability [758]. decomposable [1309, 784, 660]. Decomposition [1052, 818, 923, 986, 1133, 714, 388]. decompositions [1063, 712, 740, 229, 513, 600, 214, 717]. decreasing [718, 1096]. Dedekind [484, 314, 481]. defective [1132]. Deficiencies [693]. defined [986, 1097]. Defining [1301, 389, 217]. definite [798, 749, 105, 356, 325]. definition [194]. definitions [1267]. degenerate [128]. degree [633, 1179, 425, 233, 624, 189]. Degrees [377, 157, 418, 696, 1059, 13, 67]. delay [509]. delays [1165]. deleted [143]. deloop [760]. demi [409]. demi-treillis [409]. d’endomorphismes [237]. dendroids [1212]. Denjoy [31]. dense [235, 321]. density [895, 1222]. denumerable [1004]. dependent [329]. dérivables [530]. Derivations [1198, 843, 297, 1122, 387, 548, 959]. derivative [897]. derivatives [406, 1035, 550, 45]. Derived [526, 70, 696]. descending [682]. Descendingly [346]. description [270, 925]. Designs [849, 70, 11, 106, 671, 378, 1, 26, 920]. destabilization [730]. determinants [1152]. Determination [599]. determined [468, 1194, 1135]. Determining [602, 876]. diagonable [983]. Diagonal [849, 238]. dichromate [1049]. diffeomorphism [212]. difference [144, 671, 1283]. difference-differential [144]. differentiable [1247, 530]. Differential [1112, 909, 39, 1165, 561, 759, 1005, 144, 703, 854, 166, 1299, 740, 288, 565, 811, 929, 218, 902, 1140, 695, 722, 755, 932, 1258, 534, 276, 450, 408, 509, 996, 410, 177]. differentials [407, 949, 1133, 400, 354, 175]. differentielle [39]. Dihedral [864]. Dilworth [1023]. Dimension [879, 1113, 541, 1109, 77, 1043, 1146, 417, 427, 1220]. dimensional [424, 1185, 711, 515, 318, 929, 726, 728, 1084, 1180, 143, 965, 617]. dimensions [55, 164]. d’interpolation [1149]. Diophantine [948]. Dirac [342]. Direchlet [100]. Direchlet-finite [100]. Direct [501, 563, 1183, 839, 64, 1140, 518, 214, 276, 848, 404]. directed [417, 579, 720]. directes [518]. direction [256]. directions [1051]. Dirichlet [264, 1235, 419, 854, 1222]. Disconjugacy [218]. disconjugate [177]. disconnected [1284]. Discrete [1151, 1221, 195, 1047, 881, 764, 1277, 901]. discreteness [1208]. discrets [764]. disintegrations [1269]. disjoint [709]. disk [530]. dismantlable [650]. Dispersed [1295]. disque [530]. dissectible [536]. distance [1241]. distinct [147]. distribution [894, 93, 523, 1038]. Distributional [393, 728]. distributions [336]. Distributive [57, 244, 1002, 104]. Divisibility [848, 1175, 586, 579]. divisible [167, 556]. division [912, 1101, 333]. divisor [661]. divisors [343]. does [760]. domain [363, 216]. Domains [375, 867, 859, 988, 882, 1030, 1303, 1046, 839, 862, 389, 856, 1162, 807, 314, 1095, 540, 481, 872]. Double [546, 1196, 765, 651]. doubly [1012, 587]. down [547, 226, 691]. dual [386, 589, 305, 651, 309, 506, 243, 512, 452]. dualities [1250]. duality [980, 788, 81, 116, 958]. due [62, 888]. d’un [1211, 1240, 767, 469, 409]. d’une [812]. dyadic [1047]. dynamical [61].

Easton [607]. Edge [455]. Edgeconvex [782]. Ehrenfest [1252]. Eigenfunctions [79]. eigenvalues [325]. Einstein [822, 1178]. Elations [106]. element [922]. élémentaire [767]. elementary [87, 136, 429, 767, 863, 1041, 661, 608]. elements [387, 567, 11, 43, 438, 126, 990, 674, 271, 126]. Elimination [946]. elliptic [1060, 1243, 684, 807]. embedded [889, 398, 66, 1274]. Embedding [475, 1177, 1212, 96, 995, 657, 1199, 358, 1125]. embeddings [992, 263, 1185]. end [196]. endomorphism [1176, 320, 154]. endomorphisms [1079, 237, 823, 674]. energy [622]. engendrées [126]. Enlargement [1056]. entières [335]. Entire [766, 550, 713, 914, 1082, 3, 160]. Entropy [810]. enumerable [796]. Enumerating [536, 399]. enumeration [1252]. envelope [409]. enveloping [585]. epi [681]. epi-reflective [681]. equalities [1039]. Equality [784, 340, 720]. equation [39, 538, 1005, 1247, 408, 509, 575, 39]. Equational [104, 1002, 281, 430, 245]. Equationally [421]. equations [618, 1060, 1165, 561, 759, 86, 259, 662, 703, 690, 740, 288, 565, 123, 218, 902, 722, 755, 932, 1258, 534, 807, 165, 614, 177]. equicontinuous [559, 558, 191]. Equidistant [290, 1203]. Equidistribution [812, 812]. equimeasurability [1039]. equimeasurable [53, 897]. equivalence [851, 762, 1068, 133, 969, 242, 133]. equivalences [81]. equivalent [973, 1109, 943, 1157]. Ergodic [1187, 396, 790, 644, 1246, 1120, 837, 927]. ergodicity [569]. Erratum [1247, 1146, 977]. error [228]. error-correcting [228]. espace [812]. espaces [739]. essential [138, 657, 482, 315]. essentially [1143]. estimates [58, 136, 429, 639]. Euclidean [532, 520, 731, 184, 98, 956]. Euler [33, 725, 852]. Eulerian [748]. eutactic [1055]. Even [1050, 1128, 218, 509]. Every [966, 464, 1308, 1310, 256, 544, 451, 743, 1274]. everywhere [1142]. evolution [86, 259, 165]. exact [200, 185]. exactly [552]. examined [374]. Examples [1007, 1292, 962]. exceptional [685]. Existence [216, 86, 259, 613, 1005, 1247, 147, 400, 413, 619, 1, 26, 531, 920]. exotic [212]. expanding [153]. expansions [386]. explicit [778, 714]. exponent [183, 1104]. exponential [915, 1236, 944, 852]. exponential-type [944]. exponents [1155]. expressions [854, 929]. extendability [510]. Extended [306, 279, 82]. Extending [271, 1102, 115]. Extension [383, 625, 1002, 385, 328, 188, 1245, 207, 1095, 69, 1190]. Extensions [285, 41, 1081, 114, 359, 40, 1030, 90, 864, 656, 137, 1294, 935, 1085, 574, 1275, 1179, 80, 895, 34, 1064, 801, 462, 292, 174, 691, 607, 135, 612, 315, 1120]. Extensive [960, 715]. extent [323]. exterior [1198]. extrapolation [1107]. Extrema [606]. extremal [442, 469]. extrémales [469]. Extreme [251, 1200].

Factor [836, 1308, 1310]. Factoring [1195]. Factorization [1238, 1031, 1295, 361]. factors [1171, 591, 1186, 1000, 219, 729, 402]. faibles [886]. fails [1058]. failure [1071]. Faithful [816]. Families [1173, 445, 110, 649, 389, 1265, 735, 558, 508, 298]. Familles [298]. family [594, 570]. fences [650]. Fermat [1186]. fiber [762, 268]. fibrations [1045, 545]. fibre [1068, 1120]. field [507, 864, 935, 1085, 971, 652, 801, 188, 174, 1245, 1229, 1279]. fields [40, 1294, 1263, 895, 354, 296, 1078, 616, 1015, 1276]. Filter [686, 1066]. filtrations [620]. final [132]. finitary [1111]. Finite [395, 155, 591, 1010, 1044, 832, 929, 233, 127, 1079, 8, 724, 507, 1288, 15, 1280, 147, 103, 275, 267, 705, 516, 517, 1206, 602, 623, 1256, 1169, 72, 971, 787, 253, 254, 373, 282, 449, 389, 384, 652, 1265, 257, 726, 153, 100, 606, 1298, 313, 857, 184, 331, 663, 1233, 735, 294, 1224, 1297, 207, 815, 243, 512, 583, 7, 94, 208, 631]. finite-order [1280]. Finite-to-one [155]. Finitely [511, 1143, 980, 468, 1308, 1310, 371, 24, 602, 477, 428, 1250, 559, 1135, 816]. finiteness [1115, 792, 495]. Finsler [122]. First [223, 1006]. five [711]. five-dimensional [711]. Fixed [109, 1147, 918, 933, 1241, 1084, 1292, 196]. fixed-point [1292]. flat [341, 1185, 520, 108]. flows [380]. fluids [522]. fonctionnelles [469]. fonctions [1142, 530, 1018]. forest [747]. form [818]. Formal [1093]. formally [1015]. formation [483]. forms [1055, 40, 1047, 555, 711, 1294, 419, 971, 749, 741, 648]. formula [1083, 472, 337]. Formulas [1014, 777, 49]. formule [337]. four [1088, 1109]. Fourier [931, 304, 1032, 574, 121, 71, 821, 14, 892, 1304, 714]. fourth [902, 695]. Fox [500]. fractional [1035, 842, 91, 394, 247]. Frattini [8, 876]. Frechet [450, 1193, 794, 982, 1134]. Free [916, 712, 195, 707, 1308, 1310, 1136, 501, 1206, 787, 1010, 523, 249, 376, 354, 51, 1191, 1201, 729, 608, 262]. French [1211, 1149, 39, 886, 1142, 335, 237, 1240, 530, 767, 469, 444, 1253, 1018, 914, 812, 1226, 126, 133, 518, 764, 643, 159, 409, 891, 805, 298, 337]. Friedberg [374]. Frobenius [494, 1174, 636, 506, 948]. frontière [530]. Fuchsian [487, 870, 319, 898]. Full [140, 1126]. Full-completeness [140]. fully [167, 556]. function [713, 87, 364, 118, 231, 1121, 354, 571, 1276, 450, 610, 211, 892, 3, 160]. functional [270, 924, 178]. Functionals [1131, 1171, 469, 906, 982, 514]. Functions [121, 445, 688, 435, 699, 605, 903, 719, 18, 284, 882, 1011, 58, 875, 915, 1236, 1142, 436, 498, 530, 1282, 1014, 986, 59, 1129, 42, 1018, 172, 261, 609, 802, 1200, 411, 93, 794, 477, 1082, 1222, 484, 318, 798, 1073, 1052, 334, 99, 100, 766, 1300, 658, 581, 702, 606, 145, 293, 1150, 1138, 503, 799, 1048, 239, 311, 175, 414, 105, 833, 1190, 324, 1216, 550, 79, 941, 1039, 356, 227, 146, 905, 1134, 865, 872, 701, 45, 307]. functor [743]. functors [645, 265, 596, 664, 185]. Fundamental [255, 887, 532, 1299, 749].

Gabriel [1146, 1043]. Gaining [963].

Galois [34, 315]. gas [810]. Gasper [1098]. Gauss [472]. Gaussian [1227, 590]. GCD [540]. GCD-domains [540]. GCR [362].

Gelfand [61]. general [995, 897, 809, 807, 1048, 129, 791, 1160, 388, 1065]. Généralisation [337]. generality [209]. generalization [706, 633, 132, 869, 299, 46, 332, 342, 112, 1098, 545, 1261, 888, 504, 280, 337]. Generalizations [789, 456]. Generalized [198, 410, 987, 8, 1173, 871, 1033, 1260, 1129, 1053, 121, 123, 1099, 29, 30, 240, 594, 1202, 860, 1204, 868, 725, 82, 746, 404, 852]. generalizing [853]. generated [1106, 1270, 980, 707, 1308, 1310, 511, 1280, 647, 24, 78, 126, 659, 428, 483, 151, 816]. generating [1052]. generation [595, 726, 184, 331]. generator [7]. Generators [234, 461, 582, 469, 778, 348]. Génératrices [469]. Generic [1136, 812, 124]. générique [812]. Gentle [125]. genus [873, 626]. geodesic [590]. Geometric [150, 241, 21]. geometries [557, 921, 1260, 290]. Geometry [1089, 684, 184, 276, 25]. germs [468]. Gibbs [725]. girth [954, 1306, 151]. given [602, 19, 570]. global [1225, 895, 1015, 427]. Goethals [744]. Going [226, 691, 547]. Goldschmidt [846]. good [553, 1070]. Gottlieb [788]. Graded [1248, 985]. grading [874, 1312]. graph [1049, 1031, 936, 1000, 176, 403, 720, 139]. graphes [891]. Graphical [365, 366]. Graphs [463, 1249, 434, 6, 954, 595, 981, 1126, 1023, 873, 70, 418, 455, 1210, 748, 1100, 128, 471, 1306, 163, 891, 69, 151, 890, 970, 403, 675, 274, 626]. Green’s [674, 1025]. Grothendieck [580, 158]. Grothendiek [785]. Group [89, 384, 8, 611, 580, 564, 1240, 35, 638, 1080, 928, 1022, 767, 602, 995, 940, 1179, 870, 1256, 72, 431, 34, 1010, 300, 1199, 190, 449, 152, 348, 696, 779, 879, 736, 1075, 911, 913, 579, 1159, 874, 1312, 521, 327, 470, 757, 105, 1113, 168, 204, 472, 205, 127, 698, 1285, 113, 612, 1279, 745, 543, 139, 631]. groupe [767]. groupes [964, 1240, 492, 764, 409]. Groups [1308, 1310, 1192, 418, 494, 1088, 162, 676, 234, 461, 1139, 998, 1040, 27, 536, 138, 563, 1013, 1143, 487, 980, 707, 511, 532, 371, 921, 267, 395, 132, 705, 24, 192, 995, 991, 1260, 1032, 1069, 1168, 623, 1169, 591, 869, 372, 253, 254, 373, 282, 220, 1301, 102, 370, 157, 950, 1044, 60, 1104, 537, 863, 88, 64, 123, 119, 210, 405, 249, 217, 551, 55, 164, 426, 1266, 726, 1151, 235, 232, 585, 233, 465, 428, 184, 669, 887, 149, 174, 1180, 1245, 792, 877, 96, 97, 1004, 913, 358, 319, 451, 1224, 1020, 898]. groups [365, 366, 4, 409, 846, 1041, 821, 320, 1166, 207, 352, 85, 191, 815, 158, 243, 512, 884, 677, 404, 708, 1160, 502, 583, 382, 519, 769, 803, 7, 816, 222, 608, 742]. Growth [136, 429, 660, 1053, 1082, 1204].

H [500]. Hadamard [1260, 943, 378, 744]. Hahn [1098]. Hall [578]. Halmos [737]. Hamiltonian [1249]. handlebody [965]. Handles [730]. Hankel [699, 1099, 728, 170]. Hardy [656, 1127, 251, 355, 219]. Harmonic [1141, 264, 894, 5, 99, 100, 764]. harmonique [764]. Hausdorff [1100, 883, 1182, 922, 774, 206, 533, 966]. Hausdorff-closed [1100]. hautement [159]. having [1040, 1023, 552, 800, 1164, 874, 1312, 793]. heat [575]. height [1079]. Heilbronn [91]. Heine [241]. hereditarily [919, 1076]. hereditary [641, 1157, 488, 640, 772, 829]. Hermite [661]. Hermitian [971]. Hewitt [74, 388]. hierarchy [941]. High [387, 563, 1059]. Higher [1122, 145, 293, 1170, 297, 515, 959, 1123]. Highly [560, 159]. highly-composed [159]. Higman [578]. Hilbert [737, 1089, 161, 420, 987]. Hjelmslev [1266]. HNN [551]. Hofmann [768]. Hölder [924]. holomorphe [812]. Holomorphic [906, 731, 812, 477, 1138, 1277, 872]. Homeomorphic [199]. Homeomorphism [677, 1109]. Homeomorphisms [765, 1271, 299, 1239, 965]. homocyclic [591]. homogeneous [928, 1022, 425, 961, 560, 946]. Homology [1144, 143, 439, 223, 1160, 717]. homomophism [326]. homomorphic [1206]. homomorphism [1240, 700, 1266]. homomorphisme [1240]. Homomorphisms [632, 1016, 1031, 1096, 73, 694, 271, 359]. homotopical [775, 776]. Homotopy [44, 1068, 958, 371, 278, 974, 788, 81, 529, 850, 303, 1008]. Hopfian [1271]. Hughes [198, 860, 82]. hull [611, 678, 628]. hulls [838, 681, 262]. Hurwitz [337]. hyperbolic [216, 731]. hyperellipticity [250]. hypergraphs [1169]. hyperinvariant [925]. hyperplane [19, 124]. hyperplanes [110]. hyperspace [286]. hyperspaces [820, 1212, 1219]. hypersurfaces [448].

Ideal [137, 535, 330, 620, 1207]. Ideals [297, 959, 571, 146, 893, 1270, 244, 1194, 340, 479, 1096, 629, 521, 540, 271, 1195, 307]. idempotent [126, 85, 743]. Idempotents [431, 126]. identification [163]. identities [1232, 748, 178, 746]. identity [1278, 983]. if [300]. II [240, 828, 692, 1236, 291, 1136, 436, 974, 238, 548, 68, 1230, 253, 254, 751, 565, 1108, 835, 332, 30, 1075, 378, 1258, 365, 366, 1017, 1201, 512, 931, 1273, 829, 502, 460, 776, 948, 84]. III [180, 1167, 145, 143, 25, 474, 274]. images [806, 1046, 1206, 672]. imbedding [202]. immersed [673, 822]. Immersions [263, 749, 878, 390, 546]. implicational [601]. Implications [4]. implies [1135, 847]. imply [1039]. Imprimitive [940]. Improved [555, 551]. improving [454]. Incidence [111, 361]. Inclusion [446, 219]. incomplete [346, 11, 1, 26]. Indecomposable [277, 1287, 1143]. indéfiniment [530]. Indefinite [122]. independence [915, 1236]. Index [1088, 1183, 602, 72, 83, 84]. Indicator [967, 1220, 797]. indistinguishability [1268]. Induced [81, 710, 266]. inducing [815]. induction [840]. inductive [1217]. inequalities [699, 713, 357, 53, 247]. Inequality [1009, 897, 1307, 1237, 1127, 328, 504]. Infinite [932, 1258, 478, 407, 499, 1185, 625, 1007, 1126, 796, 1074, 224, 755, 1283]. infinite-difference [1283]. infinite-dimensional [1185]. infinitely [530]. inflation [283]. influence [8]. Initial [13, 67]. Initially [817]. Injective [262, 1013, 1054, 1311, 479, 1027, 320, 130, 640]. injectivité [237]. Injectivity [281, 1002, 237]. injectors [483]. inner [651, 815]. inseparability [1263]. inseparable [935, 1085]. insulator [1240]. integer [335]. integers [1279]. integrable [561, 759]. Integral [53, 411, 1073, 462, 336, 913, 49, 40, 859, 1030, 528, 839, 209, 1154, 897, 389, 31, 663, 534, 214, 20, 1095, 1039, 21]. integrals [1124, 625, 272, 37, 394, 21, 247]. integration [224]. interior [703]. intermediate [534]. internal [193]. interpolation [1105, 1071, 505, 1028, 852, 1001, 1149, 794]. interpolatory [553, 1070]. interpretation [905]. Intersections [819, 307]. intertwining [1309, 1134]. interval [36]. intrinsic [804]. invariance [87, 868, 1048]. Invariant [1094, 235, 585, 1118, 1213, 641, 469, 1017, 63, 648, 631]. invariantes [469]. invariants [1085, 28, 1086, 1144, 887, 500, 775, 776]. Inverse [1271, 285, 1033, 1225, 670, 1259, 944, 628]. inverses [704]. Invertibility [148]. invertible [476]. involution [795, 1181, 604, 751, 632, 910, 1094, 317, 544, 1118, 63, 746, 271]. involutions [1249, 103, 705]. Involutory [294]. irrationals [453]. Irreducible [991, 634, 940, 223, 300, 994, 55, 164, 1298, 149, 976]. Isometric [938]. Isometries [856, 401, 422, 694]. Isometry [828, 184]. Isomorphic [253, 254, 583, 395, 866, 1104, 577]. Isomorphism [985, 677]. isomorphisms [426, 526, 168, 1272, 502, 698]. isotopic [683]. isotopies [730]. isotropic [711]. Iterated [655, 607]. iteration [1007]. iterations [396]. IV [293, 493].

J [556, 759, 737, 1022, 802, 1070, 513, 1145, 1146, 959, 1019, 977, 500, 1164]. Jacobi [180, 23, 68, 414]. Jacobian [1186]. James [809]. Janko [207]. jets [485]. join [804, 1057]. joins [981]. Jordan [926, 1063, 1218, 271, 687]. Julia [942].

K. [1019]. Kaehler [668, 822, 390]. Kähler [872]. Kaplansky [768]. Kasteleyn [471]. kernel [1098, 664, 743]. kernels [663, 56]. kind [742]. Kinetic [522, 810]. Klein [225]. knot [792]. knots [851, 515, 1192, 28, 841, 977, 500]. König [853]. Korovkin [663]. Kronecker [564, 981]. Krull [1146, 1043, 343]. Kuan [959]. Kuiper [825].

Labeled [1205]. Lagrange [453]. Laguerre [1232, 386]. Lamperti [1187]. Large [675, 954, 881, 11, 1082, 229, 513, 709]. Latin [934]. lattice [1270, 187, 244, 980, 611, 986, 372, 712, 43, 438, 700, 925, 483, 245, 1191, 1201, 458]. lattice-ordered [980, 611, 372]. lattice-ordered-rings [458]. lattice-theoretic [925]. lattices [1079, 1270, 57, 236, 291, 1115, 1288, 368, 686, 1206, 1121, 1256, 430, 787, 808, 220, 75, 650, 752, 51, 104, 1150, 420, 1132, 655, 891, 890, 150]. Laufer [173]. Laurent [1275]. laws [411, 1008]. LCA [1139]. LCM [363, 988]. Lebesgue [986, 1001]. Lee [1150]. left [185]. lemma [737, 885, 280]. lemmas [1267]. length [623, 696, 612]. L’enveloppe [409]. leur [237]. leurs [126].

Leviflat [1281]. Levitzki [1290]. Lie [387, 415, 1211, 1286, 843, 1148, 844, 921, 985, 1032, 300, 2, 726, 367, 1298, 1075, 98, 470, 179, 316, 1231, 107]. Lie-like [387]. Lifting [1062, 1214, 1272, 412]. liftings [252]. like [475, 58, 387, 367, 639]. Limit [909, 288, 565, 411, 473, 861]. limits [670, 1259, 655]. Lindelöf [732, 919, 1302]. line [705, 277]. linéaires [469]. Linear [87, 951, 868, 1048, 757, 227, 745, 1171, 1236, 561, 759, 498, 308, 86, 662, 256, 995, 690, 469, 48, 623, 598, 740, 161, 348, 55, 726, 695, 233, 925, 722, 755, 56, 441, 433, 1160, 996, 1134, 165, 855, 948, 326, 222, 907]. Linearization [180, 23, 68]. linearly [831]. Lines [110]. link [621]. linkage [1086]. links [851, 28, 877, 841, 977, 500]. Liouville [145, 293]. Lipschitz [1035]. Lipschitzian [649]. L’isolateur [1240]. Littlewood [656]. Local [1305, 1064, 292, 1138, 896, 65, 40, 644, 724, 912, 981, 43, 438, 741, 188, 174, 1015, 294, 1095, 867, 133]. local-global [1015]. locale [133]. Localization [71, 629, 880, 1244, 1273, 716]. Localizations [899]. Locally [448, 436, 1185, 1058, 282, 190, 370, 635, 88, 975, 1145, 153, 922, 105, 1166, 1125, 108, 1076, 617, 966]. loci [290]. Lodato [570]. logarithmic [1155, 489]. logic [830]. longest [720]. loop [1074]. loops [1063, 939, 1244, 1273, 508]. Lower [1243, 453, 729]. Lusternik [134]. Lyndon [869, 376].

M. [625]. Mackey [221, 589]. MacRae [610]. magic [934]. Mahler [1019]. majorant [1034, 1166]. Majorization [436]. Majorization-subordination [436]. malnormal [249]. manifold [973]. manifolds [828, 997, 1185, 223, 1281, 1141, 229, 513, 668, 673, 822, 590, 355, 303, 885]. many [371, 13]. many-one [13]. map [760, 1264, 616]. mapping [882, 633, 81, 685, 83, 84]. mappings [468, 947, 1090, 155, 48, 1097, 731, 1141, 938, 1184, 814, 1078, 369, 441, 150, 141, 987]. Maps [108, 41, 338, 308, 80, 681, 339, 385, 1064, 392, 1289, 1153, 189]. Martin [1110]. martingales [473]. mass [440]. Matchings [970, 675]. Math [556, 759, 737, 1022, 802, 1070, 513, 1145, 1146, 959, 1164, 977, 500]. Mathieu [911]. Matrices [1012, 1018, 87, 1173, 707, 457, 238, 1260, 587, 748, 1127, 378, 1116, 294, 1152, 868, 1048, 239, 744, 203, 478]. matrix [538, 989, 943, 1048, 175, 450, 1218, 325]. Maximal [347, 27, 192, 1130, 761, 467, 692, 301, 88, 188, 1207, 207, 927]. Maximality [118]. maximally [616]. maximum [93, 99, 100, 3, 160, 626]. Mean [1251, 625, 121, 766]. means [1097, 29, 30, 240, 235, 336, 701]. measurable [1142]. Measure [1092, 476, 622, 1236, 593, 771, 351, 94]. measures [264, 334, 670, 559, 558, 85, 1304]. member [818]. Mercerian [258]. meromorphic [93, 702, 799, 833, 1190]. mesurables [1142]. metabelian [816]. metacompact [531, 345]. Metacyclic [28, 500]. metanilpotent [102]. metastable [212, 546]. method [206, 258]. methods [924, 1007, 489]. Metric [881, 806, 1034, 834, 397, 857, 369, 321]. metrics [1034, 973, 525]. Metrizability [480, 993, 260]. metrizable [48, 497, 531]. Metrization [780, 1067]. Meyer [853]. middle [1287]. Minimal [1156, 119, 841, 977, 1170, 230, 896, 904]. minimax [720]. minimum [767, 1179, 767]. Minkowski [1180]. Minkowskian [290]. minus [1103]. Mochizuki [950]. Model [608]. Model-completeness [608]. models [1170]. Modular [378, 988, 1256]. Modularity [801]. module [464, 949, 1133, 651, 302, 16]. Modules [772, 829, 167, 556, 237, 182, 215, 34, 537, 518, 1250, 930, 1272, 488, 640, 1209, 315, 262]. modulo [1038, 1285]. modulus [93]. monads [686]. Monocoreflective [1065]. monoids [418]. monomially [1270]. monomials [1106]. monothetic [234, 461]. monotone [338]. monotonic [42]. monotonicity [1073, 145, 293]. Montel [17]. Moore [1223, 369, 260, 672, 962, 531, 883]. Morse [689, 885]. most [1287]. motion [684, 77]. Moufang [76]. Muchnik [374]. multifunction [774]. multifunctions [196, 322]. Multilinear [239, 756]. multiple [21]. multiples [864]. Multiplication [893, 573, 482, 1074]. Multiplications [371, 762]. Multiplicative [982, 224, 1245, 27, 508]. multiplier [768]. multipliers [304]. Multivalent [702]. my [461].

Nagata [988]. Nakayama [616]. natural [44]. near [136, 429]. nearfields [111]. nearly [191]. necessary [509]. negatively [1141]. Nehari [1014]. Neighbor [1266]. neighborhoods [1185, 673]. Neighborly [997]. neighbourhoods [398, 54]. nest [582]. nested [1197]. nets [727, 186, 978, 1029]. Neumann [526, 214, 577]. Nijenhuis [490]. nil [52, 142]. nilalgebras [642]. nilpotency [142]. nilpotent [1286, 1032, 913, 544, 207, 169, 803]. nine [113]. no [705, 1022, 300, 1145, 1146, 1164, 977]. nodes [1172]. Noether [438, 1270, 43, 75]. noetherian [330, 431, 629, 880, 1158, 312, 640, 772, 829]. nombres [159]. Non [524, 707, 1104, 413, 441, 497, 1, 26, 577, 947, 265, 371, 86, 1047, 136, 429, 111, 505, 1058, 576, 635, 811, 1228, 832, 919, 938, 184, 56, 774, 1132, 880, 1269, 365, 366, 242, 165, 883]. non- [1228, 832, 1269]. Non-abelian [524, 365, 366]. non-central [265]. non-commutative [938, 880]. non-defective [1132]. non-dyadic [1047]. non-elementary [136, 429]. non-Euclidean [184]. Non-existence [413, 1, 26]. Non-free [707]. non-Hausdorff [774]. Non-isomorphic [1104, 577]. non-Lindelöf [919]. Non-linear [441, 86, 56, 165]. non-locally [1058]. non-locally-solvable [635]. Non-metrizable [497]. non-normal [242, 883]. non-planar [111]. non-poised [505]. non-selfadjoint [811]. non-standard [947, 576]. non-vanishing [371]. nonalternate [538]. noncommutative [821]. nonconvexity [724]. nonelliptic [1299]. Nonlinear [902, 259, 613, 509, 514]. nonoscillatory [690]. nonselfadjoint [218]. nonspherical [545]. nonstandard [838]. Nörlund [71, 296, 14]. Norm [699, 1096, 1087, 701, 313, 931, 247]. Normal [51, 445, 284, 1183, 270, 457, 1223, 93, 1309, 376, 525, 20, 758, 962, 1128, 531, 242, 883, 478, 1284]. normalisateur [1211]. Normality [863, 1123, 117, 477]. normalizer [1211]. Normed [256, 391, 316, 499]. norms [255]. Noshiro [456]. note [1060, 733, 310, 1223, 566, 854, 1188, 1101, 993, 584, 333, 210, 1184, 385, 124, 465, 669, 741, 558, 404, 976, 648]. Notes [1095]. Nowhere [321]. nuclear [1193]. null [1113]. Number [1019, 1243, 527, 1023, 148, 11, 1114, 348, 454, 800, 1164, 936, 1262, 747, 208, 631]. numbers [1254, 864, 523, 159, 92]. Numerical [775, 776, 594, 1123, 101, 211].

obeying [1008]. objects [278]. Obstructions [252]. odd [581, 1306, 1285]. off [302]. ominoes [454]. once [300]. One [704, 285, 487, 155, 869, 425, 217, 354, 133, 13, 552, 824, 292, 1021]. one-relator [487, 869]. One-sided [704]. ones [1103]. only [384]. onto [882]. open [806, 362, 155, 80, 906, 385, 1064, 691, 679, 575]. operad [1074]. opérateurs [643]. Operations [580, 32, 669, 781]. operator [999, 871, 662, 1129, 1278, 1175, 1299, 635, 354, 853, 1214, 1041, 213, 211, 79]. operator-valued [79]. Operators [574, 824, 542, 1243, 554, 909, 697, 908, 1183, 1063, 1112, 647, 737, 349, 1225, 1003, 1188, 197, 1053, 667, 161, 811, 1309, 61, 1187, 1176, 933, 1140, 695, 663, 944, 1107, 643, 534, 1017, 757, 20, 660, 758, 510, 9, 116, 410, 242, 1134]. orbit [195, 319]. orbital [1124]. order [753, 1170, 1242, 949, 561, 759, 1005, 1247, 1035, 864, 387, 1280, 1263, 854, 1133, 740, 384, 218, 902, 695, 1116, 509, 113, 410, 793, 177]. orderable [1020]. Ordered [88, 244, 980, 611, 1155, 372, 918, 994, 1202, 358, 592, 352, 458]. Ordering [592, 900]. orderings [992, 1233, 1259]. orders [928, 1022, 845, 1230, 148, 158]. ordinal [941]. ordinary [722, 410]. orientation [163]. orientations [1049]. oriented [1289]. Orlicz [1151]. orthoautomorphisms [420]. Orthogonal [238, 849, 1301, 164, 745, 920]. orthogonalities [839]. orthogonality [514]. orthogonally [773]. ortholattices [916]. orthomodular [236, 291, 1115, 1288, 220]. Oscillation [618, 561, 759, 425, 807, 225, 902, 695, 663, 509, 996]. oscillatory [690]. osculatory [183, 424]. outer [548, 350, 588]. outer-commutator [588]. overlap [1293]. overrings [1030, 896].

P [737]. packing [183, 424]. packings [228]. pairs [1061, 598]. paper [461, 1019]. parabolic [921, 359]. paracompact [975, 1145, 466]. paracompactness [375, 392, 867]. Parallel [472]. parfaits [492]. parity [936]. part [630, 1264, 91]. Partial [186, 1260, 992, 703, 1130, 1299, 811, 408]. Partially [722, 244, 1155, 918, 994, 592, 1, 26]. partition [516, 455]. partitions [1254, 2, 415, 791]. partout [1142]. parts [842, 1108]. path [412, 720]. paths [6, 50]. Peano [406]. Pearcy [737, 211]. Pentagon [151]. Pentagon-generated [151]. Perfect [681, 141, 947, 95, 77, 403, 543]. Periodic [1257, 165, 225, 659, 766]. permanent [27, 566]. Permanents [587]. Permutation [903, 868, 163, 1203]. permutations [507]. Perron [37]. Perturbation [273, 1140, 440]. Perturbations [1291, 554, 1183, 125, 929, 969, 758]. Petersen [1127]. Pettis [1151]. phase [742]. phenomenon [725]. Picard [887, 870]. PID [314]. piecewise [862, 324]. PL [299]. PL-homeomorphisms [299]. Planar [752, 1191, 1201, 111, 1084, 639, 683, 403, 437, 460, 474, 493, 614]. plane [1303, 76, 400, 17, 168, 113, 94]. planes [103, 275, 705, 198, 835, 257, 860, 82]. plat [851, 841, 977]. Plessner [555]. Plethysm [865, 311]. Plücker [405]. pluriharmonic [1200]. plus [1103]. poids [1119]. Poincaré [1196]. Point [36, 909, 1147, 288, 565, 918, 1265, 933, 292, 1084, 1292, 176, 196]. point-arboricity [176]. point-finite [1265]. Point-transitive [36]. pointed [976]. Points [724, 806, 445, 1243, 727, 818, 136, 429, 19, 250, 937, 109, 251, 199]. pointwise [790]. poised [505]. polarities [257]. Pollaczek [1098]. Pólya [656]. polycyclic [623]. polydisk [1149]. polydisque [1149]. polygonal [1306, 94]. polygons [1033]. polyhedra [536, 979, 143]. Polymonial [350]. polynômes [886]. Polynomes [335]. Polynomial [1204, 521, 1270, 1232, 842, 823, 547, 914, 983, 748, 226, 746, 946]. polynomiale [914, 1119]. polynomials [180, 886, 894, 5, 553, 1070, 23, 68, 1175, 585, 1150, 414, 1098, 1117, 1227, 306, 946, 335]. Polytopes [38, 997, 831, 595, 5]. polyverbal [669]. posets [220]. Positive [308, 798, 969, 469, 1200, 933, 593, 1107, 105, 972, 356, 325]. positives [469]. possessing [56]. Possibility [834]. possible [206]. potentials [1204, 1025]. pour [1149, 39]. power [1093, 832, 853, 1297]. powers [203]. pre [761]. pre-primal [761]. Prefrattini [483, 769]. prescribed [800, 1164, 1025].

Presentations [382, 1159]. presented [602]. preserve [1241]. preserving [476, 1032, 1078, 757, 168, 694, 745]. presque [1142]. primal [761, 827]. primary [138, 1042, 132, 152, 43, 438, 450, 307]. Prime [479, 540, 244, 1030, 279, 330, 230, 343, 878, 841, 977, 640, 772, 829, 1285]. primes [541]. Primitive [567, 187, 285, 535, 736, 941, 793, 222]. primitivity [204]. principal [325]. principle [99, 100]. principles [1015]. probabilities [1037]. probability [21]. problem [998, 264, 1235, 444, 71, 61, 782, 1019, 171, 97, 948, 83, 84]. problème [444]. problems [903, 871, 216, 1062, 505, 273, 465, 613, 97, 504]. procedure [595, 454]. processes [1246, 440, 837]. Product [392, 180, 875, 23, 68, 350, 295, 249, 651, 47, 905]. Products [813, 76, 370, 665, 1170, 284, 1122, 564, 839, 981, 862, 873, 667, 64, 119, 51, 232, 336, 224, 74, 143, 1277, 815, 404, 803, 577, 687, 626, 345]. projectif [812]. projectifs [964]. projection [908, 289]. Projections [1303, 362, 1100]. Projective [1179, 303, 57, 103, 275, 1091, 1161, 812, 157, 257, 1050, 822, 277, 640]. prolongations [426, 763, 208]. proof [173, 927]. proofs [1059]. proper [441, 189]. Properties [381, 1211, 753, 117, 476, 999, 380, 398, 718, 442, 1219, 455, 78, 685, 1187, 1064, 145, 293, 292, 670, 898, 1218, 1234, 562, 508, 789, 769, 1137, 608]. property [908, 974, 1155, 895, 350, 289, 313, 1084, 877, 534, 309, 1166, 847, 286]. propriétés [1211]. proximities [497, 570]. proximity [1221]. Prüfer [1046, 572]. Pseudo [264, 622, 814, 104, 1239, 107]. pseudo-algebras [107]. pseudo-arc [1239]. pseudo-complemented [104]. Pseudo-confluent [814]. Pseudo-measure [622]. pseudocompactness [820]. Pseudodifferential [635]. pseudogroup [1211]. pseudogroupe [1211]. pseudometrics [115]. Pull [850, 958].

Pull-backs [850, 958]. Pure [563, 935, 827, 1013, 451, 1027, 518]. purely [935, 1027]. purifiable [152]. Purity [907, 680, 598]. purs [964, 518].

Quadratic [1294, 453, 40, 1047, 711, 400, 1078, 741, 648]. Quantization [1080]. Quasi [1013, 237, 200, 640, 201, 138, 964, 1097, 993, 477, 589, 636, 309, 320, 708, 827, 478, 237]. Quasi- [1013]. quasi-ambivalent [708]. quasi-components [1097]. quasi-dual [589, 309]. quasi-essential [138]. quasi-Frobenius [636]. Quasi-injective [640, 320]. Quasi-injectivité [237]. Quasi-injectivity [237]. quasi-metrizability [993]. quasi-normal [478]. quasi-normality [477]. quasi-primal [827]. quasi-projective [640]. quasi-purs-projectifs [964]. Quasi-splitting [200]. quasi-uniserial [201]. quasianalytic [912]. quasiconformal [882]. Quasicyclic [1077]. Quasigraphs [955]. quasigroups [1262, 939]. quasilinear [618]. quasivarieties [1072]. quaternion [1294, 1189]. Quelques [1211]. questions [1292]. quivers [1174]. Quotient [1193, 458, 347, 363, 1046, 573, 381]. Quotients [1259, 1270, 1010, 330, 1073, 98, 324, 646].

R [737, 82, 500]. R. [62, 809]. Racine [767]. radial [701]. Radical [1061, 166, 129, 1290, 1218]. Radicals [314, 467, 692, 269]. radius [498, 59, 503]. Ramanujan [159]. Ramification [174, 615, 616]. Ramsey [747]. Random [821, 587, 1172]. range [212, 1123]. Ranges [667, 594, 394]. Rank [538, 1040, 407, 132, 952, 824, 615, 3, 160, 92, 208, 1021]. rates [1156]. Ratio [1246, 861, 65, 473, 214]. Rational [328, 979, 834, 1276, 972]. rationally [464]. rays [1200, 942]. re [374]. re-examined [374]. Real [1105, 1189, 541, 1131, 914, 484, 389, 693, 826, 1015, 757, 972, 745]. realcompact [1302, 345]. realcompactifications [74]. realcompactness [193]. reals [391]. Rearrangement [357, 1217, 897]. rearrangements [718, 840, 53]. reciprocity [506]. Reconstruction [10]. Recursive [992, 377, 941]. Recursively [659, 796]. Redfield [756]. reducing [718]. reductive [697, 678, 179]. réel [914]. reexamined [803]. Refinement [781]. reflection [1224]. reflections [76]. reflective [681]. reflexifs [805]. reflexive [527, 1003, 353, 824, 805]. regions [532, 400]. Regular [5, 1169, 596, 1111, 673, 627, 839, 630, 70, 479, 418, 405, 674, 544, 351, 365, 366, 1072, 904, 135, 1153]. régularité [39]. Regularity [653, 39, 166, 375]. reguli [967]. related [117, 1097, 378, 720]. relation [603, 217, 1266, 33, 7]. Relations [1053, 527, 1301, 12, 119, 674, 351]. relationship [185]. relative [273, 664]. Relatively [1135, 379]. relator [487, 869]. Remark [461, 62, 173, 1269, 738]. remarkable [434]. Remarks [1281, 1019, 131, 490]. remote [199]. remplissage [50]. Repartition [159, 159]. replaceable [186, 978]. Representable [1250, 1270, 185]. Representation [161, 317, 48, 1169, 72, 630, 300, 164, 1075, 1092, 976]. Representations [971, 1175, 220, 851, 246, 266, 564, 1124, 921, 1080, 89, 1129, 1179, 300, 634, 157, 1073, 55, 51, 1298, 149, 328, 841, 977, 365, 366, 1277, 710, 815, 836, 816, 742]. representative [147]. Representing [952]. Residual [495, 792]. Residually [15]. Residuals [1057]. Residue [354]. resolution [811]. resolutions [1136]. resolvent [635]. respect [956]. restricted [300, 855]. Restricting [815]. restrictions [48]. Restrictive [318]. result [1060, 791]. results [688, 1011, 711, 1121, 72, 820, 637, 689, 607, 543]. retracts [181]. revisited [583]. Ricci [341]. Rice [721, 1036].

Richardson [653]. Riemann [906, 109, 319, 251, 1153, 337]. Riemannian [828, 99, 100, 590, 885]. Riesz [1022, 240, 953, 625, 928, 845, 1230, 71, 29, 30, 289, 600, 379, 701]. Right [363, 641, 352, 1177, 330, 629, 1020, 391]. right-orderable [1020]. Right-ordered [352]. Rigid [468, 995, 1167]. ring [893, 237, 1046, 638, 823, 34, 1199, 330, 482, 230, 736, 526, 271, 946, 729]. Rings [464, 910, 313, 544, 1209, 688, 893, 1248, 1270, 347, 926, 363, 917, 52, 297, 1122, 641, 1177, 773, 580, 627, 795, 1181, 1148, 1255, 15, 912, 1047, 35, 547, 754, 95, 968, 431, 350, 479, 166, 573, 572, 421, 584, 952, 604, 751, 423, 899, 384, 863, 832, 61, 1176, 46, 332, 1163, 959, 1151, 629, 632, 1094, 131, 887, 741, 226, 294, 913, 1026, 1118, 112, 636, 880, 1158, 930, 381, 521, 470, 704, 320, 1113, 785, 805, 135, 1218, 710, 427, 746, 646, 205, 312, 661, 495, 130, 640, 772, 829, 458, 683, 146, 307]. rings [543, 154]. Robinson [888]. root [767, 465]. rooted [437, 460, 474, 493, 614]. rotation [658, 581, 702, 503]. rotund [1051, 1261]. row [239].

Saks [539]. salesman [782]. sample [781]. Sandler [62]. satisfies [983]. satisfying [88]. Saturation [944, 360]. scalar [329]. scale [489]. schemes [127, 495]. Schnirelmann [134]. Schur [72, 328, 1229, 905]. Schwarz [1019]. screenability [260]. Screening [78]. second [561, 759, 1005, 1247, 854, 749, 410, 786, 742]. second-order [854]. secrétaires [444]. secretary [444]. section [124, 1234]. sections [1188, 19]. segments [13, 67]. Seidel [744]. Selberg [1083]. selection [593]. selections [322]. self [479, 542, 213, 130]. self-adjoint [213]. self-commutator [542]. self-injective [479, 130]. selfadjoint [811]. Semi [525, 956, 492, 749, 741, 369, 20, 326, 543, 664].

Semi-compactness [956]. semi-definite [749]. semi-linear [326]. semi-local [741]. semi-metric [369]. Semi-metrics [525]. semi-normal [20]. semi-parfaits [492]. semi-perfect [543]. Semi-simplicity [664]. semigroup [285, 42, 114, 73, 1297, 628]. Semigroups [9, 1271, 201, 1165, 137, 1129, 1077, 268, 318, 674, 245, 678, 1072]. semihandles [730]. semilattice [409, 678]. semilattices [430]. semilinear [1235]. semilocal [1046, 1163]. semiprime [926, 629, 1195]. Semiregular [930]. semirings [586]. semisignatures [1163]. semisimple [1075, 1215]. sense [343]. Separability [326, 801]. separable [908, 486, 794, 589, 919, 966]. separating [666]. separation [416, 890, 891, 891]. Sequence [1252, 200, 323, 496, 1038, 433]. sequences [1287, 1103, 949, 777, 255, 553, 1070, 1052, 659, 597, 1234]. Sequential [847, 953]. Sequentially [379]. Series [386, 304, 1093, 272, 71, 853, 620, 821, 1277, 101, 14, 734, 1279]. set [244, 364, 32, 147, 928, 1022, 986, 606, 417, 214, 94]. Sets [1279, 1149, 889, 819, 818, 724, 923, 797, 1142, 499, 777, 147, 881, 1155, 1161, 110, 1028, 666, 788, 1082, 721, 1264, 918, 685, 1228, 377, 1036, 1265, 77, 994, 1151, 593, 709, 936, 857, 1292, 639, 559, 592, 1125, 833, 1190, 967, 1220, 1051, 1024, 679, 1283, 575, 108, 199]. setting [351]. several [1053]. Sharp [787]. Sharply [808]. sheaves [1054, 1311]. shifts [1183]. Shimura [1083, 1300]. sided [704]. sign [989, 734]. Signatures [1163]. Simple [1059, 1186, 98, 1276, 1088, 338, 459, 464, 595, 1255, 1280, 995, 985, 449, 751, 423, 1300, 1298, 1027, 327, 789, 1021, 927, 742]. simplex [42]. simplicial [575]. simplicity [664]. simply [1032, 400]. simultaneous [1028]. Singer [60]. Singular [272, 37, 909, 182, 20, 408, 108, 247, 504]. six [233]. Skew [744, 748, 943]. skew-equivalent [943]. Skew-Hadamard [744]. skew-symmetric [748]. skews [1148]. slit [1184]. small [413, 1152, 322]. smallest [157]. Smirnov [1216]. Smith [1252]. smooth [1212, 414, 146, 402]. Sobolev [202, 1105]. solid [765]. soluble [950]. Solution [740, 171, 1299]. Solutions [444, 538, 1005, 1247, 866, 722, 165]. solvability [8, 1041]. Solvable [1231, 844, 635, 696, 367, 149, 97, 470, 222]. Solving [662]. Some [202, 978, 380, 1124, 656, 921, 711, 576, 442, 548, 72, 1200, 553, 1070, 820, 637, 1219, 423, 778, 918, 289, 937, 771, 1015, 1159, 1218, 191, 621, 962, 661, 490, 575, 543, 139, 1236, 216, 1042, 557, 86, 259, 985, 811, 1059, 689, 551, 334, 668, 1300, 1117, 550, 1, 26, 1211]. sommantes [518]. Sous [518, 1018, 739, 764]. sous-espaces [739]. sous-groupes [764]. Sous-modules [518]. sous-stochastiques [1018]. SP [611]. SP-hull [611]. space [195, 499, 520, 371, 737, 535, 1223, 1189, 812, 740, 731, 1089, 161, 1114, 78, 1074, 1241, 305, 1180, 1207, 822, 904, 323, 1220, 47, 116, 355, 570, 883, 987, 1284, 966]. Spaces [889, 806, 202, 1105, 790, 439, 908, 122, 1182, 838, 446, 733, 270, 341, 231, 1303, 953, 1042, 532, 1081, 520, 762, 1282, 881, 517, 1091, 256, 1058, 1193, 1161, 110, 48, 794, 866, 649, 1305, 780, 1066, 81, 961, 732, 318, 715, 353, 1100, 919, 1167, 938, 1302, 975, 1145, 856, 289, 99, 100, 933, 922, 984, 1202, 809, 857, 560, 1233, 1259, 480, 1221, 319, 774, 369, 1197, 321, 466, 1113, 260, 672, 1125, 1215, 496, 962, 654, 1128, 1244, 1273, 497, 734, 1178, 1051, 287, 1006, 1261, 140, 514, 531, 1134, 165, 855, 379, 1076, 141]. spaces [781, 1110, 793, 617, 345]. span [433]. Special [269, 614, 903, 889, 1030]. species [645]. Spector [171]. spectra [999, 20]. Spectral [476, 622, 273, 811, 116, 410]. spectre [643]. spectrum [1243, 453, 194, 230, 643, 1178, 901, 1021]. spectrums [1030]. speedable [1228]. Sphere [228, 424]. spheres [828, 212, 546]. spherical [834, 54]. sphericity [515]. spin [246]. spline [1001]. splines [360, 852]. Split [1023, 862, 1287]. splits [302]. splitting [796, 537, 200, 232, 789, 612]. spots [279]. spreads [186, 797, 1130, 967]. square [121, 523, 1241]. square-distance [1241]. square-free [523]. squares [813, 968, 1050, 252]. stabilité [1226]. stability [1071, 1226]. Stable [895, 1267, 1239, 212, 676, 831, 851, 1109, 529, 989]. stably [1157]. stacking [901]. standard [947, 576, 1264]. star [58]. star-like [58]. starlike [719]. starshaped [364, 1024]. states [525, 836]. stationary [136, 429]. Steenrod [998]. Steiner [1164, 552, 709, 800, 120, 939, 945]. Stiefel [303]. Stieltjes [1032, 121, 170]. stochastic [1012, 1246, 587, 239]. stochastiques [1018]. Stone [770, 368, 321]. Strict [1282, 609, 802, 975, 1145, 351]. strictly [1130]. strong [1171, 416, 873, 296, 432, 403]. Strongly [505, 690, 70, 1111, 904]. structural [75, 877]. Structure [645, 795, 1121, 746, 1139, 926, 875, 517, 221, 397, 589, 428, 636, 191, 1215, 472, 170]. structured [817]. structures [187, 527, 1295, 1162, 133, 1111, 885]. studying [510]. Sturm [145, 293]. subalgebras [783, 571, 361]. subbase [78]. subcategories [960, 715, 1250, 1065]. subdirect [862]. subdivisions [979]. subdomains [1105]. Subfields [1085, 935, 576]. subgraphs [442]. subgroup [676, 1308, 1310, 876, 300, 1044, 249, 551, 451, 1229]. Subgroups [487, 90, 486, 217, 1297, 8, 138, 563, 195, 511, 267, 395, 602, 192, 1168, 301, 870, 253, 254, 1101, 863, 1208, 376, 764, 483, 149, 1118, 579, 207, 583, 769, 1057, 1137, 359, 1231]. subidealizers [1043, 1146]. Subinvariance [844]. Sublattices [67, 1206, 787, 483, 1191, 1201]. submanifolds [668, 355, 49]. submatrices [325]. submodule [182, 302]. submodules [751, 518]. subnormal [1077, 1101, 1214, 1137]. subordination [436]. subparacompact [345]. subplanes [275]. Subregular [797]. Subrings [1106, 142, 1094]. subschemes [127]. subsemigroups [235]. subsets [147, 1114, 78, 593, 1221, 321, 345]. subsocles [152]. subspace [983, 1274]. subspaces [310, 711, 1189, 690, 383, 925, 1213]. substitutions [631]. Substochastic [1018]. Sufficiency [619, 485]. sufficient [1041, 213, 509]. sufficiently [229, 513]. Suites [1149, 1142]. Sullivan [1170]. Sum [968, 466, 832, 1140, 101]. summabilities [71]. summability [1171, 924, 603, 1007, 569, 1156, 29, 30, 240, 296, 206, 14, 219]. summands [1287]. summation [489]. Sums [284, 563, 1183, 501, 518, 1251, 848, 205, 437, 460, 474, 493, 614]. Super [353]. Super-reflexive [353]. superadditive [1246]. supercompact [1182]. supersoluble [210]. surface [251]. Surfaces [66, 1242, 398, 906, 109, 319, 1153]. Suslinian [1090]. Suspension [134, 804]. suspensions [339]. Sylow [1044]. symbols [494]. symétries [1253]. Symmetric [1116, 909, 87, 580, 894, 192, 567, 780, 748, 856, 910, 658, 560, 1215, 271]. symmetries [1224]. Symmetrizable [1006]. Symmetrized [564, 784]. symmetry [951, 1253, 38]. symplectic [164, 757]. Synthèse [764]. synthesis [622, 764]. system [671]. Systèmes [886]. Systems [934, 123, 1027, 886, 111, 686, 61, 552, 709, 800, 1164, 670, 755, 932, 1258, 120, 996, 116, 945, 907]. Szász [1225].

table [876]. tables [443]. Takahashi [309]. tame [459]. Taming [338]. tangential [833]. Tauberian [206, 489, 533]. Taylor [725]. Tchebycheff [886]. technique [719, 1124, 510]. temperatures [386]. ten [997]. Tensor [232, 1277, 687, 1122, 873, 164, 328, 490, 577]. tensors [784]. term [1287, 3, 160]. termes [1226]. terminals [282]. terms [1226, 1293, 793]. ternary [683]. Testing [432]. th [949, 1133, 740, 832, 177]. their [536, 1257, 237, 656, 576, 1174, 573, 370, 126, 320, 1209, 21, 1231, 49, 456]. them [369]. Theorem [884, 65, 790, 39, 831, 644, 988, 923, 279, 507, 360, 1181, 656, 840, 1217, 625, 1047, 144, 555, 1058, 541, 768, 539, 422, 869, 731, 732, 225, 2, 415, 62, 1252, 376, 551, 241, 342, 933, 473, 471, 91, 120, 1107, 358, 774, 1132, 173, 1296, 399, 248, 206, 846, 489, 207, 506, 654, 374, 447, 219, 885, 825, 456, 927]. théorème [39]. theorems [202, 435, 1139, 446, 578, 882, 561, 759, 436, 216, 86, 259, 1225, 1246, 721, 918, 273, 1059, 1036, 1151, 367, 937, 663, 944, 613, 96, 861, 466, 408, 191, 756, 139, 361]. theoretic [925]. theories [524, 1108, 750, 848]. Theory [1019, 605, 1183, 980, 871, 278, 1007, 1219, 788, 1147, 2, 415, 1067, 689, 396, 922, 440, 1202, 771, 615, 1292, 850, 351, 344, 129, 621, 243, 512, 522, 810, 116, 1285, 21, 410, 315, 139]. theta [1150]. Thick [1143]. third [87, 194]. Thom [529, 545]. Three [434, 1242, 424, 1305, 726, 7]. three-dimensional [726]. three-generator [7]. three-relation [7]. tight [928, 1022, 845, 1230]. time [1056]. timelike [122]. Tit [327]. Toeplitz [349]. Tomber [173]. Topics [276]. topolgical [235]. Topological [670, 646, 716, 875, 960, 137, 1114, 586, 1064, 879, 292, 351, 514, 570, 1029, 189]. Topologies [1194, 1102, 516, 609, 802]. topology [563, 1282, 194, 975, 1145, 351, 1065]. Torsion [990, 524, 467, 692, 195, 501, 283, 848, 302, 789, 608, 262]. torsion-free [195, 501]. torsions [964]. torus [765]. total [573]. totally [711]. tout [914]. Towber [131]. trace [1307, 1083, 544]. transcendence [1019]. transfer [1043, 1146]. transferability [787]. transferable [808]. Transferring [688]. transform [931]. transformation [426, 728, 879, 925, 191, 996, 1021]. Transformations [1241, 901, 753, 87, 476, 341, 1033, 48, 598, 649, 44, 1099, 56, 868, 1048, 533, 745, 907]. transforms [1032, 121, 393, 725, 170, 1304]. transitif [1211]. Transitive [163, 1211, 36, 527, 705]. Transitivity [898]. Translation [1017, 978, 835, 168, 472]. Translation-invariant [1017]. translational [678, 628]. translations [409]. transvections [665]. transversalité [1226]. transversality [1226, 1196]. transversals [619]. traveling [782]. trees [10, 1172, 196]. Treillis [891, 409]. triangulations [437, 460, 474, 493, 614]. trigonometric [1035]. triple [770, 552, 709, 800, 1164, 945]. triples [800, 1164]. trivalent [151]. trivial [1044, 384]. Tsuji [1011]. tubular [1185]. Tutte [1252]. Two [919, 287, 717, 1287, 707, 1280, 1294, 1023, 971, 1301, 832, 249, 19, 1180]. type [578, 921, 630, 373, 529, 809, 944, 37, 441, 1227, 744, 976, 141]. types [1224, 399, 510].

UHF [1253]. ultrafilters [346]. Ultrafunctors [723]. ultrapowers [346]. ultraproducts [372]. ultraspherical [699, 272]. unbounded [1304]. unconditionally [304]. Unequivocal [754]. unified [6]. Uniform [726, 184, 331, 833, 1190, 439, 834, 693, 414, 592]. Uniformities [295, 1202, 497]. Uniformly [649, 256, 377, 1051, 1261, 379]. unimodular [172, 261]. Union [1302]. unions [724]. unipotent [976]. Unique [361, 412, 841, 977]. Uniquely [954]. Uniqueness [1139, 408, 1296, 1279]. uniserial [201]. Unit [694, 36, 1089, 913]. Unitaries [1255, 1280]. unitarized [1075]. unitary [1188, 257, 1224, 969, 85, 242]. units [963, 384, 832, 910, 1285]. univalent [435, 18, 882, 436, 1014, 1117, 550, 227]. universal [960, 585, 309]. universality [432]. unrestricted [1234]. unsolvability [1059, 67]. Unsolvable [97]. unsymmetric [419]. unsymmetrical [1, 26]. upcrossing [1037]. upper [454, 1166].

V [614]. valency [1306]. valent [477].

valeurs [335, 812]. valuation [1047, 208, 298]. Valuations [572, 615, 33, 1102, 298]. Value [894, 871, 625, 93, 613, 534, 504]. valued [875, 1282, 986, 609, 802, 559, 607, 79]. values [335, 812, 942, 1251, 146, 872, 1025]. vanishing [371, 47]. variable [216, 354]. variables [1032]. Variation [718]. variational [719]. variations [1009]. varietal [682]. Varieties [236, 291, 939, 1008, 858, 102, 549, 1083, 1300, 1072, 1290, 827]. variety [682, 19, 124]. various [1073, 734]. Veblen [113]. vector [1282, 986, 1189, 609, 802, 268, 334, 990, 559, 558, 277, 472, 514]. vector-valued [1282, 609, 802]. vectors [496]. versions [555]. Vertex [149]. vertices [997, 979]. very [1135]. via [573, 1100, 607]. view [243, 512]. visibility [364]. Vitali [1293]. volumes [558]. vs [801]. Vulikh [379].

W [959, 1019]. Waerden [566]. Watson [393, 280]. ways [463]. Weak [117, 953, 1002, 1026, 890, 216, 1058, 637, 1135, 654, 886]. Weakly [1090, 839, 683, 505, 574, 678, 1051, 1006]. Wedderburn [113, 447]. Wei-eihn [959]. Weierstrass [912, 250]. weighing [1173]. weighted [1183, 433, 140, 247]. Weyl [585]. where [747]. Which [532, 27, 889, 464, 1308, 1310, 482, 832, 377, 1241, 544, 451, 496, 196, 1039, 531, 575, 1021]. Whitehead [1170, 1210, 1050]. whose [1287, 818, 864, 705, 983, 910]. width [16]. Wiener [892]. wild [338]. without [201, 1185, 881, 1193, 1084]. Witt [1047, 1163, 1132]. word [97]. words [588]. wreath [119, 905].

Yosida [388]. Young [892].

Zassenhaus [783]. Zeller [853]. Zero [849, 617, 201, 182, 238, 190, 370, 336, 94]. Zerodimensional [617]. zeros [1222, 484, 45]. zeta [1300]. zonotopes [562].

References

Shrikhande:1963:NES

[1] S. S. Shrikhande, D. Raghavarao, and S. K. Tharthare. Non-existence of some unsymmetrical partially balanced incomplete block designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 686–701, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [26].

Hughes:1968:TCP

[2] J. W. B. Hughes. A theorem concerning partitions and its consequences in the theory of Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 698–700, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [415].

Sreenivasulu:1969:MTR

[3] V. Sreenivasulu. The maximum term and the rank of an entire function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):257–261, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [160].

Pedersen:1970:IBC

[4] Franklin D. Pedersen. Implications between conditions on l-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):1–6, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Flatto:1970:RPH

[5] Leopold Flatto and Sister Margaret M. Wiener. Regular polytopes and harmonic polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):7–21, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Biondi:1970:CPC

[6] E. Biondi, L. Divieti, and G. Guardabassi. Counting paths, circuits, chains, and cycles in graphs: A unified approach. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):22–35, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wamsley:1970:CTG

[7] J. W. Wamsley. A class of threegenerator, three-relation, finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):36–40, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beidleman:1970:IGF

[8] James C. Beidleman. The influence of generalized Frattini subgroups on the solvability of a finite group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):

41–46, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sentilles:1970:SO

[9] F. Dennis Sentilles. Semigroups of operators in C(S). Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):47–54, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Manvel:1970:RT

[10] Bennet Manvel. Reconstruction of trees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):55–60, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hanani:1970:BIB

[11] Haim Hanani. On balanced incomplete block designs with large number of elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):61–65, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hilton:1970:ARA

[12] P. J. Hilton and Y. c. Wu. On the addition of relations in an abelian category. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):66–74, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lachlan:1970:ISM

[13] A. H. Lachlan. Initial segments of many-one degrees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):75–85, ????

1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sahney:1970:NSC

[14] Badri N. Sahney. On the Nørlund summability of a class of Fourier series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):86–91, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chew:1970:RFR

[15] Kim Lin Chew and Sherry Lawn. Residually finite rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):92–101, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wichman:1970:WM

[16] Michael Wichman. The width of a module. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):102–115, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meyers:1970:MAP

[17] W. E. Meyers. Montel algebras on the plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):116–122, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Basgoze:1970:CUF

[18] T. Basgöze, J. L. Frank, and F. R. Keogh. On convex univalent functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):123–127, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuan:1970:HST

[19] Wei eihn Kuan. On the hyperplane sections through two given points of an algebraic variety. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):128–133, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putnam:1970:SSN

[20] C. R. Putnam. The spectra of semi-normal singular integral operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):134–150, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Streit:1970:MIG

[21] Franz Streit. On multiple integral geometric integrals and their applications to probability theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 151–163, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Corson:1970:C

[22] H. H. Corson and I. Glicksberg. Compactness in Hom(G, H). Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 164–170, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gasper:1970:LPJa

[23] George Gasper. Linearization of the product of Jacobi polynomials. I.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):171–175, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dark:1970:CCF

[24] Rex Dark and Akbar H. Rhemtulla. On R_{0}-closed classes, and finitely generated groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):176–184, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rigby:1970:AAG

[25] J. F. Rigby. Axioms for absolute geometry. III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):185–190, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shrikhande:1970:NES

[26] S. S. Shrikhande, D. Raghavarao, and S. K. Tharthare. “Non-existence of some unsymmetrical partially balanced incomplete block designs”: Corrigendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):191, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [1].

Beasley:1970:MGW

[27] LeRoy B. Beasley. Maximal groups on which the permanent is multiplicative. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):192, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fox:1970:MIK

[28] R. H. Fox. Metacyclic invariants of knots and links. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):193–201, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [500].

Korle:1970:ASRa

[29] H.-H. Körle. On absolute summability by Riesz and generalized Cesàro means. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):202–208, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Korle:1970:ASRb

[30] H.-H. Körle. On absolute summability by Riesz and generalized Cesàro means. I, II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):209–218, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kubota:1970:CAC

[31] Yôto Kubota. A characterization of the approximately continuous Denjoy integral. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):219–226, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bressler:1970:CSO

[32] D. W. Bressler and A. H. Cayford. Compositions of set operations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):

227–234, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Perles:1970:CCV

[33] M. A. Perles and G. T. Sallee. Cell complexes, valuations, and the Euler relation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):235–241, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garfinkel:1970:GEM

[34] Gerald Garfinkel and Morris Orzech. Galois extensions as modules over the group ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):242–248, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coleman:1970:GR

[35] D. B. Coleman. On group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 249–254, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borrego:1970:PTA

[36] J. T. Borrego and E. E. DeVun. Pointtransitive actions by the unit interval. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):255–259, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pfeffer:1970:SIP

[37] W. F. Pfeffer. Singular integrals are Perron integrals of a certain type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):260–264, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McMullen:1970:PAS

[38] P. McMullen and G. C. Shephard. Polytopes with an axis of symmetry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):265–287, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arminjon:1970:TRP

[39] Paul Arminjon. Un théorème de régularité pour une équation differentielle abstraite. (French) [A regularity theorem for an abstract differential equation]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):288–296, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Band:1970:IEQ

[40] Melvin Band. On the integral extensions of quadratic forms over local fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):297–307, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berg:1970:ECM

[41] I. D. Berg. Extensions of certain maps to automorphisms of m. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 308–316, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fine:1970:SCM

[42] N. J. Fine and P. H. Maserick. On the simplex of completely monotonic functions on a commutative semigroup. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):317–326, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1970:PEL

[43] E. W. Johnson and J. A. Johnson. Mprimary elements of a local Noether lattice. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):327–331, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [438].

Hardie:1970:HNT

[44] K. A. Hardie. Homotopy of natural transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):332–341, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wells:1970:ZFD

[45] James Wells. On the zeros of functions with derivatives in H_{1} and H_{∞}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):342–347, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kleinfeld:1970:GCA

[46] Erwin Kleinfeld, Margaret Humm Kleinfeld, and Frank Kosier. A generalization of commutative and alternative rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):348–362, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Singh:1970:VPS

[47] K. Singh. On the vanishing of a (G, σ) product in a (G, σ) space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 363–371, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edelstein:1970:RMC

[48] Michael Edelstein. On the representation of mappings of compact metrizable spaces as restrictions of linear transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):372–375, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yano:1970:IFS

[49] Kentaro Yano. Integral formulas for submanifolds and their applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):376–388, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1970:ICR

[50] P. M. Gauthier. it Cercles de remplissage and asymptotic behaviour along circuitous paths. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):389–393, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lakser:1970:NCR

[51] H. Lakser. Normal and canonical representations in free products of lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):394–402, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Biggs:1970:NBN

[52] R. G. Biggs. A new bound for nil U-rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):403–407, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1970:IIE

[53] G. F. D. Duff. Integral inequalities for equimeasurable rearrangements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 408–430, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Petty:1970:CSN

[54] C. M. Petty and J. M. Crotty. Characterizations of spherical neighbourhoods. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):431–435, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

King:1970:DIR

[55] R. C. King. The dimensions of irreducible representations of linear groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):436–448, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mizel:1970:CNL

[56] Victor J. Mizel. Characterization of non-linear transformations possessing kernels. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):449–471, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baker:1970:DPL

[57] Kirby A. Baker and Alfred W. Hales. Distributive projective lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 472–475, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brannan:1970:CEC

[58] D. A. Brannan, J. Clunie, and W. E. Kirwan. Coefficient estimates for a class of star-like functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 476–485, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eenigenburg:1970:RCC

[59] Paul Eenigenburg. On the radius of curvature for convex analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):486–491, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hestenes:1970:SG

[60] Marshall D. Hestenes. Singer groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):492–513, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kallman:1970:PGR

[61] Robert R. Kallman. A problem of Gelfand on rings of operators and dynamical systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):514–517, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Iden:1970:RTD

[62] Oddvar Iden. Remark to a theorem due to R. Sandler. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):518, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Orlik:1970:AII

[63] Peter Orlik. On the Arf invariant of an involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):519–524, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Irwin:1970:DPA

[64] John M. Irwin and John D. O’Neill. On direct products of Abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):525–544, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akcoglu:1970:LRT

[65] M. A. Akcoglu and R. V. Chacon. A local ratio theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):545–552, ????

1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jaco:1970:SE

[66] William Jaco. Surfaces embedded in M^{2} × S^{1}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):553–568, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thomason:1970:SIS

[67] S. K. Thomason. Sublattices and initial segments of the degrees of unsolvability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):569–581, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gasper:1970:LPJb

[68] George Gasper. Linearization of the product of Jacobi polynomials. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):582–593, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Read:1970:CGC

[69] R. C. Read and E. M. Wright. Coloured graphs: A correction and extension. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):594–596, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goethals:1970:SRG

[70] J.-M. Goethals and J. J. Seidel. Strongly regular graphs derived from combinatorial designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):597– 614, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Izumi:1970:LPA

[71] Masako Izumi and Shin ichi Izumi. Localization problem of the absolute Riesz and absolute Nörlund summabilities of Fourier series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):615–625, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ford:1970:SRS

[72] Charles Ford. Some results on the Schur index of a representation of a finite group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):626–640, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Plemmons:1970:CCS

[73] R. J. Plemmons. On a conjecture concerning semigroup homomorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):641–644, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McArthur:1970:HRP

[74] William G. McArthur. Hewitt realcompactifications of products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 645–656, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1970:SAN

[75] E. W. Johnson, J. A. Johnson, and J. P. Lediaev. A structural approach to Noether lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):657–665, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gotzky:1970:PRA

[76] K. Martin Götzky. Products of reflections in an affine Moufang plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 666–673, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaufman:1970:BMD

[77] Robert Kaufman. Brownian motion and dimension of perfect sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 674–680, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hursch:1970:SPS

[78] J. L. Hursch, Jr. Screening properties of the subbase of all closed connected subsets of a connectedly generated space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):681–685, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sherman:1970:EOV

[79] Malcolm J. Sherman. Eigenfunctions of operator-valued analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):686–690, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Franklin:1970:OEM

[80] S. P. Franklin and J. K. Kohli. On open extensions of maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):691–696, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heath:1970:IHE

[81] Philip R. Heath. Induced homotopy equivalences on mapping spaces and duality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):697–700, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rao:1970:CDR

[82] M. L. Narayana Rao. Conjecture of D. R. Hughes extended to generalized André planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):701–704, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wakae:1970:MPIa

[83] Masami Wakae and Oma Hamara. A mapping problem and J_{p}-index. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):705–712, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wakae:1970:MPIb

[84] Masami Wakae and Oma Hamara. A mapping problem and J_{p}-index. I, II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):713–718, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rider:1970:CIM

[85] Daniel Rider. Central idempotent measures on unitary groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 719–725, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clements:1970:ETS

[86] John C. Clements. Existence theorems for some non-linear equations of evolution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):726–745, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [259].

Beasley:1970:LTM

[87] LeRoy B. Beasley. Linear transformations on matrices: The invariance of the third elementary symmetric function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):746–752, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hursey:1970:OGS

[88] R. J. Hursey, Jr. and A. H. Rhemtulla. Ordered groups satisfying the maximal condition locally. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):753–758, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deliyannis:1970:GRC

[89] Platon C. Deliyannis. Group representations and cardinal algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 759–772, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burroughs:1970:SCC

[90] John E. Burroughs and James A. Schafer. Subgroups of conjugate classes in extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):773–783, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Liu:1970:THC

[91] Ming chit Liu. On a theorem of Heilbronn concerning the fractional part of θn^{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):784–788, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Turgeon:1970:RNA

[92] J. Turgeon. On the rank numbers of an arc. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):789–802, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1970:MMN

[93] Paul Gauthier. The maximum modulus of normal meromorphic functions and applications to value distribution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):803–814, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ward:1970:SPM

[94] D. J. Ward. A set of plane measure zero containing all finite polygonal arcs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):815–821, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dlab:1970:CPR

[95] Vlastimil Dlab. On a class of perfect rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):822–826, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCool:1970:ETC

[96] James McCool. Embedding theorems for countable groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 827–835, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCool:1970:UPG

[97] James McCool. Unsolvable problems in groups with solvable word problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):836–838, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moody:1970:SQE

[98] Robert V. Moody. Simple quotients of Euclidean Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 839–846, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kwon:1970:MPB

[99] Y. K. Kwon and L. Sario. A maximum principle for bounded harmonic functions on Riemannian spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 847–854, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kwon:1970:MPD

[100] Y. K. Kwon and L. Sario. A maximum principle for direchlet-finite harmonic functions on Riemannian spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):855–862, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ruckle:1970:ACS

[101] William H. Ruckle. An abstract concept of the sum of a numerical series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):863–874, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gupta:1970:MVG

[102] Narain Gupta. On metanilpotent varieties of groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):875–877, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cofman:1970:BIF

[103] Judita Cofman. On Baer involutions of finite projective planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 878–880, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:1970:ECD

[104] K. B. Lee. Equational classes of distributive pseudo-complemented lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):881–891, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Price:1970:PDF

[105] J. F. Price. On positive definite functions over a locally compact group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):892–896, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kantor:1970:ED

[106] William M. Kantor. Elations of designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):897–904, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanQue:1970:CLP

[107] Ngô van Quê. On the classification of Lie pseudo-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):905–915, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Timourian:1970:MLF

[108] J. G. Timourian. Maps with locally flat singular sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):916–921, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moore:1970:FPA

[109] M. J. Moore. Fixed points of automorphisms of compact Riemann surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):922–932, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edelstein:1970:LHA

[110] M. Edelstein. Lines and hyperplanes associated with families of closed and bounded sets in conjugate Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):933–938, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davis:1970:ISA

[111] Elwyn H. Davis. Incidence systems associated with non-planar nearfields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):939–952, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morgan:1970:GAR

[112] Raymond V. Morgan, Jr. On a generalization of alternative rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):

953–966, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stevenson:1970:CGV

[113] Frederick W. Stevenson. The collineation group of the Veblen– Wedderburn plane of order nine. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 967–973, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lallement:1970:EBS

[114] Gérard Lallement and Mario Petrich. Extensions of a Brandt semigroup by another. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):974–983, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shapiro:1970:MEC

[115] H. L. Shapiro. More on extending continuous pseudometrics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 984–993, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stampfli:1970:STD

[116] J. G. Stampfli. A spectral theory for duality systems of operators on a Banach space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):994–996, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ball:1970:WNR

[117] Eugene S. Ball. Weak normality and related properties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):997–1001, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blumenthal:1970:MFA

[118] Robert G. Blumenthal. Maximality in function algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1002–1004, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1970:MRC

[119] D. L. Johnson. Minimal relations for certain wreath products of groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):1005–1009, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mendelsohn:1970:TSS

[120] N. S. Mendelsohn. A theorem on Steiner systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1010–1012, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Henniger:1970:FBM

[121] J. Henniger. Functions of bounded mean square, and generalized Fourier– Stieltjes transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1016–1034, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beem:1970:IFS

[122] John K. Beem. Indefinite Finsler spaces and timelike spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1035–1039, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1970:SEG

[123] I. M. Isaacs. Systems of equations and generalized characters in groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 1040–1046, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuan:1970:NGH

[124] Wei eihn Kuan. A note on a generic hyperplane section of an algebraic variety. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1047–1054, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Derzko:1970:GPA

[125] N. A. Derzko. Gentle perturbations of i, d/dx with application to −d^{2}/dx^{2}+q. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):1055–1070, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Keimel:1970:ACE

[126] Klaus Keimel. Algèbres commutatives engendrées par leurs éléments idempotents. (French) [Commutative algebras generated by their idempotent elements]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1071–1078, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shatz:1970:FSG

[127] Stephen S. Shatz. Finite subschemes of group schemes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1079–1081, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lick:1970:DG

[128] Don R. Lick and Arthur T. White. kdegenerate graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1082–1096, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ortiz:1970:CGR

[129] Augusto H. Ortiz. A construction in general radical theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1097– 1100, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Singh:1970:CSI

[130] Surjeet Singh and Kamlesh Wasan. Commutative self-injective rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 1101–1108, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lissner:1970:RTR

[131] David Lissner and Anthony Geramita. Remarks on OP and Towber rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):1109–1117, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cutler:1970:GFR

[132] Doyle O. Cutler and Paul F. Dubois. A generalization of final rank of primary Abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1118–1122, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kumpera:1970:ELS

[133] A. Kumpera. Équivalence locale des structures de contact de codimension un. (French) [Local equivalence of contact structures of codimension one]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):1123–1128, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilbert:1970:SLS

[134] William J. Gilbert. Suspension of the Lusternik–Schnirelmann category. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):1129–1132, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Raphael:1970:AEC

[135] R. M. Raphael. Algebraic extensions of commutative regular rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 1133–1155, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coleman:1970:GDE

[136] Courtney Coleman. Growth and decay estimates near non-elementary stationary points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1156–1167, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [429].

Christoph:1970:IET

[137] Francis T. Christoph, Jr. Ideal extensions of topological semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 1168–1175, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benabdallah:1970:QES

[138] Khalid Benabdallah and John M. Irwin. On quasi-essential subgroups of primary abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 1176–1184, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yap:1970:SAT

[139] H. P. Yap. Some addition theorems of group theory with applications to graph theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1185–1195, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Summers:1970:FCW

[140] W. H. Summers. Full-completeness in weighted spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1196–1207, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vaughan:1970:PMS

[141] J. E. Vaughan. Perfect mappings and spaces of countable type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 1208–1210, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fisher:1970:NNS

[142] Joe W. Fisher. On the nilpotency of nil subrings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1211–1216, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Patty:1970:HDP

[143] C. W. Patty. Homology of deleted products of contractible 2-dimensional polyhedra. III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1217–1223, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohn:1970:DDB

[144] Richard M. Cohn. A differencedifferential basis theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 1224–1237, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorch:1970:HMP

[145] Lee Lorch, M. E. Muldoon, and Peter Szego. Higher monotonicity properties of certain Sturm–Liouville functions.

III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):1238–1265, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taylor:1970:IRA

[146] B. A. Taylor and D. L. Williams. Ideals in rings of analytic functions with smooth boundary values. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??): 1266–1283, ???? 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clements:1970:EDR

[147] G. F. Clements. On existence of distinct representative sets for subsets of a finite set. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(6):1284–1292, December 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gorman:1971:ICN

[148] Howard Gorman. Invertibility and class number of orders. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):1–11, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Malzan:1971:VSI

[149] J. Malzan. Vertex subgroups of irreducible representations of solvable groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):12–21, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sachs:1971:GMG

[150] David Sachs. Geometric mappings on geometric lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):22–35, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robertson:1971:PGT

[151] Neil Robertson. Pentagon-generated trivalent graphs with girth 5. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 36–47, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Irwin:1971:PSP

[152] John Irwin and James Swanek. On purifiable subsocles of a primary abelian group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):48–57, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Krajewski:1971:ELF

[153] Lawrence L. Krajewski. On expanding locally finite collections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 58–68, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zelmanowitz:1971:CER

[154] J. Zelmanowitz. Commutative endomorphism rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):69–76, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duda:1971:FOO

[155] Edwin Duda and W. Hugh Haynsworth. Finite-to-one open mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):77–83, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burde:1971:BC

[156] Gerhard Burde. On branched coverings of S^{3}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):84–89, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harris:1971:SDP

[157] Morton E. Harris and Christoph Hering. On the smallest degrees of projective representations of the groups PSL(n, q). Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):90–102, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roggenkamp:1971:GGB

[158] Klaus W. Roggenkamp. Grothendieck groups of Bass orders. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):103– 115, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nicolas:1971:RNH

[159] Jean-Louis Nicolas. Répartition des nombres hautement composés de Ramanujan. (French) [Repartition of highly-composed Ramanujan numbers]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):116–130, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sreenivasulu:1971:MTR

[160] V. Sreenivasulu. “The maximum term and the rank of an entire function”: Corrigendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):131, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [3].

Harvey:1971:RCL

[161] Bernard Niel Harvey. Representation of certain linear operators in Hilbert space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):132–150, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Amberg:1971:GCC

[162] Bernhard Amberg. On groups with chain conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):151–159, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pnueli:1971:TOG

[163] A. Pnueli, A. Lempel, and S. Even. Transitive orientation of graphs and identification of permutation graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):160–175, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

King:1971:DIT

[164] R. C. King. The dimensions of irreducible tensor representation of the orthogonal and symplectic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 176–188, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ton:1971:PSN

[165] Bui An Ton. Periodic solutions of non-linear evolution equations in Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):189–196, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gorman:1971:RRD

[166] Howard E. Gorman. Radical regularity in differential rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):197–201, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beachy:1971:BCF

[167] John A. Beachy. Bicommutators of cofaithful, fully divisible modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 202–213, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [556].

Rado:1971:CPI

[168] F. Radó. Congruence-preserving isomorphisms of the translation group associated with a translation plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):214–221, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ravisankar:1971:CNA

[169] T. S. Ravisankar. Characteristically nilpotent algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):222–235, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schwartz:1971:SAH

[170] Alan Schwartz. The structure of the algebra of Hankel transforms and the algebra of Hankel–Stieltjes transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):236–246, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lachlan:1971:SPS

[171] A. H. Lachlan. Solution to a problem of Spector. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):247–256, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fisher:1971:AUFa

[172] Stephen Fisher. Approximation by unimodular functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):257–269, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [261].

Myung:1971:RPT

[173] Hyo Chul Myung. A remark on the proof of a theorem of Laufer and Tomber. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):270, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marshall:1971:RGA

[174] Murray A. Marshall. Ramification groups of abelian local field extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):271–281, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Powers:1971:DCM

[175] David L. Powers. On the differentials of certain matrix functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 282–286, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mitchem:1971:PAG

[176] John Mitchem. On the point-arboricity of a graph and its complement. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 287–292, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Willett:1971:ABD

[177] D. Willett. Asymptotic behaviour of disconjugate nth order differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):293–314, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mercer:1971:CFI

[178] A. McD. Mercer. On certain functional identities in E^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):315–324, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sagle:1971:RLA

[179] Arthur A. Sagle. On reductive Lie admissible algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):325–331, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Askey:1971:LPJ

[180] Richard Askey and George Gasper. Linearization of the product of Jacobi polynomials. III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):332–338, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cramer:1971:BAR

[181] Timothy Cramer. Boolean algebra retracts. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):339–344, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cateforis:1971:MSS

[182] Vasily C. Cateforis and Francis L. Sandomierski. On modules of singular submodule zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):345–354, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boyd:1971:EOP

[183] David W. Boyd. On the exponent of an osculatory packing. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):355–363, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lowenthal:1971:UFG

[184] Franklin Lowenthal. Uniform finite generation of isometry groups of Euclidean and non-Euclidean geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):364–373, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stauffer:1971:RBL

[185] H. B. Stauffer. A relationship between left exact and representable functors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):374–380, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruen:1971:PSR

[186] A. Bruen. Partial spreads and replaceable nets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):381–391, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Atherton:1971:LPC

[187] C. R. Atherton, Jr. On the lattice of primitive convergence structures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):392–397, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marshall:1971:MEL

[188] Murray A. Marshall. The maximal p-extension of a local field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 398–402, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wong:1971:TDP

[189] H. Ship fah Wong. The topological degree of A-proper maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 403–412, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hanson:1971:ALC

[190] T. H. McH. Hanson. Actions of a locally compact group with zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 413–420, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roberson:1971:STS

[191] Fred A. Roberson. Some theorems on the structure of nearly equicontinuous transformation groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 421–425, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dixon:1971:MAS

[192] John D. Dixon. Maximal abelian subgroups of the symmetric groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):

426–438, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harris:1971:ICR

[193] Douglas Harris. An internal characterization of realcompactness. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 439–444, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gardner:1971:TDT

[194] L. Terrell Gardner. On the “third definition” of the topology on the spectrum of a C^{∗}-algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):445–450, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Best:1971:TFD

[195] L. A. Best. On torsion-free discrete subgroups of PSL(2, C) with compact orbit space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):451–460, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schirmer:1971:BMT

[196] Helga Schirmer. Biconnected multifunctions of trees which have an end point as fixed point or coincidence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):461–467, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Friedman:1971:AO

[197] N. A. Friedman and A. E. Tong. On additive operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):468–480, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dembowski:1971:GHP

[198] Peter Dembowski. Generalized Hughes planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):481–494, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Woods:1971:HSR

[199] R. Grant Woods. Homeomorphic sets of remote points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):495–502, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hou:1971:QSE

[200] Hsiang dah Hou. Quasi-splitting exact sequence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):503–506, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Behrens:1971:AQU

[201] Ernst-August Behrens. The arithmetic of the quasi-uniserial semigroups without zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):507–516, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adams:1971:SIT

[202] R. A. Adams and John Fournier. Some imbedding theorems for Sobolev spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):517–530, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Turyn:1971:MAP

[203] Richard J. Turyn. On C-matrices of arbitrary powers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):531–535, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rosenberg:1971:PGA

[204] Alan Rosenberg. On the primitivity of the group algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):536–540, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sehgal:1971:CSA

[205] Sudarshan K. Sehgal. On class sums in p-adic group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):541–543, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Peterson:1971:BPT

[206] G. E. Peterson. A best possible Tauberian theorem for the collective continuous Hausdorff summability method. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):544–549, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Randolph:1971:ETJ

[207] John W. Randolph. An extension of a theorem of Janko on finite groups with nilpotent maximal subgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 550–552, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wright:1971:NPF

[208] Michael J. Wright. On the number of prolongations of a finite rank valuation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):553–556, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cross:1971:GI

[209] G. E. Cross. On the generality of the AP-integral. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):557–561, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1971:NSG

[210] D. L. Johnson. A note on supersoluble groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):562–564, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Salinas:1971:FBP

[211] Norberto Salinas. On the η function of Brown and Pearcy and the numerical function of an operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(4):

565–578, August 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Antonelli:1971:SDE

[212] P. L. Antonelli. On stable diffeomorphism of exotic spheres in the metastable range. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):579–587, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Radjavi:1971:SCO

[213] Heydar Radjavi and Peter Rosenthal. A sufficient condition that an operator algebra be self-adjoint. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 588–597, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nielsen:1971:ARS

[214] Ole A. Nielsen. The asymptotic ratio set and direct integral decompositions of a von Neumann algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 598–607, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fieldhouse:1971:CM

[215] David J. Fieldhouse. Characterizations of modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):608–610, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carroll:1971:ETS

[216] Robert Carroll and Emile State. Existence theorems for some weak abstract variable domain hyperbolic problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):611–626, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Karrass:1971:SGG

[217] A. Karrass and D. Solitar. Subgroups of HNN groups and groups with one defining relation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):627–643, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kreith:1971:DCN

[218] Kurt Kreith. Disconjugacy criteria for nonselfadjoint differential equations of even order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):644–652, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thorpe:1971:ITB

[219] B. Thorpe. An inclusion theorem for Bohr–Hardy summability factors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):653–658, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gudder:1971:RGA

[220] Stanley P. Gudder. Representations of groups as automorphisms on orthomodular lattices and posets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):

659–673, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gardner:1971:MBS

[221] L. Terrell Gardner. On the Mackey Borel structure. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):674–678, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wright:1971:PSL

[222] C. R. B. Wright. On primitive solvable linear groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):679–685, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Evans:1971:FHI

[223] Benny Evans. First homology of irreducible 3-manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):686–691, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lovelady:1971:MII

[224] David Lowell Lovelady. Multiplicative integration of infinite products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 692–698, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Howe:1971:KOT

[225] A. Howe. Klein’s oscillation theorem for periodic boundary conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):699–703, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McAdam:1971:GPR

[226] Stephen McAdam. Going down in polynomial rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):704–711, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stump:1971:LCU

[227] Robert K. Stump. Linear combinations of univalent functions with complex coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):712–717, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leech:1971:SPE

[228] John Leech and N. J. A. Sloane. Sphere packings and error-correcting codes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):718–745, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heil:1971:MSL

[229] Wolfgang Heil. On 3-manifolds with sufficiently large decompositions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):746–748, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [513].

Hochster:1971:MPS

[230] M. Hochster. The minimal prime spectrum of a commutative ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 749–758, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borges:1971:CFS

[231] Carlos J. R. Borges. Connectivity of function spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):759–763, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lawver:1971:TPS

[232] D. A. Lawver and E. H. Toubassi. Tensor products and the splitting of abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):764–770, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lindsey:1971:FLG

[233] J. H. Lindsey II. Finite linear groups of degree six. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):771–790, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Armacost:1971:GMG

[234] D. L. Armacost. Generators of monothetic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):791–796, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See remark [461].

Lau:1971:IMD

[235] Anthony To ming Lau. Invariant means on dense subsemigroups of topolgical groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):797–801, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruns:1971:VOL

[236] Günter Bruns and Gudrun Kalmbach. Varieties of orthomodular lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):802–810, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cailleau:1971:QIC

[237] Annie Cailleau. Quasi-injectivité de certains modules sur leur anneau d’endomorphismes. (French) [Quasiinjectivity of certain modules on their ring of endomorphisms]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 811–815, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Delsarte:1971:OMZ

[238] P. Delsarte, J.-M. Goethals, and J. J. Seidel. Orthogonal matrices with zero diagonal. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):816–832, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pierce:1971:MFR

[239] Stephen Pierce. Multilinear functions of row stochastic matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 833–843, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Korle:1971:ASR

[240] Hans-Heinrich Körle. “On absolute summability by Riesz and generalized Cesàro means. II”: Addendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (5):844, October 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Killgrove:1971:GAH

[241] R. B. Killgrove, Jason Frand, William Giles, and Henry Bray. A geometric approach to the Heine–Borel theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):845–848, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tam:1971:UEC

[242] P. K. Tam. On the unitary equivalence of certain classes of non-normal operators. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):849–856, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roth:1971:DVC

[243] Richard L. Roth. A dual view of the Clifford theory of characters of finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):857–865, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Balbes:1971:POS

[244] Raymond Balbes. On the partially ordered set of prime ideals of a distributive lattice. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):866–874, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nelson:1971:LEC

[245] Evelyn Nelson. The lattice of equational classes of commutative semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):875–895, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1971:CSR

[246] Robert B. Brown. A characterization of spin representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):896–906, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Walsh:1971:WNI

[247] T. Walsh. On weighted norm inequalities for fractional and singular integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):907–928, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Parmenter:1971:TB

[248] M. M. Parmenter. On a theorem of Bovdi. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):929–932, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Karrass:1971:FPT

[249] A. Karrass and D. Solitar. The free product of two groups with a malnormal amalgamated subgroup. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 933–959, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Larcher:1971:WPC

[250] H. Larcher. Weierstrass points at the cusps of Γ_{0}, (16p) and hyperellipticity of Γ_{0}, (n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):960–968, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pranger:1971:EPH

[251] Walter Pranger. Extreme points in the Hardy class H^{1} of a Riemann surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):969–976, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pressman:1971:OLC

[252] Irwin S. Pressman. Obstructions to liftings in commutative squares. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 977–982, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glauberman:1971:ISFa

[253] George Glauberman. Isomorphic subgroups of finite p-groups. I, II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):

983–1022, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glauberman:1971:ISFb

[254] George Glauberman. Isomorphic subgroups of finite p-groups. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 1023–1039, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Crone:1971:FBS

[255] L. Crone, D. J. Fleming, and P. Jessup. Fundamental biorthogonal sequences and K-norms on phi. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 1040–1050, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Day:1971:NLS

[256] M. M. Day, R. C. James, and S. Swaminathan. Normed linear spaces that are uniformly convex in every direction. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):1051–1059, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kantor:1971:UPF

[257] William M. Kantor. On unitary polarities of finite projective planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 1060–1077, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Skerry:1971:MCM

[258] H. B. Skerry. Mercerian conditions for the method (F, d_{n}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):1078–1085, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clements:1971:ETS

[259] John C. Clements. “Existence theorems for some nonlinear equations of evolution”: Corrigendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):1086, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [86].

Reed:1971:SMM

[260] G. M. Reed. On screenability and metrizability of Moore spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 1087–1092, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fisher:1971:AUFb

[261] Stephen Fisher. “Approximation by unimodular functions”: Corrigendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23 (??):1093, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [172].

Zelmanowitz:1971:IHT

[262] J. Zelmanowitz. Injective hulls of torsion free modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):1094–1101, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1971:IEC

[263] Richard L. W. Brown. Immersions and embeddings up to cobordism. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??): 1102–1115, ???? 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arsove:1971:PHM

[264] Maynard Arsove and Heinz Leutwiler. Pseudo harmonic measures and the Dirichlet problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(6):1116– 1120, December 1971. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chapman:1972:CNC

[265] G. R. Chapman. A class of noncentral E-functors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1–4, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Busby:1972:CIR

[266] Robert C. Busby and Irwin Schochetman. Compact induced representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):5–16, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Currano:1972:CSF

[267] John J. Currano. Conjugate psubgroups of finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):

17–28, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hanson:1972:AFV

[268] T. H. McH. Hanson. Actions that fiber and vector semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):29–37, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McKnight:1972:SR

[269] J. D. McKnight, Jr. and Gary L. Musser. Special (p; , q) radicals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 38–44, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Binz:1972:FAD

[270] E. Binz and W. Feldman. A functional analytic description of normal spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):45–49, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1972:EJI

[271] Kirby C. Smith. Extending Jordan ideals and Jordan homomorphisms of symmetric elements in a ring with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):50–59, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dunkl:1972:SIU

[272] Charles F. Dunkl. Singular integrals on ultraspherical series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):60–71, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hughes:1972:PTR

[273] Edward Hughes. Perturbation theorems for relative spectral problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):72–81, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wright:1972:CCG

[274] E. M. Wright. Counting coloured graphs. III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):82–89, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cofman:1972:BSF

[275] Judita Cofman. Baer subplanes in finite projective and affine planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 90–97, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Park:1972:TDD

[276] Ralph Park. Topics in direct differential geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):98–148, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roberts:1972:IVB

[277] Leslie G. Roberts. Indecomposable vector bundles on the projective line. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):149–154, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cherenack:1972:BOA

[278] Paul Cherenack. Basic objects for an algebraic homotopy theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 155–166, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1972:ATE

[279] Ron Brown. An approximation theorem for extended prime spots. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 167–184, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wong:1972:GWL

[280] R. Wong and M. Wyman. A generalization of Watson’s lemma. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 185–208, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Day:1972:IEC

[281] Alan Day. Injectivity in equational classes of algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):209–220, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gruenberg:1972:ATC

[282] K. W. Gruenberg and J. E. Roseblade. The augmentation terminals of certain locally finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):221–238, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dobbs:1972:ITA

[283] David E. Dobbs. On inflation and torsion of Amitsur cohomology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 239–260, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bash:1972:SPN

[284] David W. Bash. Sums and products of normal functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):261–269, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ault:1972:EOP

[285] Janet E. Ault. Extensions of one primitive inverse semigroup by another. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):270–278, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rogers:1972:CHP

[286] James T. Rogers, Jr. The cone = hyperspace property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):279–285, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stephenson:1972:TCS

[287] R. M. Stephenson. Two R-closed spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):286–292, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hinton:1972:LPC

[288] Don Hinton. Limit point criteria for differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):293–305, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kutty:1972:SCP

[289] K. K. Kutty and J. Quinn. Some characterizations of the projection property in Archimedean Riesz spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 306–311, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Phadke:1972:ELM

[290] B. B. Phadke. Equidistant loci and the Minkowskian geometries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 312–327, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruns:1972:VOL

[291] Günter Bruns and Gudrun Kalmbach. Varieties of orthomodular lattices. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):328–337, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mack:1972:LTP

[292] John Mack, Marlon Rayburn, and Grant Woods. Local topological properties and one point extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 338–348, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorch:1972:HMP

[293] Lee Lorch, Martin E. Muldoon, and Peter Szego. Higher monotonicity properties of certain Sturm–Liouville functions. IV. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):349–368, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McDonald:1972:IMF

[294] B. R. McDonald. Involutory matrices over finite local rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 369–378, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hager:1972:UP

[295] Anthony W. Hager. Uniformities on a product. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):379–389, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuttner:1972:SNS

[296] Brian Kuttner and Brian Thorpe. On strong Nörlund summability fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):390–399, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1972:IHD

[297] William C. Brown and Wei eihn Kuan. Ideals and higher derivations in commutative rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):400–415, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See addendum [959].

Wong:1972:FVC

[298] Cécile M. Wong. Familles de valuations à caractère compact. (French) [Valuation families with compact character]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):416–425, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hodgson:1972:GCP

[299] J. P. E. Hodgson. A generalization of “Concordance of PL-homeomorphisms of §^{p} × §^{q}”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):426–431, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldrich:1972:CAI

[300] Fredric E. Goldrich and Eugene P. Wigner. Condition that all irreducible representations of a compact Lie group, if restricted to a subgroup, contain no representation more than once. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 432–438, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feustel:1972:MS

[301] C. D. Feustel. S-maximal subgroups of π_{1}(M^{3}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):439–449, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Teply:1972:TSC

[302] Mark L. Teply. The torsion submodule of a cyclic module splits off. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 450–464, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Strutt:1972:PHC

[303] Joseph Strutt. Projective homotopy classes of Stiefel manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 465–476, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bachelis:1972:MUC

[304] Gregory F. Bachelis and Louis Pigno. On multipliers with unconditionally converging Fourier series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 477–484, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Manes:1972:DS

[305] E. G. Manes. C(X) as a dual space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):485–491, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sobczyk:1972:ECP

[306] Andrew Sobczyk and James O. Gettys, Jr. Extended chromatic polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):492–501, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Williams:1972:IPI

[307] R. Douglas Williams. Intersections of primary ideals in rings of continuous functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):502–519, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Choi:1972:PLM

[308] Man Duen Choi. Positive linear maps on C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):520–529, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Poguntke:1972:UPT

[309] Detlev Poguntke. A universal property of the Takahashi quasi-dual. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 530–536, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berg:1972:NCS

[310] I. D. Berg. A note on complementary subspaces in c_{0}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):537–540, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Plunkett:1972:PF

[311] S. P. O. Plunkett. On the plethysm of S-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):541–552, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shock:1972:CAR

[312] Robert C. Shock. Certain Artinian rings are Noetherian. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):553– 556, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levitz:1972:RFN

[313] Kathleen B. Levitz and Joe L. Mott. Rings with finite norm property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 557–565, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Propes:1972:RPD

[314] R. E. Propes. Radicals of PID’s and Dedekind domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):566–572, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wiegand:1972:GTE

[315] Sylvia Wiegand. Galois theory of essential extensions of modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):

573–579, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vasilescu:1972:NLA

[316] F.-H. Vasilescu. Normed Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):580–591, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maxwell:1972:RAI

[317] George Maxwell. Representation of algebras with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 592–597, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hofer:1972:RSC

[318] Robert D. Hofer. Restrictive semigroups of continuous functions on Odimensional spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):598–611, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moore:1972:RSO

[319] M. J. Moore. Riemann surfaces as orbit spaces of Fuchsian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 612–616, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Poole:1972:AGQ

[320] George D. Poole and James D. Reid. Abelian groups quasi-injective over their endomorphism rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 617–621, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Porter:1972:NDS

[321] Jack R. Porter and R. Grant Woods. Nowhere dense subsets of metric spaces with applications to Stone–Čech compactifications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):622–630, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schirmer:1972:CSS

[322] Helga Schirmer. δ-continuous selections of small multifunctions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 631–635, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ruckle:1972:ESS

[323] William H. Ruckle. The extent of the sequence space associated with a basis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):636–641, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sarason:1972:APC

[324] Donald Sarason. Approximation of piecewise continuous functions by quotients of bounded analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):642–657, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thompson:1972:ECP

[325] R. C. Thompson and S. Therianos. The eigenvalues of complementary principal submatrices of a positive definite matrix. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):658–667, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Winter:1972:SAS

[326] David J. Winter. Separability in an algebra with semi-linear homomophism. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):668–671, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Parrott:1972:CTS

[327] David Parrott. A characterization of the tits’ simple group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 672–685, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1972:RTR

[328] Marvin Marcus and William Robert Gordon. Rational tensor representations of Hom(V, V) and an extension of an inequality of I. Schur. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 686–695, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coughlin:1972:SDA

[329] Raymond Coughlin and Michael Rich. On scalar dependent algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 696–702, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heinicke:1972:RQP

[330] A. G. Heinicke. On the ring of quotients at a prime ideal of a right noetherian ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(4):703–712, August 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lowenthal:1972:UFG

[331] Franklin Lowenthal. Uniform finite generation of SU(2) and SL(2, R). Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):713–727, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kleinfeld:1972:GCA

[332] Erwin Kleinfeld. A generalization of commutative and alternative rings. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):728–733, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herstein:1972:NDA

[333] I. N. Herstein and A. Ramer. A note on division algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):734–736, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Knight:1972:ACS

[334] William J. Knight. Absolute continuity of some vector functions and measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):737–746, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cahen:1972:PVE

[335] Paul-Jean Cahen. Polynomes à valeurs entières. (French) [Polynomials with integer values]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):747–754, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Linden:1972:IMZ

[336] C. N. Linden. Integral means and zero distributions of Blaschke products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):755–760, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanQue:1972:GFR

[337] Ngô van Quê. Généralisation de la formule de Riemann–Hurwitz. (French) [Generalization of the Riemann– Hurwitz formula]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):761–767, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boyd:1972:TWS

[338] W. S. Boyd and A. H. Wright. Taming wild simple closed curves with monotone maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):768–788, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoo:1972:CMS

[339] C. S. Hoo. Cyclic maps from suspensions to suspensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):789–791, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilmer:1972:EI

[340] Robert Gilmer and Anne Grams. The equality (AcapB)^{n} = A^{n}capB^{n} for ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):792–798, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bishop:1972:CFS

[341] R. L. Bishop and S. I. Goldberg. On conformally flat spaces with commuting curvature and Ricci transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):799–804, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kronk:1972:DGB

[342] Hudson V. Kronk and John Mitchem. On Dirac’s generalization of Brooks’ theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):805–807, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuntz:1972:APD

[343] Richard A. Kuntz. Associated prime divisors in the sense of Krull. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 808–818, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Naylor:1972:CT

[344] C. M. Naylor. On the β-construction in K-theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(5):819–824, October 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zenor:1972:CSP

[345] Phillip Zenor. Certain subsets of products of metacompact spaces and subparacompact spaces are realcompact. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):825–829, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adler:1972:DIU

[346] Andrew Adler and Murray Jorgensen. Descendingly incomplete ultrafilters and the cardinality of ultrapowers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):830–834, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Armendariz:1972:MQR

[347] E. P. Armendariz and Gary R. Mc-Donald. Maximal quotient rings and S-rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):835–850, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1972:NGL

[348] I. M. Isaacs. The number of generators of a linear p-group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):851–858, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deddens:1972:ATC

[349] James A. Deddens. Analytic Toeplitz and composition operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 859–865, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Geramita:1972:PRO

[350] A. V. Geramita. Polymonial rings with the outer product property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 866–872, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mosiman:1972:STC

[351] Steven E. Mosiman and Robert F. Wheeler. The strict topology in a completely regular setting: relations to topological measure theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 873–890, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rhemtulla:1972:ROG

[352] A. H. Rhemtulla. Right-ordered groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):891–895, ????

1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

James:1972:SRB

[353] Robert C. James. Super-reflexive Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):896–904, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kodama:1972:RFD

[354] Tetsuo Kodama. Residue free differentials and the Cartier operator for algebraic function fields of one variable. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):905–914, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Strichartz:1972:HSM

[355] Robert S. Strichartz. The Hardy space H^{1} on manifolds and submanifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):915–925, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stewart:1972:CCP

[356] James Stewart. Counterexample to a conjecture on positive definite functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):926–929, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Day:1972:RI

[357] Peter W. Day. Rearrangement inequalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):930–943, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Minassian:1972:ETO

[358] Donald P. Minassian. An embedding theorem for ordered groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 944–946, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wilson:1972:EHP

[359] Edward N. Wilson. Extensions of homomorphisms of parabolic subgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):947–956, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butler:1972:STS

[360] G. J. Butler and F. B. Richards. An L^{p} saturation theorem for splines. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 957–966, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yocom:1972:UFT

[361] K. L. Yocom. Unique factorization theorems for subalgebras of the incidence algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):967–977, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Digernes:1972:OPG

[362] Trond Digernes and Herbert Halpern. On open projections of GCR algebras.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):978–982, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beauregard:1972:RQR

[363] Raymond A. Beauregard. Right quotient rings of a right LCM domain. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):983–988, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beer:1972:CVF

[364] Gerald Beer. The continuity of the visibility function on a starshaped set. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):989–992, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nowitz:1972:GRRa

[365] Lewis A. Nowitz and Mark E. Watkins. Graphical regular representations of non-abelian groups. I, II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 993–1008, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nowitz:1972:GRRb

[366] Lewis A. Nowitz and Mark E. Watkins. Graphical regular representations of non-abelian groups. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1009– 1018, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leger:1972:CTB

[367] G. Leger and E. Luks. Cohomology theorems for borel-like solvable Lie algebras in arbitrary characteristic. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 1019–1026, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davey:1972:SL

[368] B. A. Davey. M-Stone lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 1027–1032, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pareek:1972:MSS

[369] C. M. Pareek. Moore spaces, semimetric spaces and continuous mappings connected with them. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 1033–1042, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hanson:1972:PLC

[370] T. H. McH. Hanson. Products of locally compact groups with zero and their actions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1043–1051, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cheng:1972:MSF

[371] Chao Kun Cheng. Multiplications on a space with finitely many non-vanishing homotopy groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1052–1062, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glass:1972:AUL

[372] A. M. W. Glass. An application of ultraproducts to lattice-ordered groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):1063–1064, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gould:1972:AGA

[373] Matthew Gould. Automorphism groups of algebras of finite type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 1065–1069, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Soare:1972:FMT

[374] Robert I. Soare. The Friedberg– Muchnik theorem re-examined. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 1070–1078, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

MacDonald:1972:DPR

[375] S. MacDonald and S. Willard. Domains of paracompactness and regularity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1079–1085, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Karrass:1972:TCL

[376] A. Karrass and D. Solitar. On a theorem of Cohen and Lyndon about free bases for normal subgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 1086–1091, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jockusch:1972:DWR

[377] Carl G. Jockusch, Jr. Degrees in which the recursive sets are uniformly recursive. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1092–1099, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marrero:1972:MHM

[378] O. Marrero and A. T. Butson. Modular Hadamard matrices and related designs. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1100–1109, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tucker:1972:SRU

[379] C. T. Tucker. Sequentially relatively uniformly complete Riesz spaces and Vulikh algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1110–1113, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Byers:1972:SPA

[380] W. Byers. Some properties of Anosov flows. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1114–1121, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Page:1972:PQR

[381] S. Page. Properties of quotient rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):1122–1128, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sunday:1972:PG

[382] J. G. Sunday. Presentations of the groups SL(2, m) and PSL(2, m). Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 1129–1131, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harris:1972:ECC

[383] Douglas Harris. Extension closed and cluster closed subspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 1132–1136, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hughes:1972:GRO

[384] Ian Hughes and Chou hsiang Wei. Group rings with only trivial units of finite order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1137–1138, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kohli:1972:NOE

[385] J. K. Kohli. A note on open extension of maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1139–1144, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haimo:1972:SED

[386] Deborah Tepper Haimo. Series expansions for dual Laguerre temperatures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):1145–1153, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chang:1972:HOD

[387] S. T. Chang. High order derivations and high order lie-like elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 1154–1163, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Traynor:1972:GHY

[388] Tim Traynor. A general Hewitt–Yosida decomposition. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1164–1169, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heinzer:1972:DFI

[389] William Heinzer and Jack Ohm. Defining families for integral domains of real finite character. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1170–1177, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ogiue:1972:KI

[390] Koichi Ogiue. On Kaehler immersions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24 (??):1178–1182, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nieto:1972:NRA

[391] José I. Nieto. Normed right alternative algebras over the reals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 1183–1186, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:1972:PMC

[392] Harold W. Martin. Product maps and countable paracompactness. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??): 1187–1190, ???? 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hsu:1972:DWT

[393] Hsing yuan Hsu. Distributional Watson transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(??):1191–1197, ???? 1972. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1972:RCF

[394] P. G. Rooney. On the ranges of certain fractional integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 24(6):1198– 1216, December 1972. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Currano:1973:FGI

[395] John J. Currano. Finite p-groups with isomorphic subgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1–13, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koliha:1973:ETA

[396] J. J. Koliha. Ergodic theory and averaging iterations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):14–23, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grant:1973:AMS

[397] Kerry E. Grant. Axioms for an nmetric structure. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):24–30, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cannon:1973:PNE

[398] J. W. Cannon. ULC properties in neighbourhoods of embedded surfaces and curves in E^{3}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):31–73, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osterweil:1973:TEC

[399] Leon Osterweil. A theorem for enumerating certain types of collections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):74–82, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grassmann:1973:EQD

[400] E. Grassmann. The existence of quadratic differentials in simply connected regions of the complex plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):83–91, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schneider:1973:I

[401] R. B. Schneider. Isometries of H^{p}(U^{n}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):92–95, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Woods:1973:CFS

[402] E. J. Woods. The classification of factors is not smooth. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):96–102, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tucker:1973:SPG

[403] Alan Tucker. The strong perfect graph conjecture for planar graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 103–114, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schick:1973:NGD

[404] Marlene Schick. A note on generalized direct products of groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 115–116, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kambayashi:1973:PCR

[405] T. Kambayashi. Plücker coordinates for regular chain groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 117–126, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bullen:1973:PD

[406] P. S. Bullen and S. N. Mukhopadhyay. On the Peano derivatives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 127–140, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1973:ADI

[407] W. C. Brown. The algebra of differentials of infinite rank. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):141–155, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramankutty:1973:UTS

[408] P. Ramankutty. Uniqueness theorems for a singular partial differential equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):156–163, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Petrich:1973:LTD

[409] Mario Petrich. L’enveloppe de translations d’un demi-treillis de groupes. (French) [The envelope of translations of a semilattice of groups]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 164–177, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sussmann:1973:GST

[410] Héctor J. Sussmann. Generalized spectral theory and second order ordinary differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):

178–193, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Galambos:1973:ILL

[411] J. Galambos. Integral limit laws for additive functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):194–203, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bell:1973:PLU

[412] Harold Bell and Gerald S. Ungar. From path lifting and unique arc lifting to unique path lifting. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):204–212, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Longyear:1973:NEC

[413] Judith Q. Longyear. Non-existence criteria for small configurations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 213–215, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Prasad:1973:UAS

[414] J. Prasad and H. Hayashi. On the uniform approximation of smooth functions by Jacobi polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 216–223, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hughes:1973:CTC

[415] J. W. B. Hughes. Corrigendum: “A theorem concerning partitions and its consequence in the theory of Lie algebras”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):224, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [2].

Cook:1973:CSS

[416] David E. Cook. Compactness and strong separation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):225–232, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mitchell:1973:CDD

[417] Barry Mitchell. The cohomological dimension of a directed set. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 233–238, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hell:1973:GMR

[418] Pavol Hell and Jaroslav Nesetril. Groups and monoids of regular graphs (and of graphs with bounded degrees). Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):239–251, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elliott:1973:UDF

[419] Joanne Elliott. On unsymmetric Dirichlet forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):252–260, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morash:1973:ABO

[420] Ronald P. Morash. Angle bisection and orthoautomorphisms in Hilbert lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):261–272, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haley:1973:ECA

[421] David K. Haley. Equationally compact Artinian rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):273–283, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Forelli:1973:TIA

[422] Frank Forelli. A theorem on isometries and the application of it to the isometries of H^{p}(S) for 2 < p < ∞. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):284–289, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hill:1973:SAS

[423] Paul Hill. Some almost simple rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):290–302, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boyd:1973:OPT

[424] David W. Boyd. The osculatory packing of a three dimensional sphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):303–322, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gustafson:1973:OCH

[425] G. B. Gustafson. Oscillation criteria for y′′ + p(t)f(y, y′) = 0 with f homogeneous of degree one. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):323–337, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

King:1973:ITG

[426] Larry King. R-isomorphisms of transformation groups and prolongations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):338–344, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ringel:1973:RGD

[427] Claus Michael Ringel. QF − 1 rings of global dimension leq2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 345–352, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Losey:1973:SFG

[428] Gerald Losey. On the structure of Q_{2}(G) for finitely generated groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):353–359, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coleman:1973:CGD

[429] Courtney Coleman. Corrigendum: “Growth and decay estimates near nonelementary stationary points”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):360, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [136].

Gaskill:1973:CSA

[430] H. S. Gaskill. Classes of semilattices associated with an equational class of lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):361–365, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Formanek:1973:ING

[431] Edward Formanek. Idempotents in Noetherian group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 366–369, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sichler:1973:TCS

[432] J. Sichler. Testing categories and strong universality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):370–385, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Siddiqi:1973:CLS

[433] Jamil A. Siddiqi. On the closure of the linear span of a weighted sequence in L^{p}(0, ∞). Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):386–396, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Biggs:1973:TRG

[434] Norman Biggs. Three remarkable graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):397–411, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aharonov:1973:CTC

[435] Dov Aharonov and W. E. Kirwan. Covering theorems for classes of univalent functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):412–419, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Campbell:1973:MST

[436] Douglas Michael Campbell. Majorizationsubordination theorems for locally univalent functions. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):420–425, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1973:CSRa

[437] W. T. Tutte. Chromatic sums for rooted planar triangulations: the cases lambda = 1 and lambda = 2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 426–447, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1973:CPE

[438] E. W. Johnson and J. A. Johnson. Corrigendum: “M-primary elements of a local Noether lattice”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):448, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [43].

Bahauddin:1973:HUS

[439] Mohammed Bahauddin and John Thomas. The homology of uniform spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):449–455, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leviatan:1973:CMP

[440] Talma Leviatan. Creation of mass processes and perturbation theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 456–467, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Potter:1973:NLP

[441] A. J. B. Potter. Non-linear A-proper mappings of the analytic type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 468–474, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edwards:1973:SEP

[442] C. S. Edwards. Some extremal properties of bipartite subgraphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 475–485, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Simpson:1973:CT

[443] William A. Simpson and J. Sutherland Frame. The character tables for SL(3, q), SU(3, q^{2}), PSL(3, q), PSU(3, q^{2}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):486–494, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dubuc:1973:SAA

[444] Serge Dubuc. Solutions asymptotiques au problème des secrétaires. (French) [Asymptotic solutions to the secretary problem]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):495–505, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Abian:1973:CPF

[445] Alexander Abian. The common points of families of normal functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 506–510, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bennett:1973:ITS

[446] G. Bennett and N. J. Kalton. Inclusion theorems for K-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 511–524, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Szeto:1973:WT

[447] George Szeto. On the Wedderburn theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):525–530, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jonker:1973:LCH

[448] L. B. Jonker and R. D. Norman. Locally convex hypersurfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 531–538, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hayden:1973:CFS

[449] John L. Hayden. A characterization of the finite simple group PSp_{4}(3). Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 539–553, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Powers:1973:FDP

[450] David L. Powers. The Frechet differential of a primary matrix function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):554–559, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moore:1973:AGW

[451] J. Douglas Moore and Edwin J. Hewett. Abelian groups in which every α-pure subgroup is β-pure. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(3): 560–566, June 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Walsh:1973:D

[452] T. Walsh. The dual of H^{p}(R_{+}^{n+1}) for p < 1. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):567–577, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davis:1973:QIL

[453] Nancy Davis and J. R. Kinney. Quadratic irrationals in the lower Lagrange spectrum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):578–584, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Klarner:1973:PIU

[454] D. A. Klarner and R. L. Rivest. A procedure for improving the upper bound for the number of n-ominoes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 585–602, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Henson:1973:EPP

[455] C. Ward Henson. Edge partition properties of graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):603–610, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yoshida:1973:GNT

[456] Hidenobu Yoshida. Generalizations of Noshiro’s theorem and their applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):611–620, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bridges:1973:CNM

[457] W. G. Bridges. A class of normal (0, 1)-matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):621–626, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steinberg:1973:QRC

[458] Stuart A. Steinberg. Quotient rings of a class of lattice-ordered-rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 627–645, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boyd:1973:SCC

[459] W. S. Boyd and A. H. Wright. A 1 − ALG simple closed curve in E^{3} is tame. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):646–656, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1973:CSRb

[460] W. T. Tutte. Chromatic sums for rooted planar triangulations. II. the case lambda = au + 1. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 657–671, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Armacost:1973:RMP

[461] D. L. Armacost. Remark on my paper: “Generators of monothetic groups”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(3):672, June 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [234].

Lindberg:1973:IEC

[462] John A. Lindberg, Jr. Integral extensions of commutative Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):673–686, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leonard:1973:GW

[463] John L. Leonard. Graphs with 6-ways. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):687–692, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1973:RWE

[464] S. H. Brown. Rings over which every simple module is rationally complete. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):693–701, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lipschutz:1973:NRD

[465] Seymour Lipschutz and Martin Lipschutz. A note on root decision problems in groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):702–705, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Potoczny:1973:STC

[466] Henry Potoczny. Sum theorems for countably paracompact spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 706–711, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beachy:1973:MTR

[467] John A. Beachy. On maximal torsion radicals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):712–726, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bochnak:1973:RFD

[468] Jacek Bochnak and Tzee-Char Kuo. Rigid and finitely V -determined germs of C^{∞}-mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):727–732, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dubois:1973:GEC

[469] Jacques Dubois. Génératrices extrémales d’un cône de fonctionnelles linéaires positives invariantes. (French) [Extremal generators of a cone of linear positive invariant functionals.]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 733–747, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Passi:1973:LSG

[470] I. B. S. Passi, D. S. Passman, and S. K. Sehgal. Lie solvable group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 748–757, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Little:1973:KTA

[471] Charles H. C. Little. Kasteleyn’s theorem and arbitrary graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 758–764, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rummler:1973:PTV

[472] Hansklaus Rummler. Parallel translation in vector bundles with abelian structure group and the Gauss– Bonnet formula. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):765–771, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lamb:1973:RLT

[473] Charles W. Lamb. A ratio limit theorem for approximate martingales. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):772–779, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1973:CSRc

[474] W. T. Tutte. Chromatic sums for rooted planar triangulations. III. the case lambda = 3. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):780–790, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ball:1973:ECL

[475] B. J. Ball and R. B. Sher. Embedding circle-like continua in E^{3}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 791–805, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Belley:1973:SPI

[476] Jean-Marc Belley. Spectral properties for invertible measure preserving transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):806–811, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haddad:1973:BBQ

[477] David C. Haddad. Boundary behavior and quasi-normality of finitely valent holomorphic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 812–819, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wiegmann:1973:IQN

[478] N. A. Wiegmann. Infinite quasinormal matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):820–828, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodearl:1973:PIR

[479] K. R. Goodearl. Prime ideals in regular self-injective rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):829–839, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:1973:MS

[480] Harold W. Martin. Metrizability of M-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):840–841, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wehlen:1973:ADD

[481] Joseph A. Wehlen. Algebras over Dedekind domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):842–855, ????

1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heinzer:1973:ERW

[482] William Heinzer and Jack Ohm. An essential ring which is not a vmultiplication ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):856–861, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Makan:1973:CSL

[483] A. R. Makan. On certain sublattices of the lattice of subgroups generated by the Prefrattini subgroups, the injectors and the formation subgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):862–869, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heilbronn:1973:RZD

[484] H. Heilbronn. On real zeros of Dedekind zeta-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 870–873, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chang:1973:SCJ

[485] S. H. Chang and Y. C. Lu. On C^{0}- sufficiency of complex jets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 874–880, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elgethun:1973:SCS

[486] E. D. Elgethun. Subgroups of central separable algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):881–887, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Best:1973:SOR

[487] L. A. Best. Subgroups of one-relator Fuchsian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):888–891, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shrikhande:1973:HCM

[488] M. S. Shrikhande. On hereditary and cohereditary modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 892–896, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Phillips:1973:TTS

[489] R. Phillips. A Tauberian theorem for a scale of logarithmic methods of summation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):897–902, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stone:1973:SRN

[490] A. P. Stone. Some remarks on the Nijenhuis tensor. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):903–907, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Larmore:1973:C

[491] Lawrence L. Larmore. The cohomology of (L^{2}X, DX). Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):908–921, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goursaud:1973:AGS

[492] J.-M. Goursaud. Sur les anneaux de groupes semi-parfaits. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):922– 928, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1973:CSRd

[493] W. T. Tutte. Chromatic sums for rooted planar triangulations. IV. the case λ = ∞. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):929–940, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Abramsky:1973:FSG

[494] Y. J. Abramsky, H. A. Jahn, and R. C. King. Frobenius symbols and the groups S_{s}, GL(n), O(n) and Sp(n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):941–959, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Simis:1973:RFC

[495] Aron Simis. Residual finiteness of commutative rings and schemes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):

960–972, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ruckle:1973:SWS

[496] William H. Ruckle. F K spaces in which the sequence of coordinate vectors is bounded. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):973–978, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sharma:1973:NMU

[497] P. L. Sharma. Non-metrizable uniformities and proximities on metrizable spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):979–981, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Campbell:1973:RCL

[498] Douglas Michael Campbell. The radius of convexity of a linear combination of functions in K, CV_{k}(η), S or U_{α}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):982–985, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Calder:1973:CIB

[499] J. R. Calder, W. P. Coleman, and R. L. Harris. Centers of infinite bounded sets in a normed space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):986–999, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reyner:1973:CMI

[500] Steven W. Reyner. Corrigendum: “Metacyclic invariants of knots and links” (Canad. J. Math. 22 (1970), 193–201) by R. H. Fox. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1000– 1001, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [28].

Cheatham:1973:DST

[501] Thomas Cheatham. Direct sums of torsion-free covers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1002–1005, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Solazzi:1973:IBC

[502] Robert Solazzi. On the isomorphisms between certain congruence groups. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):1006–1014, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Noonan:1973:CRC

[503] J. W. Noonan. Curvature and radius of curvature for functions with bounded boundary rotation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1015–1023, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Willett:1973:GCI

[504] D. Willett. A generalization of Caplygin’s inequality with applications to singular boundary value problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):1024–1039, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DeVore:1973:SWN

[505] R. DeVore, A. Meir, and A. Sharma. Strongly and weakly non-poised H − B interpolation problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 1040–1050, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1973:DFR

[506] G. de B. Robinson. The dual of Frobenius’ reciprocity theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 1051–1059, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruen:1973:TPF

[507] A. Bruen and B. Levinger. A theorem on permutations of a finite field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 1060–1065, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steinberger:1973:MPF

[508] Meïr Steinberger. On multiplicative properties of families of complexes of certain loops. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1066–1077, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Singh:1973:NSC

[509] Bhagat Singh. A necessary and sufficient condition for the oscillation of an even order nonlinear delay differential equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1078–1089, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1973:TSB

[510] P. G. Rooney. A technique for studying the boundedness and extendability of certain types of operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 1090–1102, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burns:1973:FGS

[511] R. G. Burns. Finitely generated subgroups of HNN groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 1103–1112, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roth:1973:DVC

[512] Richard L. Roth. A dual view of the Clifford theory of characters of finite groups. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1113–1119, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heil:1973:CMS

[513] Wolfgang Heil. Corrigendum: “On 3-manifolds with sufficiently large decompositions” (Canad. J. Math. 23 (1971), 746–748). Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(5):1120, October 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [229].

Sundaresan:1973:ONF

[514] K. Sundaresan and O. P. Kapoor. T - orthogonality and nonlinear functionals on topological vector spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 1121–1131, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dyer:1973:SHD

[515] Eldon Dyer and A. T. Vasquez. The sphericity of higher dimensional knots. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):1132–1136, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Das:1973:PFT

[516] Shawpawn Kumar Das. A partition of finite T_{0} topologies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1137–1147, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Das:1973:SFS

[517] Shawpawn Kumar Das. On the structure of finite T_{0} + T_{5} spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 1148–1158, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lemaire:1973:SMP

[518] Claude Lemaire. Sous-modules purs et sommantes directes. (French) [Pure submodules and direct sums]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 1159–1164, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tam:1973:CCA

[519] P. K. Tam. On the commutant of certain automorphism groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 1165–1169, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:1973:CFS

[520] Bang Yen Chen and Kentaro Yano. Conformally flat spaces of codimension 2 in a Euclidean space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 1170–1173, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Parmenter:1973:PIG

[521] M. M. Parmenter, I. B. S. Passi, and S. K. Sehgal. Polynomial ideals in group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1174–1182, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schnute:1973:KTB

[522] Jon Schnute and Marvin Shinbrot. Kinetic theory and boundary conditions for fluids. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1183–1215, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hooley:1973:DSF

[523] C. Hooley. On the distribution of square-free numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1216– 1223, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barr:1973:NAT

[524] Michael Barr. Non-abelian torsion theories. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1224–1237, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Promislow:1973:SMN

[525] David Promislow. Semi-metrics on the normal states of a W^{∗}-algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 1238–1253, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miers:1973:DRI

[526] C. Robert Miers. Derived ring isomorphisms of von Neumann algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):1254–1268, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Broughan:1973:NSR

[527] K. A. Broughan. On the number of structures of reflexive and transitive relations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):1269–1273, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bullen:1973:II

[528] P. S. Bullen and C. M. Lee. The SC_{n}P - integral and the P^{n+1}-integral. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??): 1274–1284, ???? 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Held:1973:SHT

[529] R. P. Held and D. Sjerve. On the stable homotopy type of Thom complexes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25 (??):1285–1294, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chollet:1973:CFA

[530] A.-M. Chollet. Sur des classes de fonctions analytiques dans le disque et indéfiniment dérivables à la frontière. (French) [On classes of analytic functions on the disk and infinitely differentiable at the border]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(6):1295– 1312, December 1973. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tall:1974:ENM

[531] Franklin D. Tall. On the existence of normal metacompact Moore spaces which are not metrizable. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):1–6, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chang:1974:WAG

[532] Bai Ching Chang. Which abelian groups can be fundamental groups of regions in Euclidean spaces? Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):7–18, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sherif:1974:ATC

[533] Soraya Sherif. Absolute Tauberian constants for Hausdorff transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):19–26, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muldowney:1974:IVP

[534] J. S. Muldowney and D. Willett. An intermediate value property for operators with applications to integral and differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):27–41, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dauns:1974:PIS

[535] John Dauns. The primitive ideal space of a C^{∗}-algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):42–49, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beineke:1974:EDP

[536] L. W. Beineke and R. E. Pippert. Enumerating dissectible polyhedra by their automorphism groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):50–67, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hill:1974:SMA

[537] Paul Hill. On the splitting of modules and abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):68–77, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Buckhiester:1974:RSM

[538] Philip G. Buckhiester. Rank r solutions to the matrix equation XAX^{T} = C, A nonalternate, C alternate, over GF(2^{y}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):78–90, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Farnum:1974:BST

[539] Nicholas R. Farnum. The Banach– Saks theorem in C(S). Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):91–97, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sheldon:1974:PIG

[540] Philip B. Sheldon. Prime ideals in GCD-domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):98–107, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dubois:1974:DTR

[541] D. Dubois and G. Efroymson. A dimension theorem for real primes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 108–114, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pearcy:1974:OCS

[542] Carl Pearcy and Norberto Salinas. Operators with compact self-commutator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):115–120, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Woods:1974:SRS

[543] S. M. Woods. Some results on semiperfect group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):121–129, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Montgomery:1974:RIW

[544] Susan Montgomery. Rings with involution in which every trace is nilpotent or regular. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):130–137, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schultz:1974:GTC

[545] Reinhard Schultz. A generalization of Thom classes and characteristic classes to nonspherical fibrations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):

138–144, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wells:1974:DCM

[546] Robert Wells. Double covers and metastable immersions of spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):145–176, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dawson:1974:GPR

[547] Jeffrey Dawson and David E. Dobbs. On going down in polynomial rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):177–184, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elliott:1974:SAO

[548] George A. Elliott. Some C^{∗}-algebras with outer derivations. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 185–189, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gupta:1974:CV

[549] Narain Gupta, Frank Levin, and Akbar Rhemtulla. Chains of varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 190–206, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shah:1974:EFS

[550] S. M. Shah and S. Y. Trimble. Entire functions with some derivatives univalent. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):207–213, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Karrass:1974:IST

[551] A. Karrass, A. Pietrowski, and D. Solitar. An improved subgroup theorem for HNN groups with some applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):214–224, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lindner:1974:CST

[552] Charles C. Lindner. Construction of Steiner triple systems having exactly one triple in common. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 225–232, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Freud:1974:SGS

[553] G. Freud and A. Sharma. Some good sequences of interpolatory polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):233–246, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1974:CPO

[554] Joel Anderson. On compact perturbations of operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):247–250, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Colwell:1974:IVF

[555] Peter Colwell. Improved versions of forms of Plessner’s theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 251–255, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beachy:1974:CBC

[556] John A. Beachy. Corrigendum: “Bicommutators of cofaithful, fully divisible modules” (Canad. J. Math. 23 (1971), 202–213). Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):256, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [167].

Chen:1974:CSG

[557] Yi Chen. A characterization of some geometries of chains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):257–272, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Oberle:1974:NEF

[558] Richard Alan Oberle. A note on equicontinuous families of volumes with an application to vector measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):273–280, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Oberle:1974:CCE

[559] Richard Alan Oberle. Characterization of a class of equicontinuous sets of finitely additive measures with an application to vector valued Borel measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):281–290, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mann:1974:HSH

[560] L. N. Mann. Highly symmetric homogeneous spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):291–293, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butler:1974:OCT

[561] G. Butler and J. W. Macki. Oscillation and comparison theorems for second order linear differential equations with integrable coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 294–301, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [759].

Shephard:1974:CPA

[562] G. C. Shephard. Combinatorial properties of associated zonotopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 302–321, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benabdallah:1974:PHS

[563] Khalid Benabdallah and John Irwin. Pure N-high subgroups, p-adic topology and direct sums of cyclic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):322–327, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butler:1974:SKP

[564] P. H. Butler and R. C. King. Symmetrized Kronecker products of group representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):328–339, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hinton:1974:LPC

[565] Don Hinton. Limit point criteria for differential equations. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 340–351, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dubois:1974:NVW

[566] Jacques Dubois. A note on the van der Waerden permanent conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 352–354, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edwards:1974:PES

[567] Gordon Edwards. Primitive elements in symmetric algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):355–364, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wasileski:1974:C

[568] J. S. Wasileski. Compactifications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):365–371, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duran:1974:ACS

[569] J. Peter Duran. Almost convergence, summability and ergodicity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 372–387, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thron:1974:FLP

[570] W. J. Thron and R. H. Warren. The family of Lodato proximities compatible with a given topological space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):388–404, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lund:1974:ISF

[571] Bruce Lund. Ideals and subalgebras of a function algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):405–411, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Griffin:1974:VPR

[572] Malcolm Griffin. Valuations and Prüfer rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):412–429, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Griffin:1974:MRT

[573] Malcolm Griffin. Multiplication rings via their total quotient rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 430–449, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dunkl:1974:OFA

[574] Charles F. Dunkl and Donald E. Ramirez. Operators on the Fourier algebra with weakly compact extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):450–454, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taylor:1974:SOS

[575] Peter D. Taylor. Some open sets for which the heat equation is simplicial. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):455–472, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dubrovsky:1974:SST

[576] Diana L. Dubrovsky. Some subfields of Q_{p} and their non-standard analogues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):473–491, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Williams:1974:NIT

[577] J. J. Williams. Non-isomorphic tensor products of von Neumann algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):492–512, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berger:1974:HHT

[578] T. R. Berger. Hall–Higman type theorems. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):513–531, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mott:1974:CDS

[579] Joe L. Mott. Convex directed subgroups of a group of divisibility. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 532–542, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burroughs:1974:OGR

[580] John Burroughs. Operations in Grothendieck rings and the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):543–550, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leach:1974:OFB

[581] Ronald J. Leach. On odd functions of bounded boundary rotation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 551–564, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Longstaff:1974:GNA

[582] W. E. Longstaff. Generators of nest algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):565–575, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Specht:1974:ISF

[583] William Specht. Isomorphic subgroups of finite p-groups revisited. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 576–579, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haley:1974:NCA

[584] David K. Haley. A note on compactifying Artinian rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):580–582, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:1974:IPW

[585] C. Y. Lee. Invariant polynomials of Weyl groups and applications to the centres of universal enveloping algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):583–592, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Karvellas:1974:ADC

[586] P. H. Karvellas. Additive divisibility in compact topological semirings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 593–599, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Griffiths:1974:PRD

[587] R. C. Griffiths. Permanents of random doubly stochastic matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):

600–607, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wilson:1974:OCW

[588] Jeremy Wilson. On outer-commutator words. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):608–620, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Halpern:1974:MBS

[589] Herbert Halpern. Mackey Borel structure for the quasi-dual of a separable C^{∗}-algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):621–628, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rummler:1974:GGC

[590] Hansklaus Rummler. On Gaussian and geodesic curvature of Riemannian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):629–635, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gallian:1974:FGH

[591] Joseph A. Gallian. Finite p-groups with homocyclic central factors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 636–643, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Redfield:1974:OUC

[592] R. H. Redfield. Ordering uniform completions of partially ordered sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 644–664, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Larman:1974:SCS

[593] D. G. Larman. On the selection of compact subsets of positive measure from analytic sets of positive measure. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 665–677, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lin:1974:FGN

[594] C. S. Lin. On a family of generalized numerical ranges. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):678–685, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butler:1974:GPS

[595] Jean W. Butler. A generation procedure for the simple 3-polytopes with cyclically 5-connected graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 686–708, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herrlich:1974:RCR

[596] Horst Herrlich. Regular categories and regular functors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):709–720, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ruckle:1974:CBS

[597] William H. Ruckle. On the classification of biorthogonal sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 721–733, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fixman:1974:PCP

[598] Uri Fixman and Frank A. Zorzitto. A purity criterion for pairs of linear transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):734–745, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Puttaswamaiah:1974:DBC

[599] B. M. Puttaswamaiah. Determination of Brauer characters. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 746–752, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nagel:1974:RD

[600] Alexander Nagel and Walter Rudin. Riesz decompositions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):753– 760, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ahmad:1974:IC

[601] Shafaat Ahmad. On implicational completeness. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):761–768, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dietze:1974:DSG

[602] Anke Dietze and Mary Schaps. Determining subgroups of a given finite index in a finitely presented group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 769–782, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bustoz:1974:RBS

[603] Joaquin Bustoz. On the relation between [f, g] and A_{λ} summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 783–793, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herstein:1974:RI

[604] I. N. Herstein. On rings with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):794–799, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Applebaum:1974:CTF

[605] C. H. Applebaum. Counterexamples in the theory of ω-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 800–805, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lebensold:1974:ECF

[606] Kenneth W. Lebensold. Extrema of a class of functions on a finite set. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):

806–819, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pelletier:1974:ERI

[607] Donald H. Pelletier. Easton’s results via iterated Boolean-valued extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):820–828, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zakon:1974:MCE

[608] Elias Zakon. Model-completeness and elementary properties of torsion free abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):829–840, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fontenot:1974:STV

[609] Robert A. Fontenot. Strict topologies for vector-valued functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 841–853, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [802].

Rush:1974:FM

[610] David E. Rush. The G-function of MacRae. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):854–865, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bleier:1974:MSH

[611] Roger D. Bleier. The sp-hull of a lattice-ordered group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):866– 878, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Toubassi:1974:GES

[612] E. H. Toubassi. The group of extensions and splitting length. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 879–883, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCandless:1974:ETN

[613] W. L. McCandless. Existence theorems for nonlinear boundary value problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):884–892, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1974:CSR

[614] W. T. Tutte. Chromatic sums for rooted planar triangulations. V. special equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):893–907, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marshall:1974:RTV

[615] Murray A. Marshall. Ramification theory for valuations of arbitrary rank. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):908–916, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marshall:1974:NMR

[616] Murray A. Marshall. The Nakayama map and ramification for maximally complete fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):917–919, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Woods:1974:ZDC

[617] R. Grant Woods. Zero-dimensional compactifications of locally compact spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):920–930, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allegretto:1974:OCQ

[618] W. Allegretto. Oscillation criteria for quasilinear equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):931– 947, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Milner:1974:SCE

[619] E. C. Milner and S. Shelah. Sufficiency conditions for the existence of transversals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):948–961, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Losey:1974:SFA

[620] Gerald Losey. N-series and filtrations of the augmentation ideal. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):

962–977, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rolfsen:1974:SCL

[621] Dale Rolfsen. Some counterexamples in link theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):978–984, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benedetto:1974:PME

[622] John J. Benedetto. Pseudo-measure energy and spectral synthesis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 985–1001, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fisher:1974:PLL

[623] R. K. Fisher. The polycyclic length of linear and finite polycyclic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):1002–1009, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sharma:1974:BD

[624] Sneh Sharma. A bound for the degree of H^{2}(G, Z_{p}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1010–1015, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Choudhary:1974:EMR

[625] B. Choudhary and Ashok Kumar. An extension of M. Riesz’s mean value theorem for infinite integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1016–1024, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zaks:1974:MGC

[626] Joseph Zaks. The maximum genus of Cartesian products of graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1025–1035, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carson:1974:ACR

[627] Andrew B. Carson. Algebraically closed regular rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1036–1049, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reilly:1974:THI

[628] N. R. Reilly. The translational hull of an inverse semigroup. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1050– 1068, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lambek:1974:LRN

[629] J. Lambek and G. Michler. Localization of right Noetherian rings at semiprime ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1069–1085, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Formanek:1974:TPR

[630] Edward Formanek. The type I part of the regular representation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1086–1089, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanZanten:1974:NCF

[631] A. J. van Zanten and E. de Vries. On the number of classes of a finite group invariant for certain substitutions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1090–1097, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lanski:1974:HRI

[632] Charles Lanski. Homomorphisms of rings with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1098– 1108, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bourgin:1974:GMD

[633] D. G. Bourgin. A generalization of the mapping degree. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1109–1117, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodaire:1974:IRA

[634] Edgar G. Goodaire. Irreducible representations of algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1118– 1129, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoel:1974:PRC

[635] C. Hoel. Pseudodifferential resolvent for a certain non-locally-solvable operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1130–1140, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muller:1974:SQF

[636] Bruno J. Müller. The structure of quasi-Frobenius rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1141– 1151, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gittings:1974:SRW

[637] Raymond F. Gittings. Some results on weak covering conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1152–1156, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Conrad:1974:AGR

[638] Paul Conrad. The additive group of an f-ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1157–1168, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Minda:1974:CEP

[639] Carl David Minda. Capacity estimates for planar cantor-like sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1169–1172, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Singh:1974:QIQ

[640] Surjeet Singh. Quasi-injective and quasi-projective modules over hereditary Noetherian prime rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1173–1185, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brungs:1974:RIR

[641] H. H. Brungs. Right invariant right hereditary rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1186–1191, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Myung:1974:CAC

[642] Hyo Chul Myung. A class of almost commutative nilalgebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1192–1198, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Montador:1974:SOC

[643] Bruce Montador. Le spectre des opérateurs de composition sur C[0, 1]. (French) [The spectrum of composition operators on C[0, 1]]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1199–1205, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baxter:1974:LET

[644] J. R. Baxter and R. V. Chacon. A local ergodic theorem on L_{p}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):

1206–1216, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blanchard:1974:SSC

[645] Andrew A. Blanchard. Structure species and constructive functors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):1217–1227, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schelter:1974:TRQ

[646] William Schelter. Topological rings of quotients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1228–1233, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cima:1974:AGC

[647] J. A. Cima and W. R. Wogen. On algebras generated by composition operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1234–1241, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ware:1974:NQF

[648] Roger Ware. A note on quadratic forms and the u-invariant. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1242–1244, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goebel:1974:ULF

[649] K. Goebel, W. A. Kirk, and R. L. Thele. Uniformly Lipschitzian families of transformations in Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):1245–1256, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kelly:1974:CFD

[650] David Kelly and Ivan Rival. Crowns, fences, and dismantlable lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):1257–1271, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Paschke:1974:DDI

[651] William L. Paschke. The double Bdual of an inner product module over a C^{∗}-algebra B. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1272–1280, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jung:1974:CFF

[652] Fuanglada R. Jung. On conics over a finite field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1281–1288, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gazik:1974:RRC

[653] R. J. Gazik. Regularity of Richardson’s compactification. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1289–1293, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shapiro:1974:WBT

[654] Joel H. Shapiro. On the weak basis theorem in F -spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1294–1300, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Platt:1974:ILL

[655] Craig Platt. Iterated limits of lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):1301–1320, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chong:1974:SET

[656] Kong Ming Chong. Some extensions of a theorem of Hardy, Littlewood and Pólya and their applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1321–1340, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feustel:1974:EEA

[657] C. D. Feustel. On embedding essential annuli in M^{3}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1341–1350, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leach:1974:CSF

[658] Ronald J. Leach. Coefficients of symmetric functions of bounded boundary rotation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1351–1355, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kurshan:1974:RGP

[659] R. P. Kurshan and B. Gopinath. Recursively generated periodic sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):1356–1371, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Radjabalipour:1974:GCD

[660] Mehdi Radjabalipour. Growth conditions and decomposable operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):1372–1379, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shores:1974:SCH

[661] Thomas S. Shores and Roger Wiegand. Some criteria for Hermite rings and elementary divisor rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1380–1383, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davis:1974:SLO

[662] Chandler Davis and Peter Rosenthal. Solving linear operator equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):1384–1389, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marsden:1974:KTI

[663] M. J. Marsden and S. D. Riemenschneider. Korovkin theorems for integral operators with kernels of finite oscillation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1390–1404, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rubin:1974:SSR

[664] Robert A. Rubin. Semi-simplicity relative to kernel functors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1405– 1411, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Phadke:1974:PT

[665] B. B. Phadke. Products of transvections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1412–1417, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gomory:1974:SS

[666] R. E. Gomory, T. C. Hu, and J. M. Yohe. R-separating sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1418–1429, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grabiner:1974:RPO

[667] Sandy Grabiner. Ranges of products of operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1430–1441, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kon:1974:SCS

[668] Masahiro Kon. On some complex submanifolds in Kaehler manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1442–1449, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Macedonska-Nosalska:1974:NAP

[669] O. N. Macedońska-Nosalska. A note on associative polyverbal operations on groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1450–1454, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mallory:1974:TPL

[670] Donald J. Mallory. Topological properties of limits of inverse systems of measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1455–1465, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Majindar:1974:ADS

[671] Kulendra N. Majindar. An arithmetical difference system with application to b. i. b. designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1466–1474, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reed:1974:CIM

[672] G. M. Reed. On continuous images of Moore spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1475–1479, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maxwell:1974:RNI

[673] J. W. Maxwell. Regular neighborhoods of immersed manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??): 1480–1483, ???? 1974. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Magill:1974:GRR

[674] K. D. Magill, Jr. and S. Subbiah. Green’s relations for regular elements of semigroups of endomorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):1484–1497, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weinstein:1974:LMG

[675] J. Weinstein. Large matchings in graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1498–1508, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arad:1974:CSS

[676] Zvi Arad. A characteristic subgroup of pi-stable groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1509–1514, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Scheinberg:1974:HIA

[677] Stephen Scheinberg. Homeomorphism and isomorphism of abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26 (??):1515–1519, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Petrich:1974:THS

[678] Mario Petrich. The translational hull of a semilattice of weakly reductive semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 26(??):1520–1536, ???? 1974. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stavrakas:1975:SCO

[679] Nick M. Stavrakas. L_{n} sets and the closures of open connected sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):1–5, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fieldhouse:1975:AP

[680] David J. Fieldhouse. Absolute purity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):6–10, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hager:1975:PME

[681] Anthony W. Hager. Perfect maps and epi-reflective hulls. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):11–24, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jonsson:1975:ADV

[682] B. Jónsson, G. McNulty, and R. Quackenbush. The ascending and descending varietal chains of a variety. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 25–31, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stevenson:1975:WIP

[683] Frederick W. Stevenson. Weakly isotopic planar ternary rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 32–36, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bottema:1975:CME

[684] Oene Bottema. Cardan motion in elliptic geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):37–43, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jackson:1975:MPC

[685] H. L. Jackson. On the mapping properties of certain exceptional sets in R^{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):44–49, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Day:1975:FMC

[686] Alan Day. Filter monads, continuous lattices and closure systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 50–59, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wulfsohn:1975:TPJ

[687] Aubrey Wulfsohn. Tensor products of Jordan algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):60–74, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adler:1975:TRR

[688] Andrew Adler and R. Douglas Williams. Transferring results from rings of continuous functions to rings of analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):75–87, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kalmbach:1975:SRM

[689] Gudrun Kalmbach. On some results in Morse theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):88–105, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dolan:1975:SON

[690] J. Michael Dolan and Gene A. Klaasen. Strongly oscillatory and nonoscillatory subspaces of linear equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 106–110, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McAdam:1975:GOE

[691] Stephen McAdam. Going down and open extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):111–114, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beachy:1975:MTR

[692] John A. Beachy. On maximal torsion radicals. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):115–120, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Limaye:1975:DCR

[693] B. V. Limaye and R. R. Simha. Deficiencies of certain real uniform algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):121–132, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schneider:1975:UPI

[694] Robert Schneider. Unit preserving isometries are homomorphisms in certain L^{p}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):133–137, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lewis:1975:OFO

[695] Roger T. Lewis. The oscillation of fourth order linear differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):138–145, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1975:CDD

[696] I. M. Isaacs. Character degrees and derived length of a solvable group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 146–151, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Azoff:1975:COR

[697] Edward A. Azoff. Compact operators in reductive algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 152–154, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Spiegel:1975:IAG

[698] Eugene Spiegel. On isomorphisms of abelian group algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):

155–161, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Andersen:1975:NIU

[699] Kenneth F. Andersen. Norm inequalities for ultraspherical and Hankel conjugate functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):162–171, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:1975:CLH

[700] Jongsik Kim. The characterization of a lattice homomorphism. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 172–175, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Welland:1975:NCR

[701] G. V. Welland. Norm convergence of Riesz–Bochner means for radial functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):176–185, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leach:1975:MMF

[702] Ronald J. Leach. Multivalent and meromorphic functions of bounded boundary rotation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):186–199, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Delver:1975:BIC

[703] Robert Delver. Boundary and interior control for partial differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):200–217, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Peterson:1975:OSI

[704] Roger D. Peterson. One-sided inverses in rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):218–224, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Czerwinski:1975:FLT

[705] Terry Czerwinski. On finite line transitive affine planes whose collineation groups contain no Baer involutions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):225–230, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blair:1975:GC

[706] David E. Blair. On a generalization of the catenoid. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):231–236, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brenner:1975:NFG

[707] J. L. Brenner, R. A. MacLeod, and D. D. Olesky. Non-free groups generated by two 2 × 2 matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 237–245, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sharp:1975:QAG

[708] W. T. Sharp, L. C. Biedenharn, E. de Vries, and A. J. van Zanten. On quasi-ambivalent groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 246–255, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lindner:1975:CLS

[709] C. C. Lindner and A. Rosa. Construction of large sets of almost disjoint Steiner triple systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 256–260, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rieffel:1975:IRR

[710] Marc A. Rieffel. Induced representations of rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):261–270, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cordes:1975:SRT

[711] Craig M. Cordes. Some results on totally isotropic subspaces and five-dimensional quadratic forms over GF(q). Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):271–275, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gratzer:1975:FDL

[712] G. Grätzer and J. Sichler. Free decompositions of a lattice. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):276–285, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bajpai:1975:AEF

[713] S. K. Bajpai and Joseph Tanne. Analogues of entire function inequalities for an analytic function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):286–293, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Strichartz:1975:EFD

[714] Robert S. Strichartz. The explicit Fourier decomposition of L^{2}(SO(n)/SO(n− m)). Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):294–310, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hong:1975:ESC

[715] Sung Sa Hong. Extensive subcategories of the category of T_{0}-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 311–318, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Toomer:1975:TLC

[716] Graham Hilton Toomer. Topological localization, category and cocategory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):319–322, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Toomer:1975:TAH

[717] Graham Hilton Toomer. Two applications of homology decompositions.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):323–329, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chong:1975:VRP

[718] Kong Ming Chong. Variation reducing properties of decreasing rearrangements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):330–336, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barnard:1975:VTB

[719] R. W. Barnard. A variational technique for bounded starlike functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):337–347, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vidyasankar:1975:MER

[720] K. Vidyasankar and D. H. Younger. A minimax equality related to the longest directed path in an acyclic graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):348–351, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hay:1975:RTD

[721] Louise Hay. Rice theorems for d.r.e. sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):352–365, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lovelady:1975:PBS

[722] David Lowell Lovelady. Partially bounded solutions of linear ordinary differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):366– 371, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Magid:1975:U

[723] Andy R. Magid. Ultrafunctors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 372–375, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Breen:1975:PLN

[724] Marilyn Breen. Points of local nonconvexity and finite unions of convex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):376–383, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Powell:1975:GPG

[725] Robert E. Powell and Richard A. Shoop. The Gibbs phenomenon for generalized Taylor and Euler transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):384–395, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koch:1975:UFG

[726] R. M. Koch and Franklin Lowenthal. Uniform finite generation of threedimensional linear Lie groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):

396–417, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1975:CPN

[727] C. L. Anderson, W. H. Hyams, and C. K. McKnight. Center points of nets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):418–422, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koh:1975:DDH

[728] E. L. Koh. The n-dimensional distributional Hankel transformation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 423–433, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tyler:1975:LCF

[729] Robert Tyler. On the lower central factors of a free associative ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 434–438, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Webster:1975:HSD

[730] Dallas E. Webster. Handles, semihandles, and destabilization of isotopies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):439–445, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hahn:1975:HMH

[731] Kyong T. Hahn. Holomorphic mappings of the hyperbolic space into the complex Euclidean space and the Bloch theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):446–458, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hodel:1975:TAC

[732] R. E. Hodel. On a theorem of Arhangel’skii concerning Lindelöf pspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):459–468, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bennett:1975:NCS

[733] Harold R. Bennett and Harold W. Martin. A note on certain spaces with bases (modK). Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):469–474, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shirey:1975:CSV

[734] James Shirey. The convergence of series for various choices of sign in Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):475–480, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCandless:1975:FF

[735] Byron H. McCandless. On Σ-finite families. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):481–488, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lawrence:1975:CRP

[736] John Lawrence. The coefficient ring of a primitive group ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 489–494, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dash:1975:BHP

[737] A. T. Dash. On a Brown–Halmos– Pearcy lemma (“Commutators of operators on Hilbert space”, Canad. J. Math. 17 (1965), 695–708, by A. Brown, P. R. Halmos and C. Pearcy). Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(3):495, June 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osofsky:1975:RC

[738] Barbara L. Osofsky. A remark on colimits. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):496–499, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lohoue:1975:CSE

[739] N. Lohoué. Sur certains sous-espaces de B_{p}(G). Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):500–507, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gustafson:1975:SSD

[740] G. B. Gustafson and S. Sedziwy. Solution space decompositions for nth order linear differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):

508–512, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mandelberg:1975:NQF

[741] K. I. Mandelberg. A note on quadratic forms over arbitrary semi-local rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):513–527, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanZanten:1975:CGR

[742] A. J. van Zanten and E. de Vries. Criteria for groups with representations of the second kind and for simple phase groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):528–544, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Viola-Prioli:1975:WEK

[743] Jorge E. Viola-Prioli. When is every kernel functor idempotent? Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 545–554, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Spence:1975:SHM

[744] Edward Spence. Skew-Hadamard matrices of the Goethals–Seidel type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):555–560, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wei:1975:LTP

[745] Albert Wei. Linear transformations preserving the real orthogonal group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):561–572, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rowen:1975:SRI

[746] Louis Halle Rowen. Structure of rings with involution applied to generalized polynomial identities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 573–584, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stahl:1975:RNW

[747] Saul Stahl. On the Ramsey number r(F, K_{m}) where F is a forest. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 585–589, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hutchinson:1975:EGP

[748] Joan P. Hutchinson. Eulerian graphs and polynomial identities for skewsymmetric matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):590–609, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jonker:1975:ISD

[749] Leo B. Jonker. Immersions with semidefinite second fundamental forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):610–617, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McMaster:1975:CTC

[750] R. J. McMaster. Cotorsion theories and colocalization. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):618–628, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herstein:1975:CSS

[751] I. N. Herstein. Certain submodules of simple rings with involution. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 629–635, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kelly:1975:PL

[752] David Kelly and Ivan Rival. Planar lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):636–665, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aliprantis:1975:OPO

[753] Charalambos D. Aliprantis. On order properties of order bounded transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):666–678, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Divinsky:1975:UR

[754] N. Divinsky. Unequivocal rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):

679–690, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McClure:1975:ISL

[755] J. P. McClure and R. Wong. On infinite systems of linear differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):691–703, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

White:1975:RTM

[756] Dennis E. White. Redfield’s theorems and multilinear algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 704–714, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pierce:1975:LOP

[757] Stephen Pierce. Linear operators preserving the real symplectic group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):715–724, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Radjabalipour:1975:DCP

[758] M. Radjabalipour and H. Radjavi. On decomposability of compact perturbations of normal operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 725–735, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butler:1975:COC

[759] G. Butler and J. W. Macki. Corrigendum: “Oscillation and comparison theorems for second order linear differential equations with integrable coefficients” (Canad. J. Math. 26 (1974), 294–301). Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):736, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [561].

Clough:1975:MDD

[760] Robert R. Clough. The map SJightarrowSF does not deloop mod2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):737–745, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Froemke:1975:MPP

[761] Jon Froemke. Maximal pre-primal clusters. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):746–751, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cheng:1975:ECM

[762] Chao Kun Cheng. H-equivalence classes of multiplications on certain fiber spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):752–765, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Patterson:1975:CP

[763] Leif-Norman Patterson. Connexions and prolongations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):766–791, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lohoue:1975:SHS

[764] N. Lohoué. Synthèse harmonique des sous-groupes discrets. (French) [Harmonic synthesis of discrete subgroups]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):792–796, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Myers:1975:HSD

[765] Donald Myers. Homeomorphisms on the solid double torus. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):797– 804, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Laird:1975:EMP

[766] P. G. Laird. Entire mean periodic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):805–818, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deza:1975:RMG

[767] Mikhail Deza. Racine minimum d’un groupe élémentaire abélien. (French) [Minimum root of an elementary abelian group]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):819–826, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elliott:1975:ADH

[768] George A. Elliott. An abstract Dauns– Hofmann–Kaplansky multiplier theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):827–836, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tomkinson:1975:PSC

[769] M. J. Tomkinson. Prefrattini subgroups and cover-avoidance properties in U-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):837–851, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Balbes:1975:TCS

[770] Raymond Balbes. On the triple characterization for Stone algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 852–859, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lutzer:1975:SCA

[771] D. J. Lutzer. Some constructions in abstract measure theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 860–866, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Singh:1975:MHN

[772] Surjeet Singh. Modules over hereditary Noetherian prime rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):

867–883, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burgess:1975:COC

[773] W. D. Burgess and R. Raphael. Complete and orthogonally complete rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):884–892, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morales:1975:NHM

[774] Pedro Morales. A non-Hausdorff multifunction Ascoli theorem for k_{3}-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):893–900, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Varadarajan:1975:NIHa

[775] K. Varadarajan. Numerical invariants in homotopical algebra. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 901–934, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Varadarajan:1975:NIHb

[776] K. Varadarajan. Numerical invariants in homotopical algebra. II. applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):935–960, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chimenti:1975:CFS

[777] Frank A. Chimenti. Convergence formulas for sequences of sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 961–969, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ho:1975:SEG

[778] C. Y. Ho. Some explicit generators for SL(3, 3^{n}), SU(3, 3^{n}), Sp(4, 3^{n}) and SL(4, 3^{n}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):970–979, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kenny:1975:CAG

[779] G. Otis Kenny. The completion of an abelian l-group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):980–985, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harley:1975:MSS

[780] P. W. Harley III and G. D. Faulkner. Metrization of symmetric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):986–990, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weaver:1975:RCO

[781] Robert J. Weaver. Refinement conditions on operations in sample spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):991–999, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kalmanson:1975:ECT

[782] Kenneth Kalmanson. Edgeconvex circuits and the traveling salesman problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1000–1010, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1975:CSZ

[783] Gordon Brown. Cartan subalgebras of Zassenhaus algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1011– 1021, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Merris:1975:EDS

[784] Russell Merris. Equality of decomposable symmetrized tensors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1022–1024, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Puttaswamaiah:1975:BCG

[785] B. M. Puttaswamaiah. Brauer characters and Grothendiek rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1025–1028, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wong:1975:SCC

[786] Pak Ken Wong. The second conjugates of certain Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):

1029–1035, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gaskill:1975:STF

[787] H. S. Gaskill and C. R. Platt. Sharp transferability and finite sublattices of free lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1036–1041, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Halbhavi:1975:GSD

[788] I. G. Halbhavi and K. Varadarajan. Gottlieb sets and duality in homotopy theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1042–1055, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Teply:1975:GST

[789] Mark L. Teply. Generalizations of the simple torsion class and the splitting properties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1056–1074, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akcoglu:1975:PET

[790] M. A. Akcoglu. A pointwise ergodic theorem in L_{p}-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1075– 1082, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Richmond:1975:GAR

[791] Bruce Richmond. A general asymptotic result for partitions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1083–1091, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mayland:1975:RFC

[792] E. J. Mayland, Jr. The residual finiteness of the classical knot groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1092–1099, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wicke:1975:CSH

[793] H. H. Wicke and J. M. Worrell, Jr. A characterization of spaces having bases of countable order in terms of primitive bases. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1100–1109, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1975:ISF

[794] Paul M. Gauthier and Lee A. Rubel. Interpolation in separable Fréchet spaces with applications to spaces of analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1110–1113, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chacron:1975:SCC

[795] M. Chacron, I. N. Herstein, and S. Montgomery. Structure of a certain class of rings with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1114–1126, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Florence:1975:SIR

[796] J. B. Florence. On splitting an infinite recursively enumerable class. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1127–1140, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruen:1975:SSI

[797] A. Bruen. Subregular spreads and indicator sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1141–1148, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Husain:1975:PDF

[798] T. Husain and S. A. Warsi. Positive definite functions for the class L_{p}(G). Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):1149–1156, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Noonan:1975:CBC

[799] J. W. Noonan. Coefficient behavior of a class of meromorphic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1157–1165, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lindner:1975:STS

[800] C. C. Lindner and A. Rosa. Steiner triple systems having a prescribed number of triples in common. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):

1166–1175, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [1164].

Kreimer:1975:MVS

[801] H. F. Kreimer and N. Heerema. Modularity vs. separability for field extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1176–1182, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fontenot:1975:STV

[802] Robert A. Fontenot. “Strict topologies for vector-valued functions” (Canad. J. Math. 26 (1974), 841–853): corrigendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1183–1184, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [609].

Waldinger:1975:NPC

[803] Hermann V. Waldinger and Anthony M. Gaglione. On nilpotent products of cyclic groups reexamined by the commutator calculus. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1185–1210, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leise:1975:BSI

[804] James A. Leise. The Bott suspension and the intrinsic join. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1211–1221, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Querre:1975:ARF

[805] J. Querré. Sur les anneaux reflexifs. (French) [On reflexive rings]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1222–1228, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Abernethy:1975:CCC

[806] Kenneth C. Abernethy. Characterization of certain classes of spaces with G_{δ} points as open images of metric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):1229–1238, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Noussair:1975:OEE

[807] E. S. Noussair. Oscillation of elliptic equations in general domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1239–1245, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gaskill:1975:STL

[808] H. Gaskill, G. Grätzer, and C. R. Platt. Sharply transferable lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1246–1262, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lohman:1975:GCS

[809] Robert H. Lohman and Peter G. Casazza. A general construction of spaces of the type of R. C. James. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):1263–1270, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schnute:1975:EKT

[810] Jon Schnute. Entropy and kinetic theory for a confined gas. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1271–1315, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Huige:1975:SRS

[811] Gustavus E. Huige. The spectral resolution of some non-selfadjoint partial differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1316– 1322, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1975:EVA

[812] P. M. Gauthier. Équidistribution des valeurs d’une application holomorphe générique à valeurs dans l’espace projectif. (French) [Equidistribution of values of a generic holomorphic application in values in the projective space]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27 (??):1323–1328, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edmunds:1975:PCP

[813] Charles C. Edmunds. Products of commutators as products of squares. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1329–1335, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lelek:1975:PCM

[814] A. Lelek and E. D. Tymchatyn. Pseudo-confluent mappings and a classification of continua. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1336– 1348, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1975:RII

[815] G. de B. Robinson. Restricting and inducing on inner products of representations of finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1349–1354, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wehrfritz:1975:FRF

[816] B. A. F. Wehrfritz. Faithful representations of finitely generated metabelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1355–1360, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nel:1975:ISC

[817] L. D. Nel. Initially structured categories and Cartesian closedness. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1361–1377, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Breen:1975:MDS

[818] Marilyn Breen. An n + 1 member decomposition for sets whose lnc points form n convex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1378–1383, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Breen:1975:ICS

[819] Marilyn Breen. Intersections of mconvex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??):1384–1391, ???? 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ginsburg:1975:SRC

[820] John Ginsburg. Some results on the countable compactness and pseudocompactness of hyperspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(??): 1392–1399, ???? 1975. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Picardello:1975:RFS

[821] Massimo A. Picardello. Random Fourier series on compact noncommutative groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 27(6):1400–1407, December 1975. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nakagawa:1976:EKM

[822] Hisao Nakagawa. Einstein–Kaehler manifolds immersed in a complex projective space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1–8, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

David:1976:EPR

[823] John David. On the endomorphisms of a polynomial ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):9–18, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Longstaff:1976:ORO

[824] W. E. Longstaff. Operators of rank one in reflexive algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):19–23, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wells:1976:TK

[825] Robert Wells and Luiz A. Favaro. On a theorem of Kuiper. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):24–41, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Limaye:1976:BRB

[826] B. V. Limaye. Boundaries for real Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):42–49, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taylor:1976:PCQ

[827] Walter Taylor. Pure compactifications in quasi-primal varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 50–62, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ackerman:1976:IRM

[828] Neill H. Ackerman and C. C. Hsiung. Isometry of Riemannian manifolds to spheres. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):63–72, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Singh:1976:MHN

[829] Surjeet Singh. Modules over hereditary Noetherian prime rings. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 73–82, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Becker:1976:ALA

[830] Joseph Becker and Leonard Lipshitz. An application of logic to analysis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):83–91, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Assaf:1976:CCT

[831] David Assaf. A counterexample to a classification theorem of linearly stable polytopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):92–93, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jacobson:1976:FRW

[832] David Jacobson. Finite rings in which 1 is a sum of two non-p-th power units. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):94–103, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roth:1976:UTA

[833] Alice Roth. Uniform and tangential approximations by meromorphic functions on closed sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):104–111, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1976:PUR

[834] P. M. Gauthier, A. Roth, and J. L. Walsh. Possibility of uniform rational approximation in the spherical metric. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):112–115, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1976:TPA

[835] Norman L. Johnson and Michael J. Kallaher. On translation planes with affine central collineations. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 116–129, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schoen:1976:FRF

[836] James A. Schoen. Factor representations and factor states on a C^{∗}-algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):130–134, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

delJunco:1976:CEP

[837] A. del Junco. Coding ergodic processes to approximate Bernoulli processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):135–140, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bellenot:1976:NHC

[838] Steven F. Bellenot. On nonstandard hulls of convex spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 141–147, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cornish:1976:WRA

[839] William H. Cornish and Patrick N. Stewart. Weakly regular algebras, Boolean orthogonalities and direct products of integral domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 148–153, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chong:1976:ITR

[840] Kong Ming Chong. An induction theorem for rearrangements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 154–160, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Montesinos:1976:MPR

[841] José M. Montesinos. Minimal plat representations of prime knots and links are not unique. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(1):161–167, February 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See erratum [977].

Cook:1976:FPP

[842] R. J. Cook. On the fractional parts of a polynomial. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):168–173, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berman:1976:DLA

[843] Stephen Berman. On derivations of Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):174–180, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chao:1976:SSL

[844] Chong Yun Chao and Ernest L. Stitzinger. Subinvariance in solvable Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):181–185, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glass:1976:CTR

[845] A. M. W. Glass. Compatible tight Riesz orders. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):186–200, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pettet:1976:TGA

[846] Martin R. Pettet. On a theorem of Goldschmidt applied to groups with a coprime automorphism. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 201–206, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rajagopalan:1976:SCI

[847] M. Rajagopalan and R. F. Wheeler. Sequential compactness of X implies a completeness property for C(X). Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):207–210, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sarath:1976:DDS

[848] B. Sarath and K. Varadarajan. Divisibility of direct sums in torsion theories. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):211–214, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Geramita:1976:ODZ

[849] A. V. Geramita and J. H. Verner. Orthogonal designs with zero diagonal. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (1):215–224, February 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mather:1976:PBH

[850] Michael Mather. Pull-backs in homotopy theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):225–263, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Birman:1976:SEP

[851] Joan S. Birman. On the stable equivalence of plat representations of knots and links. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):264–290, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sharma:1976:CIG

[852] A. Sharma and J. Tzimbalario. Cardinal interpolation and generalized exponential Euler splines. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):291– 300, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levikson:1976:NAO

[853] Benny Levikson. A new approximation operator generalizing Meyer– König and Zeller’s power series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 301–311, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Everitt:1976:NDC

[854] W. N. Everitt. A note on the Dirichlet condition for second-order differential expressions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):312–320, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Totten:1976:CRL

[855] Jim Totten. Classification of restricted linear spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):321–333, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koranyi:1976:ISB

[856] Adam Korányi and Stephen Vági. Isometries of H^{p} spaces of bounded symmetric domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):334–340, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Loveland:1976:CFA

[857] L. D. Loveland. Continuous finite Apollonius sets in metric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 341–347, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1976:CV

[858] Tim Anderson and Erwin Kleinfeld. A classification of 2-varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 348–364, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beauregard:1976:CAI

[859] Raymond A. Beauregard and David E. Dobbs. On a class of Archimedean integral domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):365–375, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Luneburg:1976:CGH

[860] Heinz Lüneburg. Characterizations of the generalized Hughes planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 376–402, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mucci:1976:RLT

[861] A. G. Mucci. Ratio limit theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):403–407, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fuelberth:1976:SSP

[862] John Fuelberth and James Kuzmanovich. Split subdirect products and piecewise domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):408–419, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hurley:1976:NES

[863] James F. Hurley. Normality in elementary subgroups of Chevalley groups over rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):420–428, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Callahan:1976:DFE

[864] T. Callahan. Dihedral field extensions of order 2p whose class numbers are multiples of p. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):429–439, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Usher:1976:PF

[865] A. O. Usher. Plethysm of S-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):440–445, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glasner:1976:CIS

[866] Moses Glasner. Canonically isomorphic spaces of bounded solutions of Δu = P u. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):446–448, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Warrack:1976:DPL

[867] Brian Warrack. Domains of paracompactness and local compactness. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 449–454, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ong:1976:LTM

[868] Hock Ong and E. P. Botta. Linear transformations on matrices: the invariance of generalized permutation matrices. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):455–472, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gildenhuys:1976:GLT

[869] D. Gildenhuys. A generalization of Lyndon’s theorem on the cohomology of one-relator groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):

473–480, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fine:1976:FSP

[870] Benjamin Fine. Fuchsian subgroups of the Picard group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):481–485, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1976:OTG

[871] R. C. Brown. The operator theory of generalized boundary value problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):486–512, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weinstock:1976:CBV

[872] Barnet M. Weinstock. Continuous boundary values of holomorphic functions on Kähler domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 513–522, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garman:1976:GST

[873] B. L. Garman, R. D. Ringeisen, and A. T. White. On the genus of strong tensor products of graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 523–532, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Orzech:1976:BGA

[874] Morris Orzech. On the Brauer group of algebras having a grading and an action. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):533–552, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [1312].

Broverman:1976:SCV

[875] S. Broverman. The structure of continuous 0, 1-valued functions on a topological product. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):553–559, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garrison:1976:DFS

[876] Sidney Garrison. Determining the Frattini subgroup from the character table. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):560–567, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mayland:1976:SPG

[877] E. J. Mayland, Jr. and Kunio Murasugi. On a structural property of the groups of alternating links. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 568–588, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:1976:PI

[878] John R. Martin. On prime immersions of S^{1} into R^{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):589–593, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ku:1976:DTT

[879] Hsü Tung Ku and Mei Chin Ku. Dimension of a topological transformation group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):594–599, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muller:1976:LNC

[880] Bruno J. Müller. Localization in noncommutative Noetherian rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 600–610, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Comfort:1976:MSL

[881] W. W. Comfort and Anthony W. Hager. Metric spaces without large closed discrete sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):611–626, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blevins:1976:CTU

[882] Donald K. Blevins. Covering theorems for univalent functions mapping onto domains bounded by quasiconformal circles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):627–631, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wage:1976:CHN

[883] Michael L. Wage. A collectionwise hausdorff, non-normal Moore space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):632–634, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sacerdote:1976:GBT

[884] George S. Sacerdote. On the groups of Britton’s theorem A. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 635–639, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tromba:1976:ARS

[885] A. J. Tromba. Almost-Riemannian structures on Banach manifolds: the Morse lemma and the Darboux theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(3):640–652, June 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bastien:1976:SFT

[886] Richard Bastien and Serge Dubuc. Systèmes faibles de Tchebycheff et polynômes de Bernstein. (French) [Weak Chebyshev systems and Bernstein polynomials ]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):653–658, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Magid:1976:FPC

[887] Andy R. Magid. Fundamental, picard, and class groups of rings of invariants.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):659–664, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thomas:1976:GCD

[888] Glânffrwd P. Thomas. A generalization of a construction due to Robinson. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):665–672, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blair:1976:SWS

[889] Robert L. Blair. Spaces in which special sets are z-embedded. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):673–690, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sabidussi:1976:WSL

[890] Gert Sabidussi. Weak separation lattices of graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):691–724, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Polat:1976:TSG

[891] Norbert Polat. Treillis de séparation des graphes. (French) [Separation lattices of graphs]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):725–752, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Siddiqi:1976:FYC

[892] Rafat N. Siddiqi. Fourier–Young coefficients of a function of Wiener’s class V_{p}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):753–759, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1976:MIM

[893] D. D. Anderson. Multiplication ideals, multiplication rings, and the ring R(X). Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):760–768, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DArchangelo:1976:VDB

[894] J. D’Archangelo and P. A. McCoy. Value distribution of biaxially symmetric harmonic polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 769–773, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fried:1976:SEF

[895] Michael Fried and Moshe Jarden. Stable extensions and fields with the global density property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):774–787, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Papick:1976:LMO

[896] Ira J. Papick. Local minimal overrings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):788–792, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1976:GII

[897] G. F. D. Duff. A general integral inequality for the derivative of an equimeasurable rearrangement. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 793–804, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nicholls:1976:TPF

[898] Peter J. Nicholls. Transitivity properties of Fuchsian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 805–814, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Houston:1976:LR

[899] Evan G. Houston. Localizations of H_{1} and D_{1} rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):815–819, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lewis:1976:O

[900] William J. Lewis and Jack Ohm. The ordering of Spec R. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):820–835, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

delJunco:1976:TDS

[901] Andrés del Junco. Transformations with discrete spectrum are stacking transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):836–839, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kusano:1976:NOF

[902] Takaŝi Kusano and Manabu Naito. Nonlinear oscillation of fourth order differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):840– 852, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Askey:1976:PPS

[903] Richard Askey and Mourad E. H. Ismail. Permutation problems and special functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):853–874, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pettey:1976:MRS

[904] Dix H. Pettey. A minimal regular space that is not strongly minimal regular. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):875–878, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thomas:1976:CIW

[905] Glânffrwd P. Thomas. A combinatorial interpretation of the wreath product of Schur functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):879–884, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1976:HFO

[906] P. M. Gauthier and L. A. Rubel. Holomorphic functionals on open Riemann surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):885–888, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zorzitto:1976:PCS

[907] Frank Zorzitto. Purity and copurity in systems of linear transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):889–896, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baker:1976:AOS

[908] John Warren Baker and John Wolfe. Averaging operators and C(X)-spaces with the separable projection property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):897–904, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1976:LPL

[909] Robert L. Anderson. Limit point and limit circle criteria for a class of singular symmetric differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):905–914, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lanski:1976:RIW

[910] Charles Lanski. Rings with involution whose symmetric units commute. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):915–928, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McKay:1976:MAM

[911] J. McKay. More about the Mathieu group M_{22}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):929–937, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Childress:1976:WDQ

[912] C. L. Childress. Weierstrass division in quasianalytic local rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 938–953, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Milies:1976:IGR

[913] César Polcino Milies. Integral group rings with nilpotent unit groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 954–960, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Freud:1976:APT

[914] G. Freud, A. Giroux, and Q. I. Rahman. Sur l’approximation polynomiale sur tout l’axe réel. (French) [On the polynomial approximation on the entire real axis]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):961–967, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brownawell:1976:AIC

[915] W. Dale Brownawell. The algebraic independence of certain exponential functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):968–976, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruns:1976:FO

[916] Günter Bruns. Free ortholattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):977–985, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bell:1976:CCR

[917] Howard E. Bell. A commutativity condition for rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):986–991, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoft:1976:SFP

[918] Hartmut Höft and Margret Höft. Some fixed point theorems for partially ordered sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):992–997, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Juhasz:1976:TMH

[919] I. Juhász, K. Kunen, and M. E. Rudin. Two more hereditarily separable non-Lindelöf spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):998–1005, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wolfe:1976:AOD

[920] Warren Wolfe. Amicable orthogonal designs—existence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1006–1020, ????

1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cooperstein:1976:SGA

[921] Bruce N. Cooperstein. Some geometries associated with parabolic representations of groups of Lie type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 1021–1031, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehman:1976:CET

[922] B. Lehman. Cyclic element theory in connected and locally connected Hausdorff spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1032–1050, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Breen:1976:DTC

[923] Marilyn Breen. A decomposition theorem for m-convex sets in R^{d}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 1051–1057, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1976:FCH

[924] D. Borwein and B. L. R. Shawyer. On functional Cesàro and Hölder methods of summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1058–1061, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Longstaff:1976:LTD

[925] W. E. Longstaff. A lattice-theoretic description of the lattice of hyperinvariant subspaces of a linear transformation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1062–1066, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Awtar:1976:JSS

[926] Ram Awtar. On Jordan structure in semiprime rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1067–1072, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

delaTorre:1976:SPM

[927] Alberto de la Torre. A simple proof of the maximal ergodic theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 1073–1075, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davis:1976:CTR

[928] Gary Davis and Colin D. Fox. Compatible tight Riesz orders on the group of automorphisms of an 0 − 2homogeneous set. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1076–1081, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kemp:1976:FDP

[929] R. R. D. Kemp and S. J. Lee. Finite dimensional perturbations of differential expressions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1082–1104, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nicholson:1976:SMR

[930] W. K. Nicholson. Semiregular modules and rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1105–1120, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Russo:1976:NFT

[931] Bernard Russo. The norm of the L^{p}- fourier transform. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1121–1131, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McClure:1976:ISD

[932] J. P. McClure and R. Wong. Infinite systems of differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 1132–1145, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lam:1976:FPT

[933] Ling Fai Lam. A fixed point theorem for positive operators on KB spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):1146–1152, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arkin:1976:SML

[934] Joseph Arkin, Verner E. Hoggatt, Jr., and E. G. Straus. Systems of magic Latin k-cubes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1153–1161, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deveney:1976:PSP

[935] James K. Deveney. Pure subfields of purely inseparable field extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):1162–1166, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Little:1976:NPS

[936] Charles H. C. Little. On the number of parity sets in a graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 1167–1171, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leighton:1976:SCT

[937] Walter Leighton. Some comparison theorems for conjugate and σ-points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):1172–1179, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Katavolos:1976:IMN

[938] A. Katavolos. Isometric mappings of non-commutative L_{p} spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 1180–1186, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Quackenbush:1976:VSL

[939] Robert W. Quackenbush. Varieties of Steiner loops and Steiner quasigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):1187–1198, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Djokovic:1976:IIC

[940] Dragomiz Z. Djoković and Jerry Malzan. Imprimitive, irreducible complex characters of the alternating group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1199–1204, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stahl:1976:HCP

[941] Stanley H. Stahl. A hierarchy on the class of primitive recursive ordinal functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1205–1209, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1976:AVA

[942] P. M. Gauthier and J. S. Hwang. Asymptotic values along Julia rays. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):1210–1215, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Longyear:1976:CHM

[943] Judith Q. Longyear. Criteria for a Hadamard matrix to be skewequivalent. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1216–1223, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

May:1976:SIT

[944] C. P. May. Saturation and inverse theorems for combinations of a class of exponential-type operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 1224–1250, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tannenbaum:1976:AST

[945] Peter Tannenbaum. Abelian Steiner triple systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1251–1268, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stevens:1976:EHP

[946] John G. Stevens. Elimination from homogeneous polynomials over a polynomial ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1269–1276, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Button:1976:NSC

[947] Robert Warren Button. A nonstandard characterization of perfect mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1277–1279, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vitek:1976:BLD

[948] Yehoshua Vitek. Bounds for a linear Diophantine problem of Frobenius. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28 (??):1280–1288, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1976:BSM

[949] William C. Brown. Blow up sequences and the module of nth order differentials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):1289–1301, ???? 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartley:1976:CBM

[950] Brian Hartley. A conjecture of Bachmuth and Mochizuki on automorphisms of soluble groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 1302–1310, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cummings:1976:LSC

[951] L. J. Cummings and R. W. Robinson. Linear symmetry classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 1311–1319, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Handelman:1976:RRC

[952] David Handelman. Representing rank complete continuous rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??): 1320–1331, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burkinshaw:1976:WSC

[953] Owen Burkinshaw and Peter Dodds. Weak sequential compactness and completeness in Riesz spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):

1332–1339, ???? 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bollobas:1976:UCG

[954] Béla Bollobás and Norbert Sauer. Uniquely colourable graphs with large girth. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(6):1340–1344, December 1976. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lane:1977:Q

[955] Norman D. Lane, Peter Scherk, and Jean M. Turgeon. Quasigraphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):1–28, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wagner:1977:SCR

[956] Daniel H. Wagner. Semi-compactness with respect to a Euclidean cone. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):29–36, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wasileski:1977:ABC

[957] J. S. Wasileski. On Alexandroff base compactifications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):37–44, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Walker:1977:HPB

[958] Marshall Walker. Homotopy pull-backs and applications to duality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 45–64, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuan:1977:AIH

[959] Wei Eihn Kuan. An addendum to: “Ideals and higher derivations in commutative rings” (Canad. J. Math. 24 (1972), 400–415) by W. C. Brown and Wei-eihn Kuan. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):65–70, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [297].

Choe:1977:ESU

[960] T. H. Choe and Y. H. Hong. Extensive subcategories in universal topological algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):71–76, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hirschorn:1977:CHS

[961] Ronald M. Hirschorn. Curves in homogeneous spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):77–83, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rudin:1977:SEN

[962] Mary Ellen Rudin and Michael Starbird. Some examples of normal Moore spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):84–92, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bernstein:1977:GUU

[963] Leon Bernstein. Gaining units from units. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):93–106, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benabdallah:1977:GQP

[964] Khalid Benabdallah and Robert Bradley. Sur les groupes quasi-purs-projectifs torsions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):107–110, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Suzuki:1977:HDH

[965] Shin’ichi Suzuki. On homeomorphisms of a 3-dimensional handlebody. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 111–124, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanMill:1977:EHC

[966] J. van Mill. Every Hausdorff compactification of a locally compact separable space is a GA compactification. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 125–131, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sherk:1977:ISR

[967] F. A. Sherk and Günther Pabst. Indicator sets, reguli, and a new class of spreads. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):132–154, ????

1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fine:1977:SSR

[968] Benjamin Fine. Sum of squares rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):155–160, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putnam:1977:PPU

[969] C. R. Putnam. Positive perturbations and unitary equivalence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 161–164, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steffens:1977:MCG

[970] K. Steffens. Matchings in countable graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):165–168, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fulton:1977:RHF

[971] John D. Fulton. Representations by Hermitian forms in a finite field of characteristic two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):169–179, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:1977:RAP

[972] Q. I. Rahman and G. Schmeisser. On rational approximation on the positive real axis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):180–192, ????

1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fischer:1977:MCE

[973] Arthur E. Fischer and Jerrold E. Marsden. The manifold of conformally equivalent metrics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):193–209, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Conway:1977:CAC

[974] John B. Conway. On the Calkin algebra and the covering homotopy property. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):210–215, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Khurana:1977:STP

[975] Surjit Singh Khurana. Strict topology on paracompact locally compact spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):216–219, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [1145].

Trushin:1977:NRT

[976] David Trushin. A note on the representation type of pointed irreducible coalgebras and unipotent algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 220–223, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Montesinos:1977:EMP

[977] José M. Montesinos. Erratum: “Minimal plat representations of prime knots and links are not unique” (Canad. J. Math. 28 (1976), no. 1, 161–167). Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):224, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [841].

Bruen:1977:SNR

[978] A. A. Bruen. Some new replaceable translation nets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):225–237, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beynon:1977:RSP

[979] W. M. Beynon. On rational subdivisions of polyhedra with rational vertices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):238–242, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beynon:1977:ADT

[980] W. M. Beynon. Applications of duality in the theory of finitely generated lattice-ordered abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 243–254, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Farzan:1977:KPL

[981] M. Farzan and D. A. Waller. Kronecker products and local joins of graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):255–269, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Husain:1977:MFF

[982] T. Husain and J. Liang. Multiplicative functionals on Fréchet algebras with bases. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):270–276, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodaire:1977:ADS

[983] E. G. Goodaire. Algebras with a diagonable subspace whose centralizer satisfies a polynomial identity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 277–283, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levy:1977:AS

[984] Ronnie Levy. Almost-P -spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 284–288, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dokovic:1977:ISS

[985] Dragomir Ž. – Doković. Isomorphism of some simple 2-graded Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):289–294, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dence:1977:LDV

[986] Thomas P. Dence. A Lebesgue decomposition for vector valued additive set functions defined on a lattice. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 295–298, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wicker:1977:GBM

[987] Fletcher D. Wicker. Generalized Bloch mappings in complex Hilbert space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):299–306, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beauregard:1977:ANT

[988] Raymond A. Beauregard. An analog of Nagata’s theorem for modular LCM domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):307–314, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jeffries:1977:WMS

[989] Clark Jeffries, Victor Klee, and Pauline van den Driessche. When is a matrix sign stable? Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):315–326, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Little:1977:TEC

[990] Robert D. Little. Torsion elements and the classification of vector bundles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):327–332, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Djokovic:1977:IAC

[991] D. Z. Djoković and J. Malzan. Irreducible automorphisms of certain p-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):333–348, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Apt:1977:REP

[992] K. R. Apt. Recursive embeddings of partial orderings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):349–359, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gruenhage:1977:NQM

[993] Gary Gruenhage. A note on quasimetrizability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):360–366, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kelly:1977:IPO

[994] David Kelly. The 3-irreducible partially ordered sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):367–383, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dixon:1977:RES

[995] John D. Dixon. Rigid embedding of simple groups in the general linear group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):384–391, ????

1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

StMary:1977:TOL

[996] Donald F. St. Mary. On transformation and oscillation of linear differential systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):392–399, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Altshuler:1977:NPN

[997] A. Altshuler. Neighborly 4-polytopes and neighborly combinatorial 3manifolds with ten vertices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 400–420, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arnold:1977:SPC

[998] James E. Arnold, Jr. On Steenrod’s problem for cyclic p-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 421–428, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bezak:1977:CPO

[999] Nicholas J. Bezak and Martin Eisen. Continuity properties of operator spectra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):429–437, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mahmoodian:1977:FG

[1000] Ebad Mahmoodian. On factors of a graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):438–440, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tzimbalario:1977:LCC

[1001] J. Tzimbalario. Lebesgue constants for cardinal L-spline interpolation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 441–448, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davey:1977:WIC

[1002] Brian A. Davey. Weak injectivity and congruence extension in congruencedistributive equational classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 449–459, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feintuch:1977:RCO

[1003] Avraham Feintuch. On reflexive compact operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):460–465, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McKenzie:1977:AGD

[1004] Ralph McKenzie. Automorphism groups of denumerable Boolean algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):466–471, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butler:1977:ECS

[1005] G. J. Butler. The existence of continuable solutions of a second order differential equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):472–479, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See erratum [1247].

Stephenson:1977:SWF

[1006] R. M. Stephenson, Jr. Symmetrizable, F-, and weakly first countable spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):480–488, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Diaconis:1977:ETI

[1007] Persi Diaconis. Examples for the theory of infinite iteration of summability methods. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):489–497, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taylor:1977:VOH

[1008] Walter Taylor. Varieties obeying homotopy laws. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):498–527, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clarke:1977:ICC

[1009] Frank H. Clarke. Inequality constraints in the calculus of variations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):

528–540, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilman:1977:FQA

[1010] Robert Gilman. Finite quotients of the automorphism group of a free group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):541–551, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bonar:1977:CRT

[1011] D. D. Bonar, F. W. Carroll, and Peter Colwell. Category results for Tsuji functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):552–558, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fischer:1977:MDS

[1012] Pal Fischer and John A. R. Holbrook. Matrices doubly stochastic by blocks. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):559–577, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benabdallah:1977:QPI

[1013] Khalid Benabdallah and Adèle Laroche. Quasi-p-pure-injective groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 578–586, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DeTemple:1977:FNC

[1014] Duane W. De Temple and David B. Oulton. Formulas for the Nehari coefficients of bounded univalent functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):587–605, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marshall:1977:SLG

[1015] M. Marshall. Some local-global principles for formally real fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 606–614, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Magill:1977:H

[1016] Kenneth D. Magill, Jr. Homomorphisms from S(X) into S(Y). Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 615–625, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Oberlin:1977:TIO

[1017] Daniel M. Oberlin. Translationinvariant operators on L^{p}(G), 0 < p < 1. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):626–630, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fischer:1977:MSS

[1018] Pal Fischer and John A. R. Holbrook.

Matrices sous-stochastiques et fonctions convexes. (French) [Substochastic matrices and convex functions]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 631–637, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kubota:1977:TPK

[1019] K. K. Kubota. On a transcendence problem of K. Mahler: “Remarks on a paper by W. Schwarz” (J. Number Theory 1 (1969), 512–521). Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 638–647, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mura:1977:CRO

[1020] R. T. Botto Mura and A. H. Rhemtulla. A class of right-orderable groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):648–654, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

delJunco:1977:TSS

[1021] Andrés del Junco. A transformation with simple spectrum which is not rank one. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):655–663, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davis:1977:CTR

[1022] Gary Davis and Colin D. Fox. “Compatible tight Riesz orders on the group of automorphisms of an 0 − 2homogeneous set” (Canad. J. Math. 28 (1976), no. 5 1076–1081): addendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):664–665, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Foldes:1977:SGH

[1023] Stéphane Földes and Peter L. Hammer.

Split graphs having Dilworth number two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):666–672, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stanek:1977:CSS

[1024] Jean Chan Stanek. A characterization of starshaped sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):673–680, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wu:1977:GPP

[1025] Jang Mei G. Wu. Green’s potentials with prescribed boundary values. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):681–686, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mohamed:1977:WR

[1026] Saad Mohamed and Surjeet Singh.

Weak q-rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):687–695, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Okoh:1977:SPS

[1027] Frank Okoh. Systems that are purely simple and pure injective. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 696–700, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1977:CAS

[1028] P. M. Gauthier and W. Hengartner. Complex approximation and simultaneous interpolation on closed sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):701–706, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vincent:1977:TCC

[1029] Paul A. Vincent. A topological characterization of conjugate nets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 707–721, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boisen:1977:CEO

[1030] Monte B. Boisen, Jr. and Philip B. Sheldon. CPI-extensions: overrings of integral domains with special prime spectrums. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):722–737, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fawcett:1977:CFG

[1031] Barry Fawcett. A canonical factorization for graph homomorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 738–743, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Drury:1977:CVP

[1032] S. W. Drury. Changes of variables preserving Fourier–Stieltjes transforms on simply connected nilpotent Lie groups.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):744–755, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davis:1977:CTP

[1033] Philip J. Davis. Cyclic transformations of polygons and the generalized inverse. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):756–770, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burbea:1977:CMM

[1034] Jacob Burbea. The Carathéodory metric and its majorant metrics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 771–780, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butzer:1977:BTA

[1035] P. L. Butzer, H. Dyckhoff, E. Görlich, and R. L. Stens. Best trigonometric approximation, fractional order derivatives and Lipschitz classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 781–793, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1977:RTS

[1036] Nancy Johnson. Rice theorems for S_{n}^{−1} sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):794–805, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McDonald:1977:UP

[1037] David R. McDonald. On upcrossing probabilities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):806–818, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schmidt:1977:DMS

[1038] Wolfgang M. Schmidt. On the distribution modulo 1 of the sequence αn^{2}+βn. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (4):819–826, August 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stephenson:1977:CIE

[1039] Kenneth Stephenson. Certain integral equalities which imply equimeasurability of functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):827–844, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Atkinson:1977:RGH

[1040] M. D. Atkinson. On rank 3 groups having lambda = 0. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):845–847, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pettet:1977:SCS

[1041] Martin R. Pettet. A sufficient condition for solvability in groups admitting elementary abelian operator groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):848–855, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Casazza:1977:SCP

[1042] P. G. Casazza, C. A. Kottman, and Bor Luh Lin. On some classes of primary Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):856– 873, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Krause:1977:TKD

[1043] Günter Krause and Mark L. Teply. The transfer of the Krull dimension and the Gabriel dimension to subidealizers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):874–888, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See erratum [1146].

Hayden:1977:FGA

[1044] John L. Hayden and David L. Winter.

Finite groups admitting an automorphism trivial on a Sylow 2-subgroup. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):889–896, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Husch:1977:AAF

[1045] L. S. Husch. Approximating approximate fibrations by fibrations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 897–913, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chuchel:1977:CQR

[1046] John Chuchel and Norman Eggert.

The complete quotient ring of images of semilocal Prüfer domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 914–927, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:1977:WTQ

[1047] David Mordecai Cohen. Witt’s theorem for quadratic forms over nondyadic discrete valuation rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 928–936, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ong:1977:LTM

[1048] Hock Ong. Linear transformations on matrices: the invariance of a class of general matrix functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 937–946, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berman:1977:DOG

[1049] Gerald Berman. The dichromate and orientations of a graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 947–956, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Milgram:1977:EWS

[1050] R. James Milgram and Peter Zvengrowski. Even Whitehead squares are not projective. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):957–962, ????

1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1977:BSU

[1051] Mark A. Smith. Banach spaces that are uniformly rotund in weakly compact sets of directions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):963–970, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jackson:1977:DBG

[1052] D. M. Jackson and R. Aleliunas. Decomposition based generating functions for sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):971–1009, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Furuta:1977:RBG

[1053] Takayuki Furuta. Relations between generalized growth conditions and several classes of convexoid operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):1010–1030, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dobbs:1977:CIS

[1054] David E. Dobbs. On characterizing injective sheaves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1031–1039, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [1311].

Ash:1977:EF

[1055] Avner Ash. On eutactic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 1040–1054, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baxter:1977:EAC

[1056] J. R. Baxter and R. V. Chacon. Enlargement of σ-algebras and compactness of time changes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1055–1065, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Whitehead:1977:RJA

[1057] Jennifer Whitehead. Residuals of the join of ascendant subgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 1066–1068, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Drewnowski:1977:WBT

[1058] L. Drewnowski. The weak basis theorem fails in non-locally convex F -spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1069–1071, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jockusch:1977:SPS

[1059] Carl G. Jockusch, Jr. Simple proofs of some theorems on high degrees of unsolvability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1072–1080, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allegretto:1977:NCR

[1060] W. Allegretto. A note on a comparison result for elliptic equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 1081–1085, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Divinsky:1977:RP

[1061] N. Divinsky and A. Sulinski. Radical pairs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1086–1091, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Choi:1977:LPC

[1062] Man Duen Choi and Edward G. Effros. Lifting problems and the cohomology of C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1092–1111, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1977:JLD

[1063] Arlen Brown and Carl Pearcy. Jordan loops and decompositions of operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1112–1120, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kohli:1977:LTP

[1064] J. K. Kohli. Local topological properties of maps and open extensions of maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1121–1128, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Woods:1977:MSG

[1065] R. Grant Woods. Monocoreflective subcategories in general topology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 1129–1140, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harris:1977:FAS

[1066] Douglas Harris. Filter adjunction of spaces and compactifications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 1141–1144, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hung:1977:CTM

[1067] H. H. Hung. A contribution to the theory of metrization. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1145–1151, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heath:1977:HEC

[1068] Philip R. Heath. Homotopy equivalence of a cofibre fibre composite. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):1152–1156, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellers:1977:BCG

[1069] Erich W. Ellers. Bireflectionality in classical groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1157–1162, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Freud:1977:ASG

[1070] G. Freud and A. Sharma. Addendum: “Some good sequences of interpolatory polynomials” (Canad. J. Math. 26 (1974), 233–246). Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1163–1166, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Connett:1977:FSU

[1071] William C. Connett and Alan L.

Schwartz. A failure of stability under complex interpolation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1167– 1170, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Petrich:1977:CVQ

[1072] Mario Petrich. Certain varieties and quasivarieties of completely regular semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1171–1197, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ismail:1977:IRC

[1073] Mourad E. H. Ismail. Integral representations and complete monotonicity of various quotients of Bessel functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):1198–1207, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lada:1977:OAI

[1074] Thomas Lada. An operad action on infinite loop space multiplication. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 1208–1216, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lipsman:1977:UAR

[1075] Ronald L. Lipsman. On the unitarized adjoint representation of a semisimple Lie group. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1217–1222, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tymchatyn:1977:CHL

[1076] E. D. Tymchatyn. Compactification of hereditarily locally connected spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):1223–1229, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frankfurt:1977:QSS

[1077] Richard Frankfurt. Quasicyclic subnormal semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1230–1246, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lester:1977:CPM

[1078] J. A. Lester. Cone preserving mappings for quadratic cones over arbitrary fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1247–1253, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adams:1977:BEL

[1079] M. E. Adams and J. Sichler. Bounded endomorphisms of lattices of finite height. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1254–1263, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cressman:1977:QGR

[1080] R. Cressman. Quantization and group representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1264–1276, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chattopadhyay:1977:ECS

[1081] K. C. Chattopadhyay and W. J. Thron.

Extensions of closure spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 1277–1286, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hansen:1977:GEF

[1082] Lowell J. Hansen. On the growth of entire functions bounded on large sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29 (??):1287–1291, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Langlands:1977:SVS

[1083] R. P. Langlands. Shimura varieties and the Selberg trace formula. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 1292–1299, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:1977:DPC

[1084] John R. Martin. On a 1-dimensional planar continuum without the fixed point property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1300–1303, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deveney:1977:SII

[1085] James K. Deveney and John N. Mordeson. Subfields and invariants of inseparable field extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1304– 1311, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartley:1977:CLI

[1086] Richard Hartley and Kunio Murasugi. Covering linkage invariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??): 1312–1339, ???? 1977. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Luecke:1977:NC

[1087] Glenn R. Luecke. Norm convergence of T^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(6):1340–1344, December 1977. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alex:1978:IFS

[1088] Leo J. Alex and Dean C. Morrow. Index four simple groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):1–21, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hahn:1978:GUB

[1089] Kyong T. Hahn. Geometry on the unit ball of a complex Hilbert space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):22–31, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cook:1978:WCM

[1090] H. Cook and A. Lelek. Weakly confluent mappings and atriodic Suslinian curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):32–44, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davis:1978:BCP

[1091] Donald M. Davis. The BP-coaction for projective spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):45–53, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sultan:1978:MCR

[1092] Alan Sultan. Measure, compactification and representation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 54–65, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Behrens:1978:FPS

[1093] Ernst-August Behrens. Formal power series over commutative N-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):66–84, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lanski:1978:ISR

[1094] Charles Lanski. Invariant subrings in rings with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):85–94, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ratliff:1978:NLI

[1095] L. J. Ratliff, Jr. Notes on local integral extension domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):95–101, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kalton:1978:NDH

[1096] N. J. Kalton and G. V. Wood. Norm decreasing homomorphisms between ideals of L^{p}(G). Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):102–111, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grispolakis:1978:CRM

[1097] Joachim Grispolakis. Confluent and related mappings defined by means of quasi-components. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):112–132, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:1978:GGK

[1098] Mizan Rahman. A generalization of Gasper’s kernel for Hahn polynomials: application to Pollaczek polynomials.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):133–146, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kerman:1978:BCG

[1099] R. A. Kerman. Boundedness criteria for generalized Hankel conjugate transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):147–153, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Joseph:1978:CCH

[1100] James E. Joseph. Characterizations of completely hausdorff-closed spaces via graphs and projections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 154–160, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Greenfield:1978:NSS

[1101] Gary R. Greenfield. A note on subnormal subgroups of division algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):161–163, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rigo:1978:TEV

[1102] Thomas Rigo and Seth Warner.

Topologies extending valuations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 164–169, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1978:CSP

[1103] D. Borwein and W. Gawronski. On certain sequences of plus and minus ones. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):170–179, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hickin:1978:NIB

[1104] K. K. Hickin and R. E. Phillips. Non-isomorphic Burnside groups of exponent p^{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):180–189, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adams:1978:RIS

[1105] R. A. Adams and J. J. F. Fournier. Real interpolation of Sobolev spaces on subdomains of R^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):190–214, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1978:SGM

[1106] David F. Anderson. Subrings of k[X, Y] generated by monomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 215–224, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miller:1978:ETP

[1107] H. D. B. Miller. An extrapolation theorem for positive operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):

225–230, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isbell:1978:APA

[1108] J. R. Isbell, M. I. Klun, and S. H.

Schanuel. Affine parts of algebraic theories. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):231–237, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Friberg:1978:SHC

[1109] Bjorn Friberg. The stable homeomorphism conjecture in dimension four—an equivalent conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 238–242, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weiss:1978:CCS

[1110] William Weiss. Countably compact spaces and Martin’s axiom. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 243–249, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Matthiessen:1978:RSF

[1111] Günter Matthiessen. Regular and strongly finitary structures over strongly algebroidal categories. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 250–261, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1978:DOA

[1112] R. C. Brown. Differential operators with abstract boundary conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):262–288, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Promislow:1978:DNS

[1113] David Promislow. Dimension of null spaces with applications to group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):289–300, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hodel:1978:NCS

[1114] R. E. Hodel. The number of closed subsets of a topological space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 301–314, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruns:1978:FCO

[1115] Günter Bruns. A finiteness criterion for orthomodular lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 315–320, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mathon:1978:SCM

[1116] Rudolf Mathon. Symmetric conference matrices of order pq^{2} + 1. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):

321–331, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:1978:SCU

[1117] Q. I. Rahman and J. Szynal. On some classes of univalent polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 332–349, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Montgomery:1978:ISR

[1118] Susan Montgomery. Invariant subgroups in rings with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 350–357, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Freud:1978:APA

[1119] G. Freud, A. Giroux, and Q. I. Rahman. Sur l’approximation polynomiale avec poids exp(−midxmid). Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 358–372, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zimmer:1978:CEE

[1120] Robert J. Zimmer. Continuous ergodic extensions and fibre bundles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 373–391, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duffus:1978:SRF

[1121] D. Duffus, B. Jónsson, and I. Rival. Structure results for function lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):392–400, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1978:HDT

[1122] W. C. Brown. Higher derivations and tensor products of commutative rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):401–418, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1978:NHN

[1123] Marvin Marcus, Benjamin N. Moyls, and Ivan Filippenko. Normality and the higher numerical range. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 419–430, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Callahan:1978:SOI

[1124] T. Callahan. Some orbital integrals and a technique for counting representations of GL_{2}(F). Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):431–448, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roberts:1978:ECC

[1125] James W. Roberts. The embedding of compact convex sets in locally convex spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):449–454, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fawcett:1978:IFC

[1126] Barry Fawcett. On infinite full colourings of graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):455–457, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1978:HDP

[1127] P. D. Johnson, Jr. and R. N. Mohapatra. A Hardy–Davies–Petersen inequality for a class of matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 458–465, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shapiro:1978:ECC

[1128] H. L. Shapiro and F. A. Smith.

Even covers and collectionwise normal spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):466–473, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Faulkner:1978:CRA

[1129] G. D. Faulkner and R. W. Shonkwiler.

Cosine representations of abelian ∗- semigroups and generalized cosine operator functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):474–482, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ebert:1978:MSP

[1130] Gary L. Ebert. Maximal strictly partial spreads. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):483–489, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Farnum:1978:FR

[1131] Nicholas Farnum and Robert Whitley. Functionals on real C(S). Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 490–498, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morin-Strom:1978:WTN

[1132] Karl A. Morin-Strom. A Witt theorem for non-defective lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 499–511, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fischer:1978:DMO

[1133] Klaus G. Fischer. The decomposition of the module of n-th order differentials in arbitrary characteristic. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 512–517, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thomas:1978:ACL

[1134] Marc P. Thomas. Automatic continuity for linear functions intertwining continuous linear operators on Fréchet spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):518–530, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahe:1978:RFD

[1135] M. Rahe. Relatively finitely determined implies relatively very weak Bernoulli. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):531–548, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Buchsbaum:1978:GFR

[1136] David A. Buchsbaum. Generic free resolutions. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):549–572, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Whitehead:1978:CPS

[1137] Jennifer Whitehead. On certain properties of subnormal subgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 573–582, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nagel:1978:LBB

[1138] Alexander Nagel and Walter Rudin.

Local boundary behavior of bounded holomorphic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):583– 592, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Armacost:1978:UST

[1139] D. L. Armacost and W. L. Armacost.

Uniqueness in structure theorems for LCA groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):593–599, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:1978:PDS

[1140] S. J. Lee. Perturbation of direct sum differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 600–630, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

HarEl:1978:HMN

[1141] Zvi Har’El. Harmonic mappings of negatively curved manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 631–637, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bucchioni:1978:CPP

[1142] D. Bucchioni and A. Goldman. Sur la convergence presque partout des suites de fonctions mesurables. (French) [On the convergence almost everywhere of sets of measurable functions]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 638–649, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benabdallah:1978:TGE

[1143] Khalid Benabdallah and Robert Wilson. Thick groups and essentially finitely indecomposable groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):

650–654, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartley:1978:HI

[1144] Richard Hartley and Kunio Murasugi. Homology invariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 655–670, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Khurana:1978:CST

[1145] Surjit Singh Khurana. Corrigendum: “Strict topology on paracompact locally compact spaces” (Canad. J. Math. 29 (1977), no. 1, 216–219). Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):671, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [975].

Krause:1978:ETK

[1146] Günter Krause and Mark L. Teply.

Erratum: “The transfer of the Krull dimension and the Gabriel dimension to subidealizers” (Canad. J. Math. 29 (1977), no. 4, 874–888). Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):672, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [1043].

Hazewinkel:1978:AFP

[1147] Michiel Hazewinkel and Marcel Van de Vel. On almost-fixed-point theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):673–699, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chacron:1978:LAC

[1148] M. Chacron. Lie action of certain skews in ∗-rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):700–710, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Amar:1978:SIP

[1149] Éric Amar. Suites d’interpolation pour les classes de Bergman de la boule et du polydisque de C^{n}. (French) [Interpolation sets for Bergman classes of the ball and the polydisk of C^{n}]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 711–737, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maher:1978:LPC

[1150] David P. Maher. Lee polynomials of codes and theta functions of lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):738–747, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Labuda:1978:DCR

[1151] Iwo Labuda. Discrete copies of rings of sets in groups and Orlicz– Pettis theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):748–755, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Newman:1978:CMS

[1152] Morris Newman. Combinatorial matrices with small determinants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 756–762, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wilson:1978:RSR

[1153] Stephen E. Wilson. Riemann surfaces over regular maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):763–782, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cross:1978:AI

[1154] George Cross. Additivity of the P^{n}-integral. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):783–796, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duffus:1978:LPE

[1155] Dwight Duffus and Ivan Rival. A logarithmic property for exponents of partially ordered sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):797–807, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fridy:1978:MRS

[1156] J. A. Fridy. Minimal rates of summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):808–816, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Platzeck:1978:ASE

[1157] María Inés Platzeck. On algebras stably equivalent to an hereditary Artin algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):817–829, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Murase:1978:CAR

[1158] Ichiro Murase. A condition for Artinian rings to be Noetherian. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 830–837, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Neumann:1978:SGP

[1159] B. H. Neumann. Some group presentations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):838–850, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Snaith:1978:HGL

[1160] Victor Snaith. On the homology of the general linear groups over Z/4. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 851–855, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ebert:1978:BSP

[1161] Gary L. Ebert. Blocking sets in projective spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):856–862, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kronstadt:1978:ASC

[1162] Eric P. Kronstadt. Analytic structures for H^{∞} of certain domains in bfC^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):863–871, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kleinstein:1978:SSA

[1163] Jerrold L. Kleinstein and Alex Rosenberg. Signatures and semisignatures of abstract Witt rings and Witt rings of semilocal rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):872–895, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lindner:1978:CST

[1164] C. C. Lindner and A. Rosa. Corrigendum: “Steiner triple systems having a prescribed number of triples in common” (Canad. J. Math. 27 (1975), no. 5, 1166–1175). Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(4):896, August 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [800].

Bernier:1978:SCD

[1165] C. Bernier and A. Manitius. On semigroups in R^{n}×L^{p} corresponding to differential equations with delays. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 897–914, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rains:1978:UMP

[1166] M. Rains. On the upper majorant property for locally compact abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):915–925, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kannan:1978:CAR

[1167] V. Kannan and M. Rajagopalan. Constructions and applications of rigid spaces. III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):926–932, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Faber:1978:CGA

[1168] V. Faber, R. Laver, and R. McKenzie. Coverings of groups by abelian subgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):933–945, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Foldes:1978:RRF

[1169] Stéphane Földes and Navin M. Singhi. Regular representation of finite groups by hypergraphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):946–960, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Andrews:1978:SMM

[1170] Peter Andrews and Martin Arkowitz. Sullivan’s minimal models and higher order Whitehead products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):

961–982, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Balser:1978:LFS

[1171] W. Balser, W. B. Jurkat, and A. Peyerimhoff. On linear functionals and summability factors for strong summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):983–996, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meir:1978:ANR

[1172] A. Meir and J. W. Moon. On the altitude of nodes in random trees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 997–1015, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berman:1978:FGW

[1173] Gerald Berman. Families of generalized weighing matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1016–1028, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Green:1978:FAT

[1174] Edward L. Green. Frobenius algebras and their quivers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1029–1044, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gohberg:1978:RDO

[1175] I. Gohberg, P. Lancaster, and L. Rodman. Representations and divisibility of operator polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 1045–1069, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Khuri:1978:BER

[1176] Soumaya M. Khuri. Baer endomorphism rings and closure operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):1070–1078, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brungs:1978:ERC

[1177] H.-H. Brungs and G. Törner. Embedding right chain rings in chain rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):1079–1086, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Simon:1978:CBS

[1178] Udo Simon. Curvature bounds for the spectrum of closed Einstein spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):1087–1091, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feit:1978:PRM

[1179] Walter Feit and Jacques Tits. Projective representations of minimum degree of group extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1092–1102, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maxwell:1978:SGT

[1180] George Maxwell. The space groups of two dimensional Minkowski space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):1103–1120, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chacron:1978:CTR

[1181] M. Chacron. A commutativity theorem for rings with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 1121–1143, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bell:1978:CCS

[1182] Murray G. Bell. A cellular constraint in supercompact Hausdorff spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 1144–1151, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berg:1978:ITP

[1183] I. D. Berg. Index theory for perturbations of direct sums of normal operators and weighted shifts. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1152–1165, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kirwan:1978:NCS

[1184] W. E. Kirwan and R. W. Pell. A note on a class of slit conformal mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):1166–1173, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chapman:1978:CLF

[1185] T. A. Chapman. Constructing locally flat embeddings of infinitedimensional manifolds without tubular neighborhoods. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1174–1182, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koblitz:1978:SFJ

[1186] Neal Koblitz and David Rohrlich. Simple factors in the Jacobian of a Fermat curve. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1183–1205, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kan:1978:EPL

[1187] Charn Huen Kan. Ergodic properties of Lamperti operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1206–1214, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fialkow:1978:NUC

[1188] Lawrence A. Fialkow. A note on unitary cross sections for operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 1215–1227, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dlab:1978:RSQ

[1189] Vlastimil Dlab and Claus Michael Ringel. Real subspaces of a quaternion vector space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1228–1242, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roth:1978:UAM

[1190] Alice Roth. Uniform approximation by meromorphic functions on closed sets with continuous extension into the boundary. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1243–1255, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rival:1978:PSF

[1191] Ivan Rival and Bill Sands. Planar sublattices of a free lattice. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??): 1256–1283, ???? 1978. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feustel:1978:GCK

[1192] C. D. Feustel and Wilbur Whitten.

Groups and complements of knots. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):1284–1295, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dubinsky:1978:QSB

[1193] Ed Dubinsky and Boris Mitiagin. Quotient spaces without bases in nuclear Fréchet spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1296–1305, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Broverman:1978:TDI

[1194] S. Broverman, J. Ginsburg, K. Kunen, and F. D. Tall. Topologies determined by σ-ideals on ω_{1}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1306–1312, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vaughan:1978:FIS

[1195] N. H. Vaughan and R. W. Yeagy. Factoring ideals into semiprime ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (??):1313–1318, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hambleton:1978:PTD

[1196] I. Hambleton and R. J. Milgram. Poincaré transversality for double covers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1319–1330, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Phillips:1978:CBN

[1197] Thomas M. Phillips. Centered bases, nested bases, and completability of Aronszajn spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30(??):1331–1335, ???? 1978. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dokovic:1978:DAE

[1198] Dragomir Ž. – Doković. Derivations and automorphisms of exterior algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 30 (6):1336–1344, December 1978. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hager:1979:ERA

[1199] Anthony W. Hager and Lewis C.

Robertson. On the embedding into a ring of an Archimedean l-group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):1–8, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Forelli:1979:SER

[1200] Frank Forelli. Some extreme rays of the positive pluriharmonic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):9–16, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rival:1979:PSF

[1201] Ivan Rival and Bill Sands. Planar sublattices of a free lattice. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 17–34, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lindgren:1979:TUG

[1202] W. F. Lindgren and P. Fletcher. A theory of uniformities for generalized ordered spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):35–44, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vanstone:1979:ABE

[1203] S. A. Vanstone. The asymptotic behaviour of equidistant permutation arrays. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):45–48, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCoy:1979:PAG

[1204] Peter A. McCoy. Polynomial approximation and growth of generalized axisymmetric potentials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):49–59, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harary:1979:LBB

[1205] F. Harary and R. W. Robinson. Labeled bipartite blocks. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):60–68, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Day:1979:CFL

[1206] Alan Day. Characterizations of finite lattices that are boundedhomomorphic images of sublattices of free lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):69–78, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McDonald:1979:MIS

[1207] Gerard McDonald. The maximal ideal space of H^{∞} + C on the ball in C^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):79–86, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jorgensen:1979:CDC

[1208] Troels Jørgensen. Comments on a discreteness condition for subgroups of SL(2, C). Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):87–92, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1979:RCT

[1209] P. F. Smith. Rings characterized by their cyclic modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):93–111, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoare:1979:CGW

[1210] A. H. M. Hoare. Coinitial graphs and Whitehead automorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 112–123, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Abib:1979:QPN

[1211] Odinete Renée Abib. Quelques propriétés sur le normalisateur d’un pseudogroupe de Lie transitif. (French) [Some properties of the normalizer of a transitive Lie pseudogroup.]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 124–129, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grispolakis:1979:ESD

[1212] J. Grispolakis and E. D. Tymchatyn. Embedding smooth dendroids in hyperspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):130–138, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thomas:1979:ICC

[1213] M. P. Thomas. Invariant complements to closed invariant subspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 139–147, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Olin:1979:LCS

[1214] Robert F. Olin and James E. Thomson. Lifting the commutant of a subnormal operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):148–156, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rossmann:1979:SSS

[1215] W. Rossmann. The structure of semisimple symmetric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 157–180, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schiff:1979:CAF

[1216] J. L. Schiff. On a class of analytic functions of Smirnov. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):181–183, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chong:1979:IRT

[1217] Kong Ming Chong. An inductive rearrangement theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):184–188, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rich:1979:SRP

[1218] Michael Rich. Some radical properties of Jordan matrix rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 189–196, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grispolakis:1979:SPH

[1219] J. Grispolakis, Sam B. Nadler, Jr., and E. D. Tymchatyn. Some properties of hyperspaces with applications to continua theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):197–210, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sherk:1979:ISA

[1220] F. A. Sherk. Indicator sets in an affine space of any dimension. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 211–224, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mattson:1979:DSP

[1221] Don A. Mattson. Discrete subsets of proximity spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):225–230, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heath-Brown:1979:DZD

[1222] D. R. Heath-Brown. The density of zeros of Dirichlet’s L-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 231–240, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Devlin:1979:NNM

[1223] Keith J. Devlin and Saharon Shelah. A note on the normal Moore space conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):241–251, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morgan:1979:CTS

[1224] C. L. Morgan. On counting types of symmetries in finite unitary reflection groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):252–254, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ditzian:1979:GIT

[1225] Z. Ditzian. On global inverse theorems of Szász and Baskakov operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):255–263, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gervais:1979:CST

[1226] Jean-Jacques Gervais. Critères de Gstabilité en termes de transversalité. (French) [G-stability criteria in terms of transversality]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):264–273, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reich:1979:CPG

[1227] Daniel Reich. On certain polynomials of Gaussian type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):274–281, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jacobs:1979:SNS

[1228] Barry E. Jacobs. α-speedable and nonα-speedable sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):282–299, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Spiegel:1979:SSA

[1229] Eugene Spiegel and Allan Trojan. The Schur subgroup of a p-adic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 300–303, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glass:1979:CTR

[1230] A. M. W. Glass. Compatible tight Riesz orders. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):304–307, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Winter:1979:SST

[1231] David J. Winter. Solvable subgroups and their Lie algebras in characteristic p. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):308–311, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barrett:1979:AAL

[1232] Wayne W. Barrett. An alternative approach to Laguerre polynomial identities in combinatorics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):312– 319, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marshall:1979:CFS

[1233] Murray Marshall. Classification of finite spaces of orderings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 320–330, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sember:1979:USP

[1234] John Sember and Marc Raphael. The unrestricted section properties of sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):331–336, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Castro:1979:SDP

[1235] Alfonso Castro. A semilinear Dirichlet problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):337–340, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brownawell:1979:MLI

[1236] W. Dale Brownawell. A measure of linear independence for some exponential functions. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):341–346, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Giroux:1979:BI

[1237] A. Giroux, Q. I. Rahman, and G. Schmeisser. On Bernstein’s inequality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):347–353, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellers:1979:FA

[1238] Erich W. Ellers. Factorization of affinities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):354–362, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lewis:1979:SHP

[1239] Wayne Lewis. Stable homeomorphisms of the pseudo-arc. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):363–374, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cassidy:1979:LHG

[1240] C. Cassidy and P. J. Hilton. L’isolateur d’un homomorphisme de groupes. (French) [The insulator of a group homomorphism]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):375–391, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lester:1979:TSW

[1241] J. A. Lester. Transformations of nspace which preserve a fixed squaredistance. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):392–395, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bisztriczky:1979:BSO

[1242] Tibor Bisztriczky. Biplanar surfaces of order three. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):396–418, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allegretto:1979:LBN

[1243] Walter Allegretto. Lower bounds on the number of points in the lower spectrum of elliptic operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 419–426, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shar:1979:LALa

[1244] Albert O. Shar. Localization, algebraic loops and H-spaces. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 427–435, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

May:1979:MGU

[1245] Warren May. Multiplicative groups under field extension. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):436–440, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fong:1979:RSE

[1246] Humphrey Fong. Ratio and stochastic ergodic theorems for superadditive processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):441–447, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butler:1979:EEC

[1247] G. J. Butler. Erratum: “The existence of continuable solutions of a second order differentiable equation”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):448, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [1005].

Anderson:1979:GR

[1248] D. D. Anderson and J. Matijevic.

Graded pi-rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):449–457, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Babai:1979:HCG

[1249] László Babai, Peter Frankl, János Kollár, and Gert Sabidussi. Hamiltonian cubic graphs and centralizers of involutions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):458–464, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Macdonald:1979:RDB

[1250] Roderick N. S. Macdonald. Representable dualities between finitely closed subcategories of modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):

465–475, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Montgomery:1979:MVC

[1251] H. L. Montgomery and R. C. Vaughan. Mean values of character sums. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 476–487, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jackson:1979:SEH

[1252] D. M. Jackson and I. P. Goulden. Sequence enumeration and the de Bruijn– van Aardenne–Ehrenfest–Smith–Tutte theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):488–495, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fack:1979:CSA

[1253] Th. Fack and O. Maréchal. Sur la classification des symétries des C^{∗}- algèbres UHF. (French) [On the symmetry classifications of C^{∗}-algebras]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):496–523, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baumgartner:1979:CPC

[1254] J. Baumgartner, P. Erdős, F. Galvin, and J. Larson. Colorful partitions of cardinal numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):524–541, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chacron:1979:USA

[1255] M. Chacron. Unitaries in simple Artinian rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):542–557, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fisher:1979:CCM

[1256] Joe W. Fisher. Chain conditions for modular lattices with finite group actions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):558–564, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bernstein:1979:PAT

[1257] Leon Bernstein. Periodic algorithms and their application. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):565– 595, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McClure:1979:ISD

[1258] J. P. McClure and R. Wong. Infinite systems of differential equations. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):596–603, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marshall:1979:QIL

[1259] Murray A. Marshall. Quotients and inverse limits of spaces of orderings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):604–616, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Drake:1979:PGG

[1260] David A. Drake. Partial lambdageometries and generalized Hadamard matrices over groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):617– 627, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sullivan:1979:GUR

[1261] Francis Sullivan. A generalization of uniformly rotund Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):628–636, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McLeish:1979:NCA

[1262] Mary McLeish. On the number of conjugates of n-ary quasigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 637–654, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deveney:1979:OIF

[1263] James K. Deveney and John N. Mordeson. The order of inseparability of fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):655–662, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Henson:1979:ASB

[1264] C. Ward Henson. Analytic sets, Baire sets and the standard part map. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):

663–672, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Junnila:1979:CPF

[1265] Heikki J. K. Junnila. On countability of point-finite families of sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 673–679, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Knuppel:1979:NRN

[1266] Frieder Knüppel and Michael Kunze.

Neighbor relation and neighbor homomorphism of Hjelmslev groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 680–699, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Langlands:1979:SCD

[1267] R. P. Langlands. Stable conjugacy: definitions and lemmas. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):700–725, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Labesse:1979:I

[1268] J.-P. Labesse and R. P. Langlands. Lindistinguishability for SL(2). Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 726–785, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ngoc:1979:RDA

[1269] Nghiêm – D˘ a.ng Ngo.c. A remark on disintegrations with almost all components non-σ-additive. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):786– 788, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1979:NLR

[1270] D. D. Anderson, E. W. Johnson, and J. A. Johnson. Noether lattices representable as quotients of the lattice of monomially generated ideals of polynomial rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):789–799, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baird:1979:ISH

[1271] Bridget B. Baird. Inverse semigroups of homeomorphisms are Hopfian. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 800–807, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reiner:1979:LIM

[1272] Irving Reiner. Lifting isomorphisms of modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):808–811, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shar:1979:LALb

[1273] Albert O. Shar. Localization, algebraic loops and H-spaces. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 812–817, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanMill:1979:EES

[1274] Jan van Mill. Not every K_{1}-embedded subspace is K_{0}-embedded. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 818–823, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Farkas:1979:BCL

[1275] Daniel R. Farkas. Baire category and Laurent extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):824–830, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nyman:1979:SAR

[1276] T. Nyman and G. Whaples. Simple algebras over rational function fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):831–835, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Repka:1979:TPH

[1277] Joe Repka. Tensor products of holomorphic discrete series representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):836–844, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fong:1979:OI

[1278] C. K. Fong and A. R. Sourour. On the operator identity ∑ A_{k}XB_{k} ≡ 0. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):845–857, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wade:1979:SUG

[1279] William R. Wade. Sets of uniqueness for the group of integers of a p-series field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):858–866, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Choi:1979:SAG

[1280] Man Duen Choi. A simple C^{∗}-algebra generated by two finite-order unitaries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):867–880, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilligan:1979:RLC

[1281] B. Gilligan and A. Huckleberry. Remarks on k-Leviflat complex manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):881–889, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Choo:1979:STS

[1282] Seki A. Choo. Strict topology on spaces of continuous vector-valued functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):890–896, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stewart:1979:IDS

[1283] C. L. Stewart and R. Tijdeman. On infinite-difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):897– 910, ???? 1979. CODEN CJMAAB.

ISSN 0008-414X (print), 1496-4279 (electronic).

vanDouwen:1979:BDN

[1284] Eric K. van Douwen. A basically disconnected normal space Φ with |βΦ − Φ| = 1. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(5):911–914, October 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steiner:1979:GUT

[1285] R. J. Steiner. The group of units in K-theory modulo an odd prime. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 915–928, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benkart:1979:LAN

[1286] G. M. Benkart and I. M. Isaacs. Lie algebras with nilpotent centralizers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):929–941, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Auslander:1979:ASS

[1287] M. Auslander, R. Bautista, M. I.

Platzeck, I. Reiten, and S. O. Smalø. Almost split sequences whose middle term has at most two indecomposable summands. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):942–960, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruns:1979:BFO

[1288] Günter Bruns. Block-finite orthomodular lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):961–985, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1979:COM

[1289] W. T. Tutte. Combinatorial oriented maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):986–1004, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pokrass:1979:LRC

[1290] David Pokrass. Levitzki radical for certain varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):1005–1011, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Phillips:1979:PAA

[1291] John Phillips and Iain Raeburn. Perturbations of AF-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 1012–1016, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:1979:EQT

[1292] John R. Martin and Sam B. Nadler, Jr. Examples and questions in the theory of fixed-point sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):1017–1032, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Millet:1979:CVC

[1293] Annie Millet and Louis Sucheston.

Characterizations of Vitali conditions with overlap in terms of convergence of classes of amarts. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(5):1033– 1046, October 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cordes:1979:QFQ

[1294] Craig M. Cordes and John R. Ramsey, Jr. Quadratic forms over quadratic extensions of fields with two quaternion algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):1047–1058, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herrlich:1979:DFS

[1295] H. Herrlich, G. Salicrup, and R. Vazquez.

Dispersed factorization structures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):1059–1071, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Niimura:1979:UT

[1296] Mikio Niimura. On a uniqueness theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):1072–1076, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putcha:1979:SPS

[1297] Mohan S. Putcha. Subgroups of the power semigroup of a finite semigroup. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):1077–1083, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lemire:1979:FIR

[1298] F. Lemire and M. Pap. H-finite irreducible representations of simple Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):1084–1106, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Greiner:1979:FSN

[1299] Peter C. Greiner. A fundamental solution for a nonelliptic partial differential operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):1107–1120, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Langlands:1979:ZFS

[1300] R. P. Langlands. On the zeta functions of some simple Shimura varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):1121–1216, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gunther:1979:DRO

[1301] Georg Günther and Wolfgang Nolte. Defining relations in orthogonal groups of characteristic two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):1217–1246, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kato:1979:URS

[1302] Akio Kato. Union of realcompact spaces and Lindelöf spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(6): 1247–1268, December 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burbea:1979:PBS

[1303] Jacob Burbea. Projections of Bergman spaces over plane domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 1269–1280, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stewart:1979:FTU

[1304] James Stewart. Fourier transforms of unbounded measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):1281– 1292, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hagelgans:1979:LST

[1305] Nancy L. Hagelgans. Local spaces with three cells as H-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??): 1293–1306, ???? 1979. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Perkel:1979:BVP

[1306] Manley Perkel. Bounding the valency of polygonal graphs with odd girth. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):1307–1321, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gardner:1979:ICT

[1307] L. Terrell Gardner. An inequality characterizes the trace. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(??):1322–1328, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brunner:1979:GWE

[1308] A. M. Brunner and R. G. Burns. Groups in which every finitely generated subgroup is almost a free factor. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31 (??):1329–1338, ???? 1979. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigenda [1310].

Jafarian:1979:AIN

[1309] Ali A. Jafarian. Algebras intertwining normal and decomposable operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 31(6):1339–1334, December 1979. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brunner:1980:CGW

[1310] A. M. Brunner and R. G. Burns. Corrigenda: “Groups in which every finitely generated subgroup is almost a free factor”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(??):766, ???? 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [1308].

Dobbs:1980:CCI

[1311] David E. Dobbs. Corrigendum:

“On characterizing injective sheaves”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32 (??):1522, ???? 1980. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [1054].

Orzech:1980:CBG

[1312] Morris Orzech. Correction to: “On the Brauer group of algebras having a grading and an action”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(6): 1523–1524, December 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [874].